### v1

# A Causal Framework to Measure and Mitigate Non-binary Treatment Discrimination

Abstract
Fairness studies of algorithmic decision-making systems often simplify complex decision processes, such as bail or loan approvals, into binary classification tasks. However, these approaches overlook that such decisions are not inherently binary (e.g., approve or not approve bail or loan); they also involve non-binary treatment decisions (e.g., bail conditions or loan terms) that can influence the downstream outcomes (e.g., loan repayment or reoffending). In this paper, we argue that non-binary treatment decisions are integral to the decision process and controlled by decision-makers and, therefore, should be central to fairness analyses in algorithmic decision-making. We propose a causal framework that extends fairness analyses and explicitly distinguishes between decision-subjects’ covariates and the treatment decisions. This specification allows decision-makers to use our framework to (i) measure treatment disparity and its downstream effects in historical data and, using counterfactual reasoning, (ii) mitigate the impact of past unfair treatment decisions when automating decision-making. We use our framework to empirically analyze four widely used loan approval datasets to reveal potential disparity in non-binary treatment decisions and their discriminatory impact on outcomes, highlighting the need to incorporate treatment decisions in fairness assessments. Moreover, by intervening in treatment decisions, we show that our framework effectively mitigates treatment discrimination from historical data to ensure fair risk score estimation and (non-binary) decision-making processes that benefit all stakeholders.
**footnotetext: These authors contributed equally to this work
Keywords Algorithmic Fairness  $\cdot$ Causality  $\cdot$ Counterfactual  $\cdot$ Decision-making  $\cdot$ Discrimination  $\cdot$ Path-specific
1 Introduction
Table 1: Examples of data collected across various decision-making domains, highlighting the distinction between sensitive attributes $S$, decision-subject covariates $X$, treatment decisions $Z$, and outcomes $Y$. The multivariate, non-binary $Z$ have been represented previously in the literature as a positive binary decision (in gray).
Domain Sensitives $S$ Covariates $X$
Treatment
Decisions $Z$
Outcomes $Y$
Binary
Decisions
Lending gender, age, race
salary, savings,
employment history…
loan terms (amount,
duration, interest rate…)
loan repaid loan approved
Justice
gender, race, age,
socioeconomic status
criminal history,
type of crime…
bail terms (bail amount,
conditions…)
re-offence bail approved
Health
gender, age,
race, disability
medical history,
lab results…
medical treatment
(medication, surgery…)
re-hospitalization hospitalization
Insurance gender, age, race
goods price, salary,
employment history…
insurance terms
(cost, coverage…)
claim approval policy approved
Hiring gender, age, race
education, experience,
qualifications…
employment terms
(working hours, salary…)
job performance job offered
Data-driven approaches increasingly leverage historical data to automate decision-making processes across societal domains, such as finance [4, 49], healthcare [20], and criminal justice [13]. These systems typically reduce complex decision-making to a binary classification task, predicting either ground-truth outcomes (e.g., loan repayment, recidivism after bail) or imperfect proxies (e.g., healthcare expenditures as a proxy for health status) [44, 53]. However, real-world decisions are rarely as simple as a binary choice; they involve multiple, non-binary treatment decisions. As shown in Table 1, decision-makers often determine not only whether to give a positive decision (provide treatment, e.g., approve loan) but also its specific terms (e.g., loan amount, duration, interest rate). Similar oversimplifications occur in other domains, such as bail decisions [61]. This simplistic approach overlooks the impact of non-binary treatment decisions on outcomes, leading to fairness analyses that fail to provide a holistic view of existing systems [46].
Despite growing research on fairness in algorithmic decision-making, existing analyses simplify complex decision-making mechanisms in three key ways. First, many (if not most) studies [69, 21, 11] include non-binary treatments as covariates of the decision-subject (covariate examples in Table 1). This assumption obscures the mechanisms that assign treatment decisions, which are not controlled by decision-subjects but by decision-makers and, thus, may reveal discriminatory patterns. This assumption thus limits the ability of these works to account for all sources of unfairness and their implications. Second, alternative works [39, 12, 57] examine the fairness implications of treatment decisions but assume they are binary, ignoring the multivariate, non-binary nature of treatment decisions (such as, e.g., loan or bail amount). Finally, only a few works [2, 17, 3] have focused on disparities in non-binary treatment decisions across different groups. Still, they lack a generalizable framework, providing only empirical insights in financial domains and, more importantly, neglecting the discriminatory downstream effects of these disparities on outcomes.
To bridge gaps in the existing approaches, we propose a fundamental paradigm shift in fairness analysis of automated decision-making systems that brings fairness considerations to non-binary treatment decisions and their downstream outcome effects. Our paradigm shift leverages the following key insight: decision-makers do not simply make binary decisions (whether to give treatment) but also control the assignment of adjacent non-binary treatment decisions, which can introduce disparities and negatively impact outcomes. To incorporate this insight into fairness analysis, we introduce a broadly applicable causal framework that allows decision-makers to analyze the fairness of underlying decision-making mechanisms systematically. This systematic analysis allows our framework to provide novel measures of treatment disparities, assess their discriminatory effects on outcomes, and enable mitigation strategies to prevent biases in non-binary treatment decisions from perpetuating unfairness in automated systems.
Our approach explicitly differentiates between a decision-subject’s covariates and a decision-maker’s treatment decisions in the causal data-generating process, as illustrated in Figure 1. This causal formulation allows us to measure disparities in non-binary treatment decisions $Z$ induced by sensitive attributes $S$, both directly (solid red path) and indirectly through covariates $X$ (dotted red path). Furthermore, our framework quantifies how these disparities propagate to outcomes $Y$ via the causal path from $Z$ to $Y$ (dashed orange path), thereby inducing treatment discrimination. Our framework supports mitigation in scenarios impacted by treatment discrimination through two pre-processing strategies. These techniques generate treatment-fair counterfactual datasets, enabling fairer non-binary automated decisions and more equitable risk score estimation to inform future decision-making.
Our framework advances fairness analyses by leveraging path-specific counterfactuals to quantify direct and indirect disparities in non-binary treatment decisions and their impact on outcomes. While path-specific counterfactuals provide a powerful tool for causal fairness analyses, existing methods [52, 8] suffer from significant limitations, such as being restricted to linear data generation processes or requiring inefficient Monte Carlo estimation with strong regularization. To overcome these challenges, we build on the state-of-the-art in causal inference, the causal normalizing flows [26], and propose a novel path-specific counterfactual generation method. This advancement enables our framework to be theoretically grounded while being practical to be applied in real-world decision-making systems.
Finally, we validate our framework using four real-world lending datasets across two decision-making scenarios. Our empirical results demonstrate that explicitly accounting for non-binary treatment decisions helps provide more holistic fairness assessments, identifies systemic biases, and supports mitigation efforts. These findings underscore the importance of addressing treatment discrimination in achieving fairer outcomes, bridging a critical gap in fairness research, and laying the foundation for more robust causal fairness analyses.
$X$$Z$$Y$$S$
Figure 1: Causal graph of the data generation process, distinguishing decision-subject covariates $X$ and treatment decisions $Z$. Under the assumption that a positive binary decision has already been made. Red paths (solid: direct, dotted: indirect) show causal paths to $Z$ influenced by sensitive attribute $S$, highlighting potential sources of disparities. The orange dashed path from $Z$ to $Y$ captures the potentially discriminatory effects on the outcome.
2 Background
Our framework for measuring and mitigating treatment discrimination in decision-making relies on causal analyses. This section briefly outlines the core causal concepts used throughout the paper.
Structural causal models (SCMs). An SCM [56] $\mathcal{M}=(V,U,F)$ is defined by the observational features $V$, independent, unobserved, exogenous variables $U$ (no hidden confounding) and structural equations $F$. A structural equation for some $V_{i}$ gives the functional mapping from the corresponding exogenous $U_{i}$ and causal parents $\operatorname{pa}(V_{i})$ to the value of $V_{i}$ as $V_{i}=F_{i}(\operatorname{pa}(V_{i}),U_{i})$. These cause-effect relations in $\mathcal{M}$ can also be represented by a directed acyclic graph $\mathcal{G}$ like in Figure 1, where the nodes represent $V$, and the directed edges represent the cause-effect relations.
Performing interventions. The SCM $\mathcal{M}$ helps us estimate the effects of hypothetical interventions $\mathcal{I}$ on features. For instance, to analyze the effect of the intervention $\mathcal{I}$ setting $V_{i}=\alpha$, we use the ${\operatorname{do}}$-operator [56] ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ on $\mathcal{M}$ that generates the intervened SCM $\mathcal{M}^{\mathcal{I}}$.
Counterfactuals. In the absence of hidden confounders, this intervened SCM $\mathcal{M}^{\mathcal{I}}$ can also be used to perform counterfactual analyses for a particular factual decision-subject with features $\vartheta^{F}$ to ask “what would be the counterfactual features $\vartheta^{CF}$ had we intervened with ${\operatorname{do}}(V_{i}\rightarrow\alpha)$, and all exogenous factors $u^{F}$ remaining unchanged.” Counterfactuals $\vartheta^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))$ are computed following the three steps [56]: i). abduction: estimating $u^{F}$ from $\vartheta^{F}$ using $\mathcal{M}$, ii). action: performing ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ to get $\mathcal{M}^{\mathcal{I}}$, iii). prediction: compute $\vartheta^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))$ from $\mathcal{M}^{\mathcal{I}}$ using $u^{F}$ and setting $V_{i}=\alpha$. In simple terms, computing counterfactuals estimates the total effect of an intervention on the downstream causal features of the intervened variable in the causal graph $\mathcal{G}$ for the factual decision-subject $\vartheta^{F}$.
Path-specific counterfactuals. In certain scenarios, it is crucial to understand the effects of a particular feature $V_{i}$ on a downstream feature $V_{j}$ along different causal paths in $\mathcal{G}$. This analysis can be done by computing path-specific counterfactuals [8], which use multiple ${\operatorname{do}}$ operations to control for specific causal paths. Consider the two paths in an example causal structuring between variables $\langle V_{i},V_{j},V_{k}\rangle$, where we have $V_{i}\rightarrow V_{k}$ and $V_{i}\rightarrow V_{j}\rightarrow V_{k}$. For a factual instance $\vartheta^{F}=\langle v_{i}^{F},v_{j}^{F},v_{k}^{F}\rangle$, we can compute the path-specific counterfactual effect of ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ along these paths. For the path $V_{i}\rightarrow V_{k}$, we measure the counterfactual effect using $v_{k}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha),{\operatorname{do}}(V_{j}\rightarrow v_{j}^{F}))$. Similarly, for the path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$, we use $v_{k}^{CF}(v_{i}^{F},{\operatorname{do}}(V_{j}\rightarrow v_{j}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))))$. For the path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$, we use the counterfactual $v_{j}^{CF}$ obtained from the intervention on $V_{i}$ but keep $V_{i}=v_{i}^{F}$ fixed along the other path.
Causal fairness analyses. The SCM framework has been extensively used to causally study fairness problems in ML [35, 8]. In these scenarios, target labels $Y$ (e.g., bail, loan repayment) are considered causally affected by the sensitive feature $S$ (e.g., gender/sex, race). We can estimate the total effect of $S$ on $Y$ by finding sensitive counterfactuals $y^{\mathrm{SCF}}=y^{CF}({\operatorname{do}}(S\rightarrow\alpha))$, as was shown in counterfactual fairness [35]. Similarly, Chiappa [8] used path-specific counterfactuals $y^{\mathrm{PSCF}}$ to analyze the potentially discriminatory effect of $S$ on $Y$ along different causal paths in $\mathcal{G}$ for linear SCMs. In this work, we extend the concept of counterfactual fairness and path-specific counterfactual fairness for treatment decisions and their downstream effects.
3 Measuring treatment disparities and their downstream effects
In this section, we introduce our causal framework to address the limitations of the existing fair decision-making literature by explicitly measuring disparities in treatment decisions and their downstream effects and mitigating potential discriminatory consequences. To this end, we assume that the data generation process that produces the observed (factual) historical data $\mathcal{D}$ follows the causal graph depicted in Figure 1, where:
  1. 1.
    The treatment decision $Z$ is a causal child of both $X$ and (potentially) $S$. We assume that the causal mechanism that generated $Z$ is under the control of the decision-maker (e.g., we assume that banks decide not only whether or not to grant a loan but also the loan terms) and may be discriminatory. The direct causal path from $S$ to $Z$ is shown in solid red, while the indirect path mediated by $X$ is shown in dashed red.
  2. 2.
    The treatment decision $Z$ is a causal parent of the outcome variable $Y$ (e.g., we assume loan terms can impact the repayment ability of applicants). Hence, $S$ can influence $Y$ mediated by $Z$ (orange dashed line in Figure 1), leading to the possible propagation of treatment disparities to the outcome. Moreover, we maintain the direct $S\rightarrow Y$ since such effects may occur in specific scenarios, e.g., in healthcare.
Remarks. The edges of our causal graph $\mathcal{G}$ (Figure 1) indicate the structural causal ordering [26] among the different groups of features, where the cause-effect relation for some edges may be zero. Moreover, although the figure shows a single sensitive feature for simplicity, our $\mathcal{G}$ can consider multiple sensitive features as independent root nodes. We treat $X$ and $Z$ in $\mathcal{G}$ as multivariate blocks, focusing our causal analyses on relations between these “blocks” without making explicit assumptions about their causal relations.
3.1 Disparities in non-binary treatment decisions
Given our proposed causal graph in Figure 1, we use the SCM framework to measure disparities in treatment decisions by measuring counterfactual and path-specific counterfactual effects of the sensitive feature (e.g., gender) on the treatment decision variables. For simplicity and without loss of generalization, here we assume the sensitive feature on which we intervene to be binary, i.e., $S\in\{0,1\}$.  [^1] Our measures can easily be extended for multi-valued categorical sensitive attributes, e.g., by conducting group-pairwise comparisons. Intersectional disparity measurements across multiple sensitive attributes are left for future work. Given an observed dataset $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$, disparity function $\mathcal{F}$, and aggregation function $\mathcal{A}_{\mathcal{D}}$ regarding $\mathcal{D}$, we define two treatment disparity measures.
Definition 1.
Total Treatment Disparity (TTD), against decision-subjects with (factual) sensitive feature $s^{F}$, measures the expected difference between the treatment $z^{F}$ they received and the treatment $z^{\mathrm{SCF}}$ they would have received if they belonged to a different sensitive group, i.e.,
$\operatorname{TTD}(s^{F})=\mathcal{A}_{\mathcal{D}}\left(\mathcal{F}(z^{\mathrm{SCF}},z^{F})\right),\,\mathrm{where}\,z^{\mathrm{SCF}}=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})))$ (1)
Related to [35], the ${\operatorname{do}}$-operator on the sensitive feature $S$ gives the sensitive counterfactual treatment $z^{\mathrm{SCF}}$ a decision-subject would have received if their sensitive group were different.
Definition 2.
Direct Treatment Disparity (DTD), against decision-subjects with (factual) sensitive feature $s^{F}$, measures the expected difference between the treatment they received and the sensitive (direct) - path counterfactual treatment, which assumes a ${\operatorname{do}}(S\rightarrow(1-s^{F})))$ only on the direct path from $S$ to $Z$, i.e.,
$\begin{split}\operatorname{DTD}(s^{F})=\mathcal{A}_{\mathcal{D}}\left(\mathcal{F}(z^{\mathrm{SPCF}},z^{F})\right),\,\mathrm{where}\,z^{\mathrm{SPCF}}=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})),{\operatorname{do}}(X\rightarrow x^{F}))\end{split}$ (2)
Leveraging path-specific counterfactuals [52, 8], we compute the sensitive path counterfactual treatments $z^{\mathrm{SPCF}}$ a decision-subject would have received for different sensitive group memberships while keeping original covariates $x^{F}$ fixed.
Remarks. First, the nature of $\mathcal{A},\mathcal{F}$ in the above definitions depends on the nature of the treatment decision variables. For instance, when treatment decisions are continuous, $\mathcal{A}$ would be the expectation $\mathbb{E}$ over $\mathcal{D}$, and $\mathcal{F}$ would denote the differences across counterfactual treatments. Second, by drawing parallels with predictive fairness concepts, we can better understand the fairness implications of treatment disparities. For instance, $\operatorname{DTD}$ highlights disparities caused by directly incorporating $S$ into treatments, which may be acceptable in areas such as healthcare but undesirable in contexts such as lending, aligning with the predictive fairness concept of disparate treatment [15]. Similarly, zero $\operatorname{TTD}$ requires that treatments $Z$ be independent of $S$, showing parallels with demographic parity [15], which demands that predictions do not depend on $S$. However, like predictive fairness, the relevance of perfect treatment parity depends on context. Similar to demographic parity scenarios, when ground-truth outcome $Y$ is unavailable (e.g., admissions, hiring), treatment disparities may be considered discriminatory. On the other hand, as separation-based frameworks [21, 69] show, the outcome $Y$ can explain predictive disparities in certain contexts, e.g., when $Y$ is a ground truth as in lending. Hence, studying the downstream impact on $Y$ is imperative to understand if treatment disparities are discriminatory.
3.2 Downstream effects of treatment disparities
While Definitions 1 and 2 help measure treatment disparities, it is crucial to assess the downstream impacts of these disparities on the outcome $Y$ (e.g., when $Y$ is a ground truth as in lending) to determine whether they are discriminatory. Hence, we introduce two novel measures to quantify the counterfactual downstream outcome effects of total and direct treatment disparities.
Definition 3.
Total Treatment Disparity Effect (TTD-E) measures the downstream effect of the sensitive counterfactual treatment $z^{\mathrm{SCF}}$, computed in Eq. 1, as the probability of a label change from the factual value $y^{F}$, i.e.,
$\begin{split}\operatorname{TTD-E}(s^{F},y^{F})&=\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}\left(y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))\neq y^{F}\right)\right],\end{split}$ (3)
where $\mathbb{I}$ is an indicator function. This definition helps us understand how the total causal effect of $S$ on the treatment decisions $Z$ ends up impacting the decision outcomes $Y$. Intuitively, for any decision-subject belonging to $s^{F}$ with covariates $x^{F}$, the decision-maker provides the sensitive counterfactual $z^{\mathrm{SCF}}$ treatments to measure the outcome change.
Definition 4.
Direct Treatment Disparity Effect (DTD-E) measures the downstream effect of the sensitive path counterfactual treatment $z^{\mathrm{SPCF}}$, computed in Eq. 2, as the probability of a label change from the factual value $y^{F}$, i.e.,
$\operatorname{DTD-E}(s^{F},y^{F})$ $=\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}\left(y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SPCF}}))\neq y^{F}\right)\right],$ (4)
where $\mathbb{I}$ is an indicator function. This definition helps us understand how the direct causal effect of $S$ on the treatment decisions $Z$ ends up impacting the decision outcomes $Y$. Intuitively, for any decision-subject belonging to $s^{F}$ with covariates $x^{F}$, the decision-maker provides the sensitive path counterfactual $z^{\mathrm{SPCF}}$ treatments to measure the outcome change.
Remarks. As before, we establish connections to predictive fairness concepts to understand when treatment disparities can result in discrimination through unfair decision outcomes. For specific contexts, we can relate to the perspective of disparate treatment [15] and consider any direct use of $S$ in assigning $Z$ problematic and any consequent effect on $Y$ as discriminatory. In other scenarios, e.g., healthcare, such effects may not be discriminatory if the effect of DTD on $Y$ is beneficial. Similarly, disparities between treatment decisions $z^{\mathrm{SCF}}$ and $z^{F}$ can lead to unfairness in $Y$. For instance, in predictive fairness, separation [69] deems decision disparities justifiable if they do not result in adverse ground-truth outcomes. Similarly, treatment disparities regarding $S$ may be deemed discriminatory if they do not have a positive impact on the ground-truth $Y$, e.g., when assigning $z^{F}$ results in negative outcomes ($y^{F}$), whereas $z^{\mathrm{SCF}}$ would have led to more favorable outcomes ($y^{CF}$).
3.3 Practical implementation
Prior works that computed path-specific counterfactuals for predictive fairness [52, 8] are restricted to linear settings or must rely on imprecise Monte Carlo approximations and complex regularization. To compute path-specific counterfactuals in more practical settings, we propose a novel framework that helps estimate our proposed disparity measures. Assuming access only to factual (potentially unfair) data $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$ and the causal graph $\mathcal{G}$ (Figure 1), our framework builds on the causal normalizing flows (CNF) [26]. CNF has demonstrated accurate causal estimates in societal decision-making settings [41]. Hence, by building our framework based on CNF, we develop a practical yet theoretically grounded approach to compute path-specific counterfactuals.
3.3.1 Background on causal normalizing flows
Causal normalizing flows (CNF) approximate the true and unknown SCM $\mathcal{M}$ using a single, invertible normalizing flows model $T_{\psi}$ with parameters $\psi$. Note that CNF can handle partial causal graphs, in which each random variable in the graph may be multivariate without explicitly modeling the causal relationships across its dimensions. Moreover, given a causal ordering across the random variables, CNF is highly flexible in learning the presence or absence of causal effects from observational data. Hence, even with data collected from a (potentially unfair) decision-making process, the CNF model provides an ideal practical framework to estimate our discrimination measures regarding treatment decisions and their downstream effects. Importantly, a causal normalizing flows model fitted on observational data can be used to estimate the counterfactual datapoint $\vartheta^{CF}$ resulting from a causal intervention $({\operatorname{do}}(V_{j}\rightarrow\alpha))$ on a given factual datapoint $\vartheta^{F}=\{v_{1}^{F},\ldots,v_{K}^{F}\}$ via:
  1. 1.
    abduction step computing the exogenous values for $\vartheta^{F}$ as ${u}^{F}=T_{\psi}(\vartheta^{F})$;
  2. 2.
    action step constructing the exogenous values for $\vartheta^{CF}$ as ${u}^{CF}({\operatorname{do}}(V_{j}\rightarrow\alpha))=\{{u}^{F}_{1:j-1},T_{\psi}({\operatorname{do}}(V_{j}\rightarrow\alpha)),{u}^{F}_{j+1:K}\}$;
  3. 3.
    prediction step estimating the resulting counterfactual feature values from the exogenous variables as $\vartheta^{CF}=T_{\psi}^{-1}\left({u}^{CF}({\operatorname{do}}(V_{j}\rightarrow\alpha))\right)$.
Refer to Appendix A for more details on causal normalizing flows. Next, we show how to adapt the causal normalizing flows for our non-binary decision-making fairness analysis framework.
3.3.2 Approximating counterfactual treatments and outcomes
We can use a causal normalizing flows model trained using factual $\mathcal{D}$ to compute counterfactual quantities for some factual $\vartheta^{F}\in\mathcal{D}$. Importantly, we can readily utilize the existing counterfactual estimation method shown above to compute the sensitive total treatment counterfactual $z^{\mathrm{SCF}}$ as:
$z^{\mathrm{SCF}}=T^{-1}_{\psi}(\langle T_{\psi}({\operatorname{do}}(s\rightarrow(1-s^{F}))),{u}^{F}_{X},{u}^{F}_{Z},{u}^{F}_{Y}\rangle)$ (5)
Similarly, we can directly use the approximation process in Section 3.3.1 to compute the counterfactual labels in Definitions 3 and 4 resulting from an treatment intervention ${\operatorname{do}}(Z\rightarrow\hat{z})$, as:
$y^{CF}({\operatorname{do}}(Z\rightarrow\hat{z}))=T^{-1}_{\psi}(\langle{u}^{F}_{S},{u}^{F}_{X},T_{\psi}({\operatorname{do}}(z\rightarrow\hat{z})),{u}^{F}_{Y}\rangle),$ (6)
where $\hat{z}$ denotes respectively $z^{\mathrm{SCF}}$ from Definition 1, and $z^{\mathrm{SPCF}}$ from Definition 2. However, note that to compute the effect of the counterfactual treatments on the outcome, we need to perform counterfactual approximations in two steps, where the first step measures the counterfactual treatment and the second step measures the effect on the outcome.
Extension for path-specific disparity measures. To analyze our path-specific measures in Definitions 2 and  4, we cannot apply the existing causal normalizing flows counterfactual estimation process directly since it does not work for multiple sequential interventions. Moreover, since we want to compute path-specific measures in non-linear settings, we cannot directly apply the existing linear approach in [8] either. Hence, we provide a novel procedure that extends the existing non-linear causal normalizing flows counterfactual approach to approximate path-specific counterfactuals.
For computing $z^{\mathrm{SPCF}}$ in Definitions 2 and 4, we need to perform sequential interventions in the causal normalizing flows model following the causal ordering of the features to obtain:
$\begin{split}&z^{\mathrm{SPCF}}=T^{-1}_{\psi}(\langle{u}^{CF}_{S},\hat{{u}}_{X}^{F},{u}^{F}_{Z},{u}^{F}_{Y}\rangle),\,\mathrm{where}\\ &{u}^{CF}_{S}=T_{\psi}({\operatorname{do}}(S\rightarrow(1-s^{F}))),\,\mathrm{and}\,\,\hat{{u}}_{X}^{F}=T_{\psi}(s^{CF},{\operatorname{do}}(X\rightarrow x^{F}))\end{split}$ (7)
Hence, to compute $z^{\mathrm{SPCF}}$, we perform two interventions following the causal ordering. First we use ${\operatorname{do}}(S\rightarrow(1-s^{F}))$ on $\vartheta^{F}$ to get $\vartheta^{\mathrm{SCF}}$. Then, we use ${\operatorname{do}}(X\rightarrow x^{F})$ on $\vartheta^{\mathrm{SCF}}$, combine the different exogenous $u$ from the separate intervention steps and generate $z^{\mathrm{SPCF}}$. Further details and pseudocode regarding the different counterfactual computations using CNF can be found in Appendix B.  [^2] Code can be accessed on GitHub.
3.3.3 Assumptions and considerations
While our implementation currently utilizes causal normalizing flows, our causal framework is inherently flexible, allowing for the seamless integration of future advancements in causal generative modeling to generate counterfactual quantities. However, examining the assumptions and certain considerations specific to the current implementation is crucial for understanding its practical implications.
First, we highlight the causal assumptions central to our causal framework. Consistent with previous works in causal modeling [30, 62, 26] and causal fairness [35, 8, 31], we assume the causal data generation process has no hidden confounders. We base our analysis on the decision-maker’s perspective and assume our framework has access to all treatment decisions $Z$ and all covariates $X$ and sensitive $S$ used to assign $Z$. These considerations eliminate confounding between $Z,Y$ and $X,Z$. However, in many real-world scenarios, decision-makers may be unable to account for all $X$ that impact $Y$, leading to unseen confounding amongst $X,Y$. This hidden confounding does not impact total treatment disparity (TTD) but could influence direct treatment disparity (DTD), which relies on interventions on $X$.
Alternatively, future work could consider an alternative external auditor perspective, in which case a lack of transparency on the treatment assignment procedure leads to not accounting for all $Z$ or all $X$ that were used to assign $Z$, thereby introducing confounding across $X,Z,Y$. While some studies [31, 7] have attempted to provide fairness estimation bounds under confounding, they require specific modeling considerations or do not account for multivariate, non-binary treatment decisions. Other works [68, 67, 45] have explored causal inference using instrumental and proxy variables to address hidden confounding. However, these methods only estimate population-level average effects and lack the decision-subject specificity needed to generate counterfactuals in the presence of hidden confounders. Therefore, further research is necessary to understand the impact of confounding on treatment fairness in real-world scenarios and to develop more robust estimation methods for generating counterfactuals in the presence of confounding.
Beyond confounding, another causal assumption central to our framework and CNF is overlap [54]. We assume that intervened values in the SCM are plausible and have non-zero support in the original causal data-generating process. However, this assumption is reasonable and holds in many real-world scenarios. For instance, the treatment $Z$ distributions across different sensitive $S$ for real-world datasets (Appendix C) show significant overlap.
Second, to ensure accurate treatment fairness measures, it is essential to consider the accuracy of the generated counterfactual entities. Assuming underlying causal assumptions hold in a real-world scenario, [26] have already shown how causal normalizing flows can generate highly accurate counterfactuals. However, testing for the validity of the assumptions in real-world settings is challenging. Although we do not explicitly pursue this avenue in this work, our framework and setup offer a unique advantage: it allows decision-makers to assess and improve counterfactual quality through real-world studies. Specifically, our counterfactuals only modify treatments, which are entirely under the control of decision-makers. Hence, decision-makers can adjust treatment policies (e.g., adopting fair counterfactual treatments) and conduct real-world interventional studies on small subpopulations. This possibility enables decision-makers to conduct real-world analyses of counterfactual treatment accuracy, ultimately improving our framework’s applicability.
4 Mitigating treatment discrimination in historical data
In this section, we focus on settings where historical data serves a data-driven decision-making pipeline with one of the following purposes: i) automating the (non-binary) decision-making processes or ii) informing decision-making via risk score estimation. Unfortunately, as discussed in previous sections, historical data may have been collected under discriminatory treatment policies, where disparities in past treatments $Z$ led to undesired harm in the outcome $Y$. In such scenarios, we want to ensure that past discriminatory effects from treatment disparities do not propagate through automated decision-making processes and cause unfair future decisions. To enable future fair decisions, we propose an extension of our framework, providing an automated pre-processing procedure to transform biased historical datasets $\mathcal{D}$ to treatment-fair datasets $\mathcal{D}^{\mathrm{fair}}$. Our pre-processing approach helps prevent the propagation of past treatment discrimination in future automated systems and further improves them through fair risk score estimates.
In the following sections, we assume an observed and historically collected unfair dataset, $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$, comprising a representative sample [^3] We assume the dataset consists of independent and identically distributed samples drawn from an underlying data-generation process. Relaxations of this assumption, such as algorithmic decision-making under selective labels, are discussed in [36]. of the random variables $S$, $X$, $Z$, and $Y$, representing sensitive attributes, covariates, non-binary treatment decisions, and outcomes, respectively.
4.1 Fairness in non-binary decision-making
To automate decision-making processes, decision-makers often use historical data to train a binary classifier, $h_{y}:={S,\tilde{X}}\mapsto Y$, where $S$, $\tilde{X}$, and $Y$ represent, respectively, sensitive attributes, input features, and the target label. To the best of our knowledge, prior works assume that $\tilde{X}=\{X,Z\}$ contains the applicant covariates $X$ (e.g., income, savings, etc.) along with the non-binary treatment decisions $Z$ (e.g., loan terms). In this context, the objective is to achieve both accuracy and fairness by learning a function that produces predictions $\hat{Y}\in\{0,1\}$, which act as proxies for the true target labels $Y\in\{0,1\}$. However, if the decision-maker’s policy to assign treatments is unfair, the resulting automated decision-making process will also be unfair. We address this problem by extending our framework with an automated procedure to transform historically treatment-biased data. This procedure allows the decision-makers to assign fair treatments and ultimately make fair (non-binary) treatment decisions. The idea is that fair treatments should be assigned to lead to better, and thus fairer, outcomes $Y$. Predictive algorithms trained on these fair treatments and outcomes would lead to fairer subsequent decisions.
More in detail, we first identify a fair treatment policy $\pi_{z}:=X\mapsto Z$. However, assigning non-binary treatments is often inherently unfair, complex, or unknown, making the selection of an appropriate $\pi(Z)$ that is optimal and fair for all stakeholders a challenging task. To address this issue, we propose an approach that ensures automated treatment assignments are non-discriminatory while improving decision outcomes for disadvantaged subpopulations. Specifically, when we observe treatment disparities lead to discriminatory effects on outcomes, we leverage our causal framework and the generative capacity of the estimated SCM to augment existing potentially unfair data to construct a fair dataset $\mathcal{D}^{\mathrm{fair}}$. In this automated transformation process, for applicants in the disadvantaged group $S^{F}=s^{-}$ who receive less favorable treatments, we propose intervening in their treatment decisions $Z$ to construct $\mathcal{D}^{\mathrm{fair}}$ as follows:
$\begin{split}\mathcal{D}^{\mathrm{fair}}=\left\{s^{F},x^{F},z^{F},y^{F}\right\}_{s=s^{+}}\bigcup\left\{s^{F},x^{F},z^{\mathrm{SCF}},y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))\right\}_{s=s^{-}}\end{split}$ (8)
where $z^{\mathrm{SCF}}$ follows Definition 1. For applicants in the original (unfair) dataset $\mathcal{D}$ with $S=s^{-}$, we intervene on $Z$ by replacing it with $z^{\mathrm{SCF}}$, derived from the advantaged group. Consequently, the intervened $Z$ and the resulting $Y$ are replaced with their counterfactual values, as specified in Definitions 1 and 3. This intervention on the disadvantaged subpopulation ensures that the revised treatment assignments $z^{\mathrm{SCF}}$ remove treatment discrimination and lead to improved decision outcomes, satisfying $P[Y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))=1\mid S^{F}=s^{-}]\geq P[Y^{F}=1\mid S^{F}=s^{-}]$. This process ensures that disadvantaged applicants receive treatment decisions that eliminate bias while leading to fair decision outcomes. In summary, to automate fair non-binary decision-making processes based on an unfair dataset, one can use our causal framework to generate a fair dataset and then use it to learn a fair treatment policy $\pi_{z}$ and a fair binary classifier $h_{y}$.
4.2 Fair risk score estimation
Risk scores are a critical tool in algorithmic decision-making that decision-makers use to inform decisions and treatment policies [23, 53, 6]. For example, in lending, banks rely on risk score estimates from third-party institutions such as FICO and SCHUFA [14, 64] to understand an applicant’s creditworthiness, helping them inform the loan decisions and justify interest rate variations [65, 38]. Risk scores in decision-making are usually estimated by leveraging historically collected data $\mathcal{D}$ and past decision outcomes [40, 51]. Unfortunately, potential disparities in prior treatment decisions may have a discriminatory influence in past data and outcomes. Hence, when estimating risk scores as the probability of outcome $y^{F}$ using past data covariates $x^{F}$ and past treatments $z^{F}$, the unfairness can propagate and lead to biased risk score estimates $R_{\mathrm{unfair}}$, often modeled as:
$R_{\mathrm{unfair}}(x,z)={P}(y^{F}=0\mid X=x^{F},Z=z^{F}),$ (9)
Importantly, these risk score estimations may remain biased even if we remove treatment decisions from the estimation process due to the causal effect of the treatment $Z$ in the outcome $Y$. Therefore, directly incorporating these risk score estimates as additional applicant features in decision-making will perpetuate or amplify unfairness.
Risk score estimates should be free from potential past treatment discrimination to ensure decision-makers can make well-informed, fair decisions. For instance, in insurance underwriting, fair risk scores allow providers to offer premium rates, mitigate systemic inequities, and prevent perpetuating historical biases in insurance coverage decisions. To ensure fairness, we propose that risk scores designed to estimate an applicant’s inherent risk should exclude historical treatment decisions, as these factors go beyond an applicant’s control and may reflect past biases [33, 12]. Hence, for an applicant, we formulate fair risk score estimates $R_{\mathrm{fair}}$ as:
$\begin{split}R_{\text{fair}}(x,s)=\mathbb{E}_{z^{\prime}\sim\pi(Z)}\left[P\left(y^{CF}(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow z^{\prime}))=0\right)\right],\end{split}$ (10)
where $\pi(Z)$ denotes an interventional distribution over treatments, which we use to marginalize out the impacts of treatments on the risk estimation. These fair risk scores, $R_{\text{fair}}$, can then be integrated into decision-making pipelines with covariates ($X$) to make future fair treatment decisions.
The interventional treatment distribution $\pi(Z)$ in our risk score estimation ensures that $R_{\mathrm{fair}}$ is both fair and treatment-agnostic. Thus, $\pi(Z)$ cannot be any arbitrary distribution; instead, the treatments in $\pi(Z)$ should not systematically disadvantage any sensitive group through adverse downstream effects. As a simple approach, when we observe past treatment disparities are discriminatory, we set $\pi(Z)$ as the interventional treatment distribution of the group $S=s^{+}$ in $\mathcal{D}^{\mathrm{fair}}$. Since the group $S=s^{+}$ historically received favorable treatment assignments and outcomes, our choice of $\pi(Z)$ helps maximize positive outcomes for both groups and mitigates the effects of treatment discrimination. Moreover, since real-world scenarios show overlap across the factual treatments of the intervened groups (Appendix C), our selection of $\pi(Z)$ results in a distribution over realistic treatment values. We leave the challenging exploration of alternative fair interventional distributions that are globally fair policies for future work.
5 Use case : Lending decisions
In this section, we evaluate the effectiveness of our proposed causal framework for measuring and mitigating non-binary treatment discrimination in decision-making processes. We use four real-world datasets across diverse lending scenarios spanning different societal contexts, geographic regions, and loan structures. Given the comprehensive practical foundation for evaluation, our framework shows how historical data can contain treatment disparities and its potentially discriminatory impact on decision outcomes. Additionally, we demonstrate the application of our mitigation strategies to address this treatment discrimination. By focusing on two scenarios, our results explore potential historical biases in treatment decisions, particularly the loan terms set by banks.
  • Agreement of the decision-making process as label (S.1): The ground-truth label $Y$ represents the immediate outcome of the decision-making process, with $Y=1$ indicating a positive agreement on treatment terms between the decision-maker and the applicant. For instance, in lending scenarios, $Y=1$ means the bank approved the loan, and the applicant accepted the terms. We analyze this scenario using the HMDA (Home Mortgage Disclosure Act) dataset, focusing on data from New York and Texas for 2017 [9, 18].
  • Decision outcome as label (S.2): The ground-truth label $Y$ represents the downstream outcome of a decision, indicating whether it was right ($Y=1$) or wrong ($Y=0$). For example, in lending scenarios, $Y=1$ means the loan was fully repaid, while $Y=0$ indicates default. We analyze this scenario using the Home Credit[48] and German Credit [22] datasets as case studies.
First, we use the novel fairness metrics from our framework in Section 3 to quantify treatment disparities in historical lending data. Through our path-specific counterfactual analyses, we demonstrate how sensitive attributes influence treatment decisions both directly and totally and subsequently affect the outcomes, potentially adversely causing discrimination. Next, we mitigate the discrimination using two strategies: (i) generating treatment-fair counterfactual datasets to automate unbiased decision-making and (ii) adjusting risk score estimation to reduce the impact of unfair treatments. We detail the experimental setup and datasets in Appendix C and provide additional results on synthetic datasets in Appendix D to analyze the accuracy of our causal estimation-based framework.
5.1 Measuring treatment disparities
5.1.1 S.1: Agreement of the decision-making process as label.
Table 2: Measuring treatment disparities in S.1 and outcome effects. $Y^{F}=0$ indicates a label change from negative to positive agreement, $Y^{F}=1$ the reverse. Results show median U.S. Dollars (\$currency-dollar\\$\$).
[…]
Amount (×$\times 10^{3}$\$) $Y^{F}=0$ $Y^{F}=1$
HMDA-NY Female $\rightarrow$ Male TTD(-E) $+29.46$ $2.00$ $0.40$
DTD(-E) $+6.65$ $2.00$ $0.20$
Male $\rightarrow$ Female TTD(-E) $-29.93$ $4.00$ $0.16$
DTD(-E) $-5.88$ $2.00$ $0.14$
HMDA-TX Female $\rightarrow$ Male TTD(-E) $+19.43$ $2.00$ $0.28$
DTD(-E) $+2.67$ $2.00$ $0.13$
Male $\rightarrow$ Female TTD(-E) $-19.78$ $3.00$ $0.18$
DTD(-E) $-3.65$ $3.00$ $0.10$
In both HMDA datasets, we follow the identification and pre-processing procedures for sensitive attributes, covariates, and treatment decisions outlined in [10]. We consider “gender” a sensitive attribute, while loan amount and preapproval status are treatment decisions within the bank’s decision-making authority. We summarize the results of the treatment discrimination analyses in Table 2.
Results in HMDA. We analyze the treatment disparity measures ($\operatorname{TTD}$ and $\operatorname{DTD}$) for factual males and females across the two states. Our analyses show that female applicants, if treated like their male counterfactuals, would receive a substantial increase in loan amounts, e.g., in NY, approximately \$currency-dollar\\$\$29K more based on $\operatorname{TTD}$ (and \$currency-dollar\\$\$6K based on $\operatorname{DTD}$) on average. Conversely, if male applicants were treated like females, their loan amounts would decrease by a similar amount. Examining the effects on the outcome agreement rates, we observe similar $\operatorname{TTD-E}$ and $\operatorname{DTD-E}$ values for both factual females and males, with agreement rates increasing slightly more when males are treated as females. The HMDA-TX dataset shows similar, though less pronounced, treatment disparities compared to the NY dataset.
Takeaway. Although we find no significant differences in loan agreement rates between genders, banks tend to adopt a more conservative approach in determining treatments for female applicants, suggesting potential treatment discrimination. Our findings show that offering higher loan amounts to female applicants would not negatively impact their agreement outcomes, casting doubt on the rationale behind such differential treatment. Furthermore, the observed disparities seem primarily mediated through applicant covariates, with minimal direct disparity. The lack of direct disparities and minimal impact on the immediate outcome (agreement rates) raises important questions about the root causes of the differential treatments. Specifically, one might ask: Is the disparity in loan terms against female applicants justified by the downstream outcome, such as repayment behavior, and thus justified by business necessity? This question forms the basis for our analyses in the second scenario, where we investigate the link between treatment disparities, repayment outcomes, and discrimination.
Table 3: Measuring treatment disparities in S.2 and outcome effects. $Y^{F}=0$ indicates a label change from loan defaulted to repaid, $Y^{F}=1$ the reverse. Results show median INR (Indian Rupees) and DM (Deutsche Marks).
[…]
Annuity (INR) Amount (INR) $Y^{F}=0$ $Y^{F}=1$
Home Credit Female $\rightarrow$ Male TTD(-E) $+1008.12$ $+7885.72$ $0.24$ $0.30$
DTD(-E) $+715.91$ $-816.94$ $0.27$ $0.00$
Male $\rightarrow$ Female TTD(-E) $-942.92$ $-6659.68$ $2.30$ $0.00$
DTD(-E) $-594.56$ $+758.46$ $1.30$ $0.00$
Duration (months) Amount (DM) $Y^{F}=0$ $Y^{F}=1$
German Credit Female $\rightarrow$ Male TTD(-E) $+1.29$ $+253.34$ $2.00$ $3.50$
DTD(-E) $+0.46$ $+164.66$ $2.00$ $3.00$
Male $\rightarrow$ Female TTD(-E) $-0.79$ $-243.11$ $8.00$ $1.00$
DTD(-E) $-0.59$ $-212.39$ $8.00$ $0.60$
5.1.2 S.2: Decision outcome as label.
In the Home Credit dataset, we analyze two treatment decisions: the loan amount and annuity amount, while in the German Credit dataset, we examine three treatment decisions: loan amount, loan duration, and installment rate. In both datasets, we use “gender” as the sensitive attribute to evaluate potential disparities in treatment decisions. We summarize our treatment disparities results in Table 3.
Results in Home Credit. Our analyses show that if females were treated like their male counterfactuals, they would have received higher annuities and loan amounts. At the same time, the reverse pattern would apply to male applicants when treated as females. Interestingly, we see contrasting trends between total and direct disparities regarding loan amounts. Total disparity indicates conservative lending practices where female applicants are granted lower loan amounts and annuities. In contrast, direct disparity, attributed solely to gender, shows that female applicants receive lower loan amounts but higher annuities, suggesting elevated interest rates. These findings suggest that while banks may have implemented some affirmative treatment policies, disparities mediated through covariates influence lending decisions significantly. Importantly, we see slight downstream effects on repayment outcomes across groups, indicating that outcome differences do not justify the observed treatment disparities against female applicants, particularly in terms of $\operatorname{DTD}$. Hence, in this context, the unjustified treatment disparities may be considered discriminatory.
Results in German Credit. We observe a similar trend with the bank adopting more conservative practices toward female applicants. If females were treated like males, they would receive longer loan durations and higher loan amounts, with the reverse pattern observed for males. Significant disparity is seen in loan amounts, driven mainly by direct effects. This observation suggests that applicant covariates do not fully explain treatment disparities, and gender may have a direct influence, exposing a potential problematic mechanism. Interestingly, there is no disparity in installment rate treatments regarding $\operatorname{TTD}$ or $\operatorname{DTD}$, so we exclude this feature from Table 3 for clarity. When examining the impact of treatment disparities on outcomes, a key observation arises: treating males like their female counterparts would result in 8.0% of defaulting males repaying their loans. In contrast, only 1.0% of male repayers would default under this counterfactual treatment. Hence, our counterfactual analysis shows that the treatment disparities are not justified by outcomes, raising concern on potential treatment discrimination.
Takeaway. Our results in Scenario S.2 show that, across different datasets, female applicants experience negative treatment disparities that cannot be explained by downstream outcome differences. Furthermore, while males receive perceived positive treatment disparity, they are ultimately negatively impacted and discriminated by its downstream effects on outcomes. Notably, our findings align with similar insights observed for German Credit in a simpler linear setting [28], reinforcing the validity of our results. Despite appearing favorable, the loan terms offered to certain males may increase their default risks and lower creditworthiness over time. Applying the treatment decisions of their female counterparts to male applicants could mitigate this discrimination and improve repayment outcomes. Next, we further explore the mitigation of this discrimination in treatment decisions.
5.2 Mitigating treatment discrimination
Our results in Table 3 show that banks exhibit negative treatment disparity against female applicants. However, males are potentially discriminated against since they experience a negative downstream effect on repayment ability, and treating them as females improves repayment performance. Hence, treating all applicants conservatively as females may enhance the overall repayment ability of borrowers, benefiting both the banks and the applicants.
5.2.1 Fair non-binary decision-making
Following Section 4.1, we examine the creation of a treatment-fair dataset aimed at automating future fair non-binary decision-making tasks. Based on our Scenario S.2 analyses, we identify males as the group needing adjusted treatment to generate our fair dataset, $\mathcal{D}^{\mathrm{fair}}$. To understand the impact of $\mathcal{D}^{\mathrm{fair}}$, we aim to compare the fair dataset with the original (unfair) dataset on two key metrics: the bank’s Loss Given Default ($\operatorname{LGD}$[63] and the applicant’s Expected Simple Interest ($\operatorname{ESI}$[59]. Detailed information on the metric calculations and dataset-specific implementations can be found in Appendix C.2.
Takeaway. The results for Home Credit and German Credit, summarized in Table 4, show that adjusting treatment decisions for male applicants addresses treatment biases and yields tangible financial benefits for both stakeholders. The adjusted treatments result in an improved $\operatorname{LGD}$ for banks, indicating reduced financial risk. Simultaneously, male borrowers can reduce loan costs owing to the lower resultant $\operatorname{ESI}$. These findings underscore the potentially broad impact of fairer treatment practices to improve the conditions for all stakeholders.
5.2.2 Fair risk score estimation
Following Section 4.2, we now examine how we can more fairly inform decision-making processes by estimating treatment-fair risk scores. Based on insights from scenario S.2, we focus on male applicants and apply different treatment distributions $\pi(Z)$ to generate fair risk estimates as defined in Equation 10. We approximate $\pi(Z)$ for males under two conditions: (i) a fair interventional distribution, derived from the empirical distribution of $Z$ conditioned on female applicants to produce “fair” risk scores, and (ii) the factual “unfair” distribution, based on the empirical distribution of $Z$ conditioned on male applicants. For female applicants, $\pi(Z)$ is approximated using the empirical distribution of $Z$ conditioned on female applicants. We apply these policies to the Home Credit and German Credit datasets and compare the empirical cumulative distribution functions (CDFs) of the resulting risk scores for male and female applicants in Figure 2.
Takeaway. The figure shows that, while males generally exhibit lower risk scores than females for German Credit, the pattern reverses for Home Credit. However, historical treatment practices have led to an unfair overestimation of males’ risk scores, especially in German Credit. This bias is corrected when the risk scores are recalculated under the fair interventional distribution, adjusting for potentially unfair treatments. Integrating these recalibrated, fair risk scores into decision-making pipelines mitigates the residual impacts of historical treatment biases, paving the way for more fair decision-making processes.
Table 4: Fairness in non-binary decision-making in S.2. Comparison of bank’s LGD and applicant’s ESI (both: lower better) under different policies. $\mathcal{D}$ is factual data, $\mathcal{D}^{\mathrm{fair}}$ data is obtained by treating all applicants as females.
Dataset Sensitive Group LGD (INR/DM) ESI (INR/DM)
Home Credit Females $586692.67$ $290208.73$
Males ($\mathcal{D}$) $576590.35$ $422114.31$
Males ($\mathcal{D}^{\mathrm{fair}}$) 568157.43 391110.73
German Credit Females $1220.51$ $312.13$
Males ($\mathcal{D}$) $1163.88$ $495.65$
Males ($\mathcal{D}^{\mathrm{fair}}$) 940.03 393.86
(a) Home Credit
(b) German Credit
Figure 2: Fair risk score estimation in S.2. Comparison of risk score estimates for females and males under various treatment distributions $\pi(Z)$. Unfair estimates are based on factual distribution, while fair estimates use interventional treatment distribution where all applicants are treated as females.
6 Related work
Fairness in lending. Fairness analyses in automated lending decisions have focused mainly on statistical methods [21, 24, 37], emphasizing metrics like equality of opportunity [21], as noted in surveys [34, 19]. However, these approaches often overlook biases embedded in the ground-truth labels, which stem from unfair treatment during loan processing. While prior work has separately examined treatment discrimination [60, 1] and biased outcomes in lending decisions [50, 3, 47], the causal link between biased treatment and its impact on outcomes remains underexplored. Our novel framework bridges this gap and jointly analyzes treatment discrimination and its causal effects on lending outcomes.
Causal fairness notions in algorithmic predictive systems. Causal fairness in algorithmic decision-making examines the causal influence of sensitive features on decision labels and outcomes. Counterfactual fairness [35] introduced the study of the total causal effect of sensitive attributes, while subsequent work explored effects along specific pathways, distinguishing fair from unfair pathways [52, 8]. Later frameworks expanded to include multi-level attributes [43] and focused on fairness in decision outcomes [57]. However, these approaches primarily address biases in binary labels or outcomes, neglecting complex mechanisms like treatment decisions (e.g., loan terms) controlled by decision-makers. Although [39] introduced a causal framework for treatment fairness, it simplified multi-faceted treatments into binary variables, limiting its granularity. Our work advances this by incorporating the decision-maker’s perspective to analyze treatment decisions and their causal effects. We develop a novel framework for fine-grained fairness analyses using path-specific counterfactuals [8]. Additionally, following [41], we enhance counterfactual accuracy in societal contexts through state-of-the-art causal normalizing flows [26], surpassing previous variational methods [8, 39].
Multi-stage selection in decision-making. Multi-stage selection problems [16, 27] explore automating decision-making with binary decisions across stages, revealing additional candidate covariates for subsequent decisions. Concerning this line of work, our causal framework addresses a fundamentally different context governed by different mechanisms. Our work focuses on a single-stage decision process where binary decisions (e.g., loan approval) are made in one step, with treatments directly controlled by the decision-maker. In contrast, multi-stage problems rely on observing additional applicant covariates. We complement existing studies by analyzing fairness in automated decision-making, specifically the fairness of assigned multivariate, non-binary treatments and their causal impact on decision outcomes.
7 Discussion
This work highlights the importance of incorporating non-binary treatment decisions in fairness analyses for automating decision-making systems. We proposed a novel causal framework to measure disparities in treatment decisions across sensitive groups and its impact on downstream outcomes. We also demonstrate how our framework can help mitigate unfairness and promote fairer decisions. However, ensuring its practical and reliable implementation requires addressing and discussing several critical challenges.
Ensuring robust causal frameworks for real-world scenarios. Our analyses assume access to ground-truth labels for outcomes and the absence of hidden confounders in the causal generative process. However, these assumptions often fail in real-world settings and specific use cases (e.g., decision-maker vs auditor as user). Hidden confounders can bias treatment fairness estimates while using proxy labels for unobservable outcomes [66], which introduces additional complexities. Addressing these challenges by accounting for hidden confounders is a promising direction for future research.
Designing mitigation strategies for diverse stakeholders. Similar to pre-processing approaches in predictive fairness, our pre-processing mitigation methods for treatment discrimination prove empirically effective but they do not guarantee optimality. Moreover, these approaches may not fully align with the diverse and often conflicting objectives of all stakeholders, raising a crucial question: What defines a fair treatment policy, and how should it be formulated? For instance, banks prioritize repayment probability in loan approval, while applicants seek lower interest rates [55]. Addressing fairness in such settings requires a holistic approach that considers all stakeholders’ utilities and relies on carefully crafted utility functions to capture these objectives reliably. Hence, future work should compare our methods to existing simple treatment interventions [12] while also focusing on novel in-processing methods to develop context-aware policies that can balance fairness and diverse stakeholder utilities.
Improving data collection to enable treatment discrimination analysis. As shown in Table 1, while our framework focuses on loan datasets, it is broadly applicable across decision-making domains. However, its practical application is constrained by significant data limitations. Many datasets fail to distinguish covariates $X$ from treatments $Z$, omit critical treatment details (e.g., COMPAS [5] lacks bail terms, lending datasets exclude loan interest rates), or oversimplify multifaceted treatments (e.g., IHDP [39] reduces childcare factors to binary). Moreover, most datasets are collected under selective labeling settings [36, 32], where ground-truth outcomes are observed only for cases with a favorable agreement between the decision-maker and the decision-subject. This setting is problematic, as it overlooks unknown agreement-phase discrimination that may have occurred during the agreement process. These gaps emphasize the need for more transparent and detailed data collection practices that explicitly capture treatment decisions, including those made during the agreement phase, to enable comprehensive and fair analyses.
Concluding remarks.
Our work introduces a paradigm shift by integrating fairness considerations into non-binary treatments conditional on positive binary decisions (e.g., granting a loan). By proposing treatment disparity measures, we complement existing literature to advance a more holistic analysis of fairness. Our novel perspectives and empirical analyses demonstrate that fairness in algorithmic decision-making remains unresolved, necessitating interdisciplinary socio-technical approaches to address the needs of diverse stakeholders. By accounting for the complexity of decisions and their impact on stakeholders’ utility, our work signifies a major step toward aligning algorithmic systems with societal expectations. Our approach also paves the way for future research to adapt established predictive fairness methods to redefine and mitigate treatment unfairness.
Acknowledgements
We thank Adrián Javaloy for providing invaluable guidance regarding incorporating the causal generative models and feedback on our written manuscript. This work has been funded by the European Union (ERC-2021-STG, SAML, 101040177). However, the views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible.
Author contributions
Conceptualization: IV, DK; Formalization: All authors; Methodology: AM, IV; Related work: AM, DK; Data curation: DK, KG; Code implementation: AM, KG; Experimental evaluation: All authors; Writing - Original draft: All authors with AM as main contributor to Sections 2, 3, 6, 7, Appendix B), DK to Sections 56,Appendices CD; Writing - Review and editing: All authors.; Supervision: IV.
Appendix A Causal normalizing flows
Causal normalizing flows [26] enable causal inference from observational data under minor assumptions like invertible and differentiable structural equations, acyclic (known) graphs $\mathcal{G}$, and causal sufficiency. Under these assumptions, causal flows represent a SCM $\mathcal{M}$ as Triangular Monotonic Increasing (TMI) maps and model them using single, invertible normalizing flows models $T_{\psi}$ with parameters $\psi$. Note that causal flows can handle partial causal graphs, in which each random variable in the graph may be multivariate without explicitly modelling the causal relationships across its dimensions. Thus, it provides an ideal practical framework to estimate our measures of fairness with respect to treatment decisions and the downstream effects from available (potentially unfair) data collected from a decision-making process.
A.1 Hyper-Parameter tuning and training of causal normalising flows
For the causal normalizing flows [26], we use the original work as inspiration and tune for the flow models’ inner dimensionality and learning rate Table 5. All other parameters are fixed, which are as follows: Adam optimizer with plateau scheduler, activation as ELU, layer type was Neural Spline Flow (NSF) and all the experiments were run for a maximum of $1000$ epochs. For other parameters, we refer to the GitHub code  [^4] https://github.com/psanch21/causal-flows. The number of nodes is specified in Table 6. An example of causal graph input to causal normalizing flows is given in Figure 3.
Table 5: Hyperparameters for causal Flows. The log-prob of the observational data distribution estimate was used to select the best hyperparameters.
Dataset dim-inner lr log-prob
HMDA-NY $32\times 32$ $10^{-2}$61.67
HMDA-TX $32\times 32$ $10^{-2}$63.49
Home Credit $64$ $10^{-2}$41.87
German Credit $64$ $10^{-3}$14.68
Synthetic Loan $16\times 16\times 16\times 16$ $10^{-3}$16.51
Table 6: Number of root nodes, covariates, and treatment variables for training causal normalizing flows model. The dimensionality of the outcome label is 1.
Dataset #rootnodes #covariates #treatment
HMDA-NY 2 12 2
HMDA-TX 2 12 2
Home Credit 2 9 2
German Credit 2 15 3
Synthetic Loan 2 3 2
A.2 Hardware
All experiments of hyper-parameter tuning and training of causal normalizing flows are done on an Ubuntu workstation. The code primarily uses Pytorch CPU and Python. We set the seed for PyTorch and Python to a fixed number and ran all experiments. A fixed seed is also used to split datasets into train, validation, and test sets - 80%, 10%, 10%. This seed is fed as input in a parameter file to train and run evaluations.
AG
[…]
hammer
EISLDYSensitive $S$Outcome $Y$Block $X$Block $Z$
Figure 3: Causal graph for synthetic loan dataset introduced in Appendix D. While training the causal normalising flows, covariates $X$ and treatment decisions $Z$ are treated as multivariate blocks without any explicit assumptions on the cause-effect directionality within each block. Here, we have multiple sensitive attributes (independent root nodes) but for this work, we concentrate on intervening on just one sensitive attribute (“gender”). A similar analysis can be done by intervening on the attribute “age”.
Appendix B Counterfactual computation pseudocodes
This section provides the different function pseudocode for computing the various counterfactual and path-specific counterfactual entities using the causal normalizing flows. It is important to note and ensure that whenever we are setting a categorical feature value for intervention, the value is clamped to be within range and floored.
The algorithm names provided below match the function names used in the source code for easy understanding. Moreover, note that for our default consideration of binary sensitive attributes, $s^{CF}$ can be considered as $(1-s^{F})$ following the notation in the main paper.
B.1 Generating sensitive counterfactual
This function computes the sensitive counterfactual following the causal relationships when ${\operatorname{do}}(S\rightarrow s^{CF})$. For this computation, we perform the single ${\operatorname{do}}$ operation on $S$ as ${\operatorname{do}}(S\rightarrow s^{CF})$.
Algorithm 1 compute_sensitive_counterfactual
Input: Factual data $D^{F}$
Parameters: Flows model, data preparator, sensitive index $i_{S}$ (default: 0)
Output: Sensitive counterfactual data $D^{SCF}({\operatorname{do}}(S\rightarrow s^{CF}))$
1:Get sensitive feature values $s^{F}$ from $D^{F}$.
2:Get sensitive counterfactual feature value $s^{SCF}$ for intervention.
3:Get data scaler from preparator.
4:$D^{SCF}$leftarrow model.compute_CF ($D^{F}$, $i_{S}$, $s^{SCF}$, scaler)
5:return $D^{SCF}$. $\triangleright$ clamp and floor categorical features before saving.
B.2 Computing downstream effect of counterfactual treatments
This function helps to compute the downstream effect on $Y$ if we perform ${\operatorname{do}}(Z\rightarrow\hat{z})$, where $\hat{z}$ is some interventional value. For example, we use this function to :
  • understand the downstream effect of Total Treatment Discrimination (TTD) and
  • understand the downstream effect of Direct Treatment Discrimination (DTD).
Algorithm 2 compute_path_specific_downstream_
effect_treatment_intervention
Input: Factual data $D^{F}$, counterfactual data $\hat{D}$
Parameters: Flows model, data preparator
Output: $D(s^{F},x^{F},{\operatorname{do}}({Z\rightarrow\hat{z}}))$
1:Get treatment indices $i_{\{T\}}$ from preparator
2:Get values $\hat{z}$ from $\hat{D}$ using $i_{\{T\}}$
3:Get scaler from preparator
4:Clamp and floor $\hat{z}$ categorical features to proper value ranges
5:$D^{F}({\operatorname{do}}(Z\rightarrow\hat{z}))$ = model.compute_CF($D^{F}$, $i_{\{T\}}$, $\hat{z}$, scaler)
6:return $D(S^{F},X^{F},{\operatorname{do}}(Z\rightarrow\hat{z}))$triangleright Extract $Y(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow\hat{z}))$ and clamp, floor.
B.3 Computing direct effect of sensitive on the target
This function computes the direct effect $S$ on $Y$, i.e. ${\operatorname{do}}(S\rightarrow s^{CF})$ on the target $Y$ by computing $y({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$.
Algorithm 3 compute_path_specific_sensitive
_direct_label
Input: Factual data $D^{F}$, sensitive counterfactual data $D^{SCF}$ from Algorithm 1
Parameters: Flows model, data preparator
Output: $D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$
1:Get scaler from preparator
2:$U^{F}$ = model.flow.transform($D^{SCF}$) $\triangleright$ Helps get $U^{F}_{Y}$
3:$U^{SCF}$ = model.flow.transform ($D^{SCF}$) $\triangleright$ Helps get $U^{SCF}_{S}$
4:Get intervention indices $i_{\{X,T\}}$
5:Clamp, floor categorical dimensions in $D^{F}[i_{\{X,T\}}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $D^{F}$
6:Set data_tmp = clone($D^{SCF}$)
7:Set data_tmp$[i_{\{X,T\}}]$ = intervention_vals$[i_{\{X,T\}}]$
8:$U^{\mathcal{I}}$ = model.flow.transform (data_tmp) $\triangleright$ Helps get $U^{\mathcal{I}}_{\{X,Z\}}$
9:Get combined $U_{\text{comb}}=[U^{SCF}_{S},U^{\mathcal{I}}_{\{X,Z\}},U^{F}_{Y}]$
10:$D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F},Z\rightarrow z^{F}))$ = model.flow.transform.inverse($U_{\text{comb}}$)
11:return $D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$triangleright Extract $y({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$ and clamp, floor.
B.4 Computing path-specific treatment
With this function we calculate the path-specific counterfactuals i.e. we compute the path-specific treatment $z^{PS}=z({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$.
Algorithm 4 compute_path_specific_treatment
Input: Factual data $D^{F}$, sensitive counterfactual data $D^{SCF}$ from Algorithm 1
Parameters: Flows model, data preparator
Output: $D^{PS}=D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$
1:Get scaler from preparator
2:$U^{F}$ = model.flow.transform($D^{SCF}$) $\triangleright$ Helps get $U^{F}_{Y},U^{F}_{Z}$
3:$U^{SCF}$ = model.flow.transform ($D^{SCF}$) $\triangleright$ Helps get $U^{SCF}_{S}$
4:Get intervention indices $i_{X}$
5:Clamp, floor categorical dimensions in $D^{F}[i_{X}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $D^{F}$
6:Set data_tmp = clone($D^{SCF}$)
7:Set data_tmp[$i_{X}$] = intervention_vals[$i_{X}$]
8:$U^{\mathcal{I}}$ = model.flow.transform(data_tmp) $\triangleright$ Helps get $U^{\mathcal{I}}_{X}$
9:Get combined $U_{\text{comb}}=[U^{SCF}_{S},U^{\mathcal{I}}_{X},U^{F}_{Z},U^{F}_{Y}]$
10:$D^{PS}$ = model.flow.transform.inverse($U_{\text{comb}}$)
11:return $D^{PS}$triangleright Extract $z^{PS}$, clamp and floor categorical features before saving.
Computing downstream effect of $z^{PS}$
Note that once we have $z^{PS}$ from Algorithm 4, we can compute the downstream effect, i.e., $Y(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow z^{PS}))$ using Algorithm 2.
Appendix C Datasets, preprocessing and other details
C.1 Datasets and their preprocessing
  1. 1.
    HMDA: This dataset regarding mortgage applications covers several states in the United States, but we focused on just two states (New York and Texas) in the year 2017. Following similar preprocessing steps as done in [10], we binarized the gender column as $\{Male:1,Female:0\}$. Also, we binarized the race with originally $8$ possible values but utilized only $\{NativeWhite:0,White:1\}$. For the action taken with $8$ possible values, we filtered out ${3-8}$. We kept the outcome as a $1$ Loan originated to mean the applicant and the bank decision agreed, and $2$, which is the application approved but not accepted, to mean a disagreement. In dealing with missing values we removed columns with all NaNs and information regarding the co-applicants. After the preprocessing, we had a total dataset of size $399354$ with $17$ features.
  2. 2.
    Home Credit: We downloaded this dataset from the kaggle competition. We used the application training datasets and selected features for our study based on the feature importance as done in [42]. We utilized the resulting features selected by their random forest classifier and information entropy models. Based on this, we selected the following features : “EXT SOURCE 2, EXT SOURCE 3, EXT SOURCE 1, DAYS BIRTH, NAME EDUCATION TYPE, CODE GENDER, DAYS EMPLOYED, NAME INCOME TYPE, ORGANIZATION TYPE, AMT CREDIT, AMT ANNUITY, AMT GOODS PRICE, REGION POPULATION RELATIVE, TARGET". Based on this subset of data, we dropped the missing values in the dataset and converted the days of birth and days of employment to years using days of $\frac{\text{days of birth/employment}}{-365}$. We divided it by $-365$ because these values were negative. The complete dataset has a size of $98859$ with $14$ features.
  3. 3.
    German Credit: We obtained this dataset from the Statlog website and restructured the personal status feature by consolidating the information on gender and marital status into binary gender categories. Specifically, we grouped male single and male divorced/separated as male, and female divorced/separated and female single as female. The dataset had no missing values.
(a) German Credit
(b) Home Credit
Figure 4: Treatment decision feature distributions across social groups show overlap, indicating it as a reasonable causal assumption for real-world settings.
Remarks. We converted all categorical features using label encoding. In Table 7, we summarize our datasets by identifying treatment features, sensitive attributes, and outcomes. We note that the treatment decision features for all the lending datasets in our work are continuous features. Hence, we use the difference between treatment values as $\mathcal{F}$ and the expectation $\mathcal{A}$ over each dataset $\mathcal{D}$ when considering disparity measures in Definitions 1 and 2. Furthermore, from Figure 4, we note that treatment feature distributions show significant overlap across the two sensitive groups. This observation shows that our overlap assumption is a reasonable one and holds in real-world settings.
Table 7: Different loan approval datasets we used in the study with their sensitive attributes ($S$), treatment decisions ($Z$), outcome label ($Y$) and dataset size.
Dataset
Sensitive
Attribute (S)
Treatment
Decisions (Z)
Outcomes (Y) Dataset Size
HMDA-TX
Gender, Race
and Ethnicity
Preapproval,
Loan amount
Loan accepted or
rejected by the applicant
($399354,17$)
HMDA-NY
Gender, Race
and Ethnicity
Preapproval,
Loan amount
Loan accepted or
rejected by the applicant
($166637,17$)
Home Credit
Gender,
Days of birth (Age)
Annuity,
Credit Amount
Loan repayment ($98859,14$)
German Credit Gender, Age
Duration,
Credit Amount,
Installment rate
Loan repayment
($1000,21$)
C.2 Losses
  1. 1.
    Loss Given Default (LGD): We estimate the loss incurred by the bank in the event an applicant defaults. Quantifying the loss will require additional analysis of other variables such as the value of collateral, instalments payments etc [25]. There are a number of different ways to estimate $LGD$ however a common approach considers the exposure at default and the recovery rate that is:
[…]
    In our estimation, we assume that EAD is equivalent to the credit amount loaned out. EAD is a crucial metric that estimates the total loss a bank is exposed to when a loan defaults. However, it is important to note that the calculation of EAD can be more complex, as it may involve factors beyond the simple loan amount, such as commitment details [58].
    The recovery rate, on the other hand, is a risk-adjusted measure that reflects the portion of the loan that the bank expects to recover after default. In our analysis, we considered it as the outcome of the loan (i.e., the actual repayment).
    Finally, we formalize the expected loss given default for the bank as:
    LGD $=\mathbb{E}_{\mathcal{D}^{\prime}}[(1-y)\times\text{amount}]$ (12)
    where $y$ represents the applicant’s repayment probability.
  2. 2.
    Expected Simple Interest (ESI):
    We estimate the loss of the individuals using the simple interest formula:
    $\textbf{SI}=\frac{(\text{amount}\times\text{rate}\times\text{time})}{100}$
    For the Home Credit dataset, we used slightly modified formulae to estimate the ESI for Home Credit, as the dataset did not include information on the rate and time, making ESI estimation less straightforward. We estimate ESI using annuity and loan amount as follows: we assume duration=15 years. Annuity can be written as :
    annuity=(amount + interest)total durations(months)=(amount + $\text{annuity}=\frac{(\text{amount + interest})}{\text{total durations(months)}}=\frac{(\text{amount + interest})}{12\times\text{duration}}$
    Using this the simple interest can be calculated as :
    ESI(Home $\textbf{ESI}\text{(Home Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}[\text{annuity}\times 12\times\text{duration}-\text{amount}]$ (13)
    In our formulae for the German Credit dataset, the loan duration is in months and we assume in our implementation that everyone was given the same interest rate of 10%, hence enabling us to rewrite the ESI as follows:
    ESI(German $\textbf{ESI}\text{(German Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}\left[y\times\text{amount}\times\frac{\text{rate}}{100}\times\frac{\text{time}}{12}\right]$ (14)
Appendix D Synthetic loan
D.1 Synthetic dataset generation
Taking inspiration from [29], we consider the 7-variable loan dataset modeling the credit dataset [22] with an added outcome variable $Y$. Similar to [29], the equations are non-linear with non-additive exogenous variables. The nodes, gender $G$ and age $A$ are root-nodes (non-actionable), loan amount $L$, education-level $E$, income $I$, and savings $S$ are actionable with equations shown below. Additionally, we added a link between the savings $S$ to the loan amount and thresholded the outcome variable $Y$ directly on $G$ using two different thresholding scenarios. The symbol $\sigma(\cdot)$ represents the sigmoid function in the following equations. Distribution N (a,b) denotes the Gaussian distribution where a is the mean and b is the variance of the distribution, $\sqrt{b}$ denoting the standard deviation. Finally, Gamma (a, b) denotes the Gamma distribution where a is the concentration/shape and b is the scale, 1/b being the rate.
Root nodes : (15)
$f_{G}:G=U_{G},\quad U_{G}~{}\sim\operatorname{Bernoulli}(0.5)$
$f_{A}:A=-35+U_{A},\quad U_{A}~{}\sim\operatorname{Gamma}(10,3.5)$
Covariates:
$f_{E}:E=-0.5+\sigma(-1+0.5G+\sigma(0.1A)+U_{E}),$
$\quad\quad U_{E}\sim\mathcal{N}(0,0.25)$
$f_{I}:I=-4+0.1(A+35)+2G+GE+U_{I},U_{I}~{}\sim~{}\mathcal{N}(0,4)$
$f_{S}:S=-4+1.5\mathbb{I}\{I>0\}I+U_{S},U_{S}~{}\sim~{}\mathcal{N}(0,5)$
Treatments:
$f_{L}:L=1+0.01(A-5)(5-A)+2(1-G)+\beta S+U_{L},$
$\quad\quad U_{L}~{}\sim~{}\mathcal{N}(0,10),\beta\in\{0,0.03\}$
$f_{D}:D=-1+0.1A+3(1-G)+L+U_{D},U_{D}~{}\sim~{}\mathcal{N}(0,9)$
Outcome: two possibilities
$f_{Y}:Y=\mathbb{I}\{\sigma(\delta(-L-D)+0.3(I+S+\alpha IS))$
$\geq 0.5+\gamma(1-G)\}\quad\textbf{OR}$
$f_{Y}:Y=\mathbb{I}\{\sigma(\delta(-L-D)+0.3(I+S+\alpha IS)$
$+U_{Y}*\gamma(1-G))\geq 0.5\},U_{Y}~{}\sim~{}\mathcal{N}(0,5)$
[…]
$\alpha=1,\textbf{if}\left[\mathbb{I}\{I>0\}\land\mathbb{I}\{S>0\})\right]=1,$
$\alpha=-1,\textbf{otherwise}$
[…]
Table 8: Measuring treatment disparities and discrimination for synthetic data. Oracle uses perfect causal knowledge. Here, we have loan amount and duration as the treatment features. Results are reported as median values.
Dataset Sensitive counterfactual Measure Treatment Disparity Effect on Repayment (%)
Flows Oracle Flows Oracle
Amount Duration Amount Duration $Y^{F}=0$ $Y^{F}=1$ $Y^{F}=0$ $Y^{F}=1$
Balanced Female $\rightarrow$ Male TTD(-E) $-1.91$ $-4.44$ $-1.95$ $-4.94$ $0.19$ $0.00$ $0.18$ $0.00$
DTD(-E) $-1.62$ $-4.25$ $-2.00$ $-5.00$ $0.17$ $0.00$ $0.18$ $0.00$
Male $\rightarrow$ Female TTD(-E) $+2.14$ $+4.64$ $+1.95$ $+4.95$ $0.00$ $0.20$ $0.00$ $0.20$
DTD(-E) $+1.90$ $+4.05$ $+2.00$ $+5.00$ $0.00$ $0.17$ $0.00$ $0.20$
Imbalanced Female $\rightarrow$ Male TTD(-E) $-1.92$ $-4.61$ $-1.95$ $-4.95$ $0.09$ $0.00$ $0.09$ $0.00$
DTD(-E) $-1.55$ $-3.75$ $-2.00$ $-5.00$ $0.07$ $0.00$ $0.09$ $0.0$
Male $\rightarrow$ Female TTD(-E) $+2.00$ $+4.69$ $+1.95$ $+4.95$ $0.00$ $0.21$ $0.00$ $0.21$
DTD(-E) $+2.21$ $+5.45$ $+2.00$ $+5.00$ $0.00$ $0.24$ $0.00$ $0.21$
Table 9: Fairness in non-binary decision-making. Comparison of bank’s LGD and applicant’s ESI losses for balanced and unbalanced datasets under different policies.
Sensitive Group Balanced Dataset Unbalanced Dataset
LGD (DM) ESI (DM) LGD (DM) ESI (DM)
Males $3.09$ $0.58$ $3.30$ $0.58$
Females ($\mathcal{D}$) $4.14$ $0.36$ $3.49$ $0.30$
Females ($\mathcal{D}^{\mathrm{fair}}$) 3.14 $0.55$ 2.39 $0.47$
D.2 Results
Using the synthetic data generation process emulating the lending scenario as shown above, we generate balanced and unbalanced datasets and repeat our empirical analysis to i) use our novel notions to analyze the data for treatment discrimination, and ii) explore our mitigation strategies to enable treatment-fair decisions. Additionally, with access to the ground-truth generative process, we can consider an oracle model with the ground-truth knowledge. Using the oracle, we can perform interventions to compute the oracle treatment discrimination values. This allows us to compare the results given by our CNF-based causal framework with the oracle.
D.2.1 Balanced dataset:
Using the synthetic loan generation process described in the previous section, we set the parameters $\beta=0.03,\gamma=0.5,\delta=1,\eta=5$ to construct a balanced dataset containing $5000$ samples. The dataset was balanced concerning both gender and the outcome variable $Y$. We evaluated the performance of the causal normalizing flows (CNF) model by comparing its results to the oracle model that leverages perfect causal knowledge. The comparison, detailed in Table 8, demonstrates that the CNF model effectively captures the underlying structure of the observational data. This corroborates that CNF is a robust and reliable model for learning from observational data while accounting for causal relationships.
Results: From Table 8, we observe that females receive longer loan durations and larger loan amounts compared to their male counterparts. This significant discrimination is reflected in both $\operatorname{TTD}$ and $\operatorname{DTD}$. Remarkably, treating females as males would lead to a $0.19\%$ of defaulting females to repay, while no female applicant would have defaulted under this treatment. This suggests that a positive treatment discrimination exists towards females.
To address this discrimination, it might be beneficial to apply the treatment of males to their female counterparts. This adjustment aims to mitigate treatment discrimination and improve repayment outcomes. Using this insight, we generate the fair dataset $\mathcal{D^{\text{fair}}}$ as shown in Section 4.1. We further evaluate this treatment intervention by calculating the bank’s loss ($\operatorname{LGD}$) and applicant loss ($\operatorname{ESI}$) with results summarized in Table 9. Interestingly, while applying male treatment to females reduces the bank’s losses, it increases the applicants’ losses, highlighting that such an intervention benefits one stakeholder (the bank) but not the female applicants.
Figure 5(a) illustrates the empirical cumulative distribution function (CDF) of risk scores under different treatment distributions $\pi(Z)$: the factual distribution and the interventional treatment distribution conditioned on male applicants. This analysis reveals that males typically exhibit lower risk scores than females, with a substantial gap between the two. Applying the suggested fairer treatment reduces this gap, making the risk scores fairer across genders by mitigating the treatment discrimination.
Takeaway: While our risk score estimation strategy shows improvement in mitigating treatment discrimination, our fairness in the decision-making strategy falls short of improving conditions for all stakeholders in the decision-making setup by merely adjusting the treatment decisions. This limitation likely arises because our current approach lacks guarantees on the optimality of the designed utility functions. To address this shortcoming, we need to develop more balanced and fair treatment policies $\pi_{z}$ that not only mitigate discrimination but also optimize outcomes for all stakeholders involved. Designing such policies will enable a more fair and efficient decision-making pipeline.
D.2.2 Unbalanced dataset
To create the imbalanced dataset, we adjusted the parameters to $\beta=0.03$, $\gamma=0.5$, $\delta=2$, and $\eta=5$, generating a dataset with $5000$ samples. This configuration resulted in a dataset with a higher proportion of males compared to females and a greater number of individuals who defaulted on their loans than those who repaid them. These reflect real-world disparities often encountered in financial datasets, providing a more challenging scenario for evaluating the model’s performance.
Results: Referring to Table 8, we observe a similar trend to the balanced dataset, females receive longer loan durations and higher credit amounts compared to males. Likewise, there is significant discrimination in the impact of both $\operatorname{TTD}$ and $\operatorname{DTD}$ on repayment, just like in the balanced case. When male treatment decisions are applied to females, the risk scores of females improve, as shown in Figure 5(b) and the gap in the risk score decreases. Additionally, Table 9 highlights a consistent pattern: treating females as males reduces bank losses but increases applicant losses, following the same trend observed in the balanced dataset scenario.
Takeaway: The unbalanced dataset reveals a similar trend, highlighting the limitations of our current treatment policies $\pi_{z}$, which may be sub-optimal in addressing fairness and stakeholder utilities. The results highlight the necessity of developing treatment policies that are not only fair but also tailored to optimize utility across diverse groups in the dataset. For the unbalanced dataset, fairness treatment interventions must account for the unequal distribution of data points across groups since the imbalance can make discrimination worse and impact the stakeholder utilities. Designing fair and optimal policies $\pi_{z}$ for such scenarios requires integrating strategies that address both the imbalance in representation and the need for fair treatment decisions.
(a) Balanced
(b) Unbalanced
Figure 5: Fair risk score estimation. Comparison of risk score estimates for females and males under various treatment distributions $\pi(Z)$. Unfair estimates are based on the factual distribution, while fair estimates use interventional treatment distribution where all applicants are treated as males.

### v2

@@
 Abstract
-Fairness studies of algorithmic decision-making systems often simplify complex decision processes, such as bail or loan approvals, into binary classification tasks.
-However, these approaches overlook that such decisions are not inherently binary (e.g., approve or not approve bail or loan); they also involve non-binary treatment decisions (e.g., bail conditions or loan terms) that can influence the downstream outcomes (e.g., loan repayment or reoffending).
-In this paper, we argue that non-binary treatment decisions are integral to the decision process and controlled by decision-makers and, therefore, should be central to fairness analyses in algorithmic decision-making.
-We propose a causal framework that extends fairness analyses and explicitly distinguishes between decision-subjects’ covariates and the treatment decisions.
-This specification allows decision-makers to use our framework to (i) measure treatment disparity and its downstream effects in historical data and, using counterfactual reasoning, (ii) mitigate the impact of past unfair treatment decisions when automating decision-making.
+Fairness studies of algorithmic decision-making systems often simplify complex decision processes, such as bail or lending decisions, into binary classification tasks (e.g., approve or not approve).
+However, these approaches overlook that such decisions are not inherently binary; they also involve non-binary treatment decisions (e.g., loan or bail terms) that can influence the downstream outcomes (e.g., loan repayment or reoffending).
+We argue that treatment decisions are integral to the decision-making process and, therefore, should be central to fairness analyses.
+Consequently, we propose a causal framework that extends and complements existing fairness notions by explicitly distinguishing between decision-subjects’ covariates and the treatment decisions.
+Our framework leverages path-specific counterfactual reasoning to: (i) measure treatment disparity and its downstream effects in historical data; and (ii) mitigate the impact of past unfair treatment decisions when automating decision-making.
 We use our framework to empirically analyze four widely used loan approval datasets to reveal potential disparity in non-binary treatment decisions and their discriminatory impact on outcomes, highlighting the need to incorporate treatment decisions in fairness assessments.
-Moreover, by intervening in treatment decisions, we show that our framework effectively mitigates treatment discrimination from historical data to ensure fair risk score estimation and (non-binary) decision-making processes that benefit all stakeholders.
-**footnotetext: These authors contributed equally to this work
-Keywords Algorithmic Fairness  $\cdot$ Causality  $\cdot$ Counterfactual  $\cdot$ Decision-making  $\cdot$ Discrimination  $\cdot$ Path-specific
+Finally, by intervening in treatment decisions, we show that our framework effectively mitigates treatment discrimination from historical loan approval data to ensure fair risk score estimation and (non-binary) decision-making processes that benefit all stakeholders.
+Codehttps://github.com/ayanmaj92/fair-nonbin-treat
+Extended versionhttps://arxiv.org/abs/2503.22454
 1 Introduction
-Table 1: Examples of data collected across various decision-making domains, highlighting the distinction between sensitive attributes $S$, decision-subject covariates $X$, treatment decisions $Z$, and outcomes $Y$.
-The multivariate, non-binary $Z$ have been represented previously in the literature as a positive binary decision (in gray).
+Data-driven systems are increasingly used to automate decisions in domains such as finance, healthcare, and criminal justice (Almheiri 2023; Moscato, Picariello, and Sperlí 2021; Habehh and Gohel 2021; Dieterich, Mendoza, and Brennan 2016).
+These systems typically reduce complex decisions to binary classification tasks, for example, simply predicting for loan repayment or recidivism (Mhasawade, D’Amour, and Pfohl 2024).
+However, real-world decisions are rarely as simple as a binary choice, they involve multiple, non-binary treatment decisions which in turn affect outcomes.
+Fig.
+1 illustrates the data-driven decision-making pipeline, which involves both (a) treatment decisions, (b) binary decisions, and (c) their downstream effects.
+Figure 1: Illustration of data-driven decision-making pipelines. Parts (a)–(c) highlight the three common simplifications in existing fairness analyses.
+Despite growing research on fairness in algorithmic decision-making, existing analyses simplify complex decision-making mechanisms in three key ways.
+First, many studies (Zafar et al.
+2017; Hardt, Price, and Srebro 2016; Corbett-Davies et al.
+2023) include non-binary treatments as covariates of the decision-subject (see covariate and treatment examples in Appendix Table 4), restricting fairness analyses to binary decisions (Fig.
+1b).
+This obscures the fact that such decisions are in control of the decision-makers.
+As a result, potential sources of discrimination may be masked, limiting the scope of fairness assessments.
+Second, some works (Madras et al.
+2019; Coston et al.
+2020; Plecko and Bareinboim 2023) account for treatment decisions (Fig.
+1a) but assume they are binary, ignoring their often non-binary nature (e.g., loan or bail amounts).
+Finally, while a few studies do focus on disparities in non-binary treatment decisions, e.g., in lending (Agier and Szafarz 2013; Escalante et al.
+2018; Alesina, Lotti, and Mistrulli 2013), they only provide empirical insights into disparities in treatment but overlook the downstream effects these treatment disparities may have on outcomes (Fig.
+1c).
+In this paper, we bridge the gaps in the fairness literature by extending fairness analyses to multiple, non-binary treatment decisions and their downstream outcome effects.
+We propose a causal framework that explicitly distinguishes between a decision-subject’s covariates and the treatment decisions made by a decision-maker.
+Our framework enables us to: (i) measure disparities in non-binary treatment decisions induced by the sensitive attributes, and (ii) quantify how these disparities propagate to outcomes, thereby inducing treatment discrimination.
+Furthermore, we discuss how the proposed measures of treatment disparities and their downstream effects relate and complement existing fairness notions in binary decision-making.
+To make our framework actionable, we introduce a practical approach for generating fairness-aware counterfactual data, leveraging recent advances in causal inference (Javaloy, Sánchez-Martín, and Valera 2024).
+Unlike prior methods, which often relied on restrictive assumptions or computationally intensive procedures (Nabi and Shpitser 2018; Chiappa 2019), our approach is broadly applicable and scalable.
+Moreover, it supports treatment discrimination mitigation via pre-processing techniques.
+These techniques generate treatment-fair counterfactual datasets, enabling fairer non-binary automated decisions and more equitable risk score estimation to inform future decision-making.
+We validate our framework using four real-world lending datasets across two decision-making scenarios and empirically show that: (i) historical data encode treatment disparities that fair binary predictions fail to correct for owing to downstream effects of treatment decisions; and (ii) fairer treatment decisions yield more optimal outcomes, benefiting all stakeholders.
+Our findings thus highlight the need to address treatment decision disparities to achieve fairer outcomes, bridging a key gap in fairness research, and laying the foundation for more holistic fairness assessments and mitigation efforts.
+Thus, our analysis opens up avenues for future work, such as developing fair data-driven approaches for jointly making treatment and binary decisions.
+2 Background
+This section introduces the key causal concepts we leverage in our framework for studying treatment decision disparities.
+Structural causal models (SCMs).
+An SCM (Pearl et al.
+2000) encodes the cause-effect relations among observed features $V$ as $\mathcal{M}=(V,U,F)$, where $U$ is a set of mutually independent, unobserved exogenous variables, and $F$ is a set of functions such that each feature $V_{i}\in V$ is given by $V_{i}=F_{i}(\operatorname{pa}(V_{i}),U_{i})$, with $\operatorname{pa}(V_{i})\subseteq V$ denoting the causal parents of $V_{i}$.
+The structure of these equations induces a directed acyclic graph (DAG) $\mathcal{G}$ over $V$, as seen in Fig.
+2.
+Performing interventions.
+Interventions are external actions that forcibly fix a feature’s value, breaking its usual dependence on causal parents (Pearl et al.
+2000).
+In SCMs, such hypothetical interventions are denoted by the $\operatorname{do}$}-operator.
+For instance, the intervention $\mathcal{I}={\operatorname{do}}(V_{i}=\alpha)$ replaces $V_{i}$’s structural equation with the constant assignment $V_{i}=\alpha$, producing a modified model $\mathcal{M}^{\mathcal{I}}$.
+This intervened SCM captures the changed causal structure and enables reasoning about the intervention’s downstream effects.
+XXZZYYSS
+Figure 2: Causal graph of the data generation process, distinguishing decision-subject covariates $X$ and treatment decisions $Z$, assuming positive binary decisions.
+Red paths (solid: direct, dotted: indirect) show potential disparate influence of sensitive $S$ on $Z$.
+The orange dashed path from $Z$ to $Y$ reflects possible discriminatory effects on outcomes.
+Counterfactuals.
+In the absence of hidden confounders, the intervened SCM $\mathcal{M}^{\mathcal{I}}$ enables counterfactual analysis for a specific individual with observed features $\vartheta^{F}$.
+This analysis answers the question: “What would the counterfactual features $\vartheta^{CF}$ be if we intervened with ${\operatorname{do}}(V_{i}\rightarrow\alpha)$, all else remaining the same?” Computing counterfactuals involves three steps (Pearl et al.
+2000): (i) abduction that infers the exogenous $u^{F}$ from $\vartheta^{F}$ using the original $\mathcal{M}$; (ii) action that intervenes using ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ to produce the intervened SCM $\mathcal{M}^{\mathcal{I}}$; and (iii) prediction that computes the counterfactual features $\vartheta^{CF}$ from $\mathcal{M}^{\mathcal{I}}$ using $u^{F}$ and $V_{i}=\alpha$.
+Counterfactuals capture the total effect of an intervention on downstream causal features for the individual.
+This framework is widely used in fairness research (Kusner et al.
+2017) to quantify the total effect of sensitive attributes (e.g., gender) on binary decisions (e.g., granting loans).
+Path-specific counterfactuals.
+In certain scenarios, instead of focusing on total effects, we may want to understand how one feature influences another through specific causal pathways.
+This can be done using path-specific counterfactuals, which apply multiple $\operatorname{do}$} operations to isolate effects along selected paths.
+Consider a causal structure involving variables $\langle V_{i},V_{j},V_{k}\rangle$ with two paths: a direct path $V_{i}\rightarrow V_{k}$ and an indirect path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$.
+For a factual instance $\vartheta^{F}=\langle v_{i}^{F},v_{j}^{F},v_{k}^{F}\rangle$, we can compute the counterfactual effect of ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ along each path separately.
+To isolate the direct path $V_{i}\rightarrow V_{k}$, we block the indirect path by fixing $V_{j}$ to its factual value, yielding $v_{k}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha),{\operatorname{do}}(V_{j}\rightarrow v_{j}^{F}))$.
+To isolate the indirect path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$, we fix $V_{i}=v_{i}^{F}$ for the direct path, but update $V_{j}$ based on the intervention using $v_{k}^{CF}(v_{i}^{F},{\operatorname{do}}(V_{j}\rightarrow v_{j}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))))$.
+This approach has been effective in fairness studies (Chiappa 2019) to determine whether sensitive features influence binary decisions through potentially problematic causal paths.
+3 Measuring Treatment Disparities
+In this section, we introduce our causal framework to measure non-binary treatment disparities and their downstream effects.
+Our framework considers historical data $\mathcal{D}$ generated by the causal graph $\mathcal{G}$ (Fig.
+2), where:
+  1. 1.
+    Treatment decisions $Z$ are assigned by decision-makers (e.g., banks setting loan terms) and are causally influenced by decision-subject covariates $X$ (e.g., income, savings) and possibly the sensitive attribute $S$ (e.g., gender, race).
+This mechanism may exhibit disparities via direct (solid red) and indirect (dashed red) effects of $S$.
+  2. 2.
+    Treatment decisions $Z$ causally affect outcomes $Y$ (e.g., loan terms affect repayments).
+Thus, $S$ may influence $Y$ indirectly through $Z$ (orange dashed line in Fig.
+2), potentially propagating treatment disparities to outcomes.
+Remarks.
+Although Fig.
+2 presents a single $S$ for simplicity, our framework supports modeling multiple as independent root nodes.
+We treat $X$ and $Z$ as multivariate blocks, focusing on causal relations between blocks without making assumptions about causal structures within them.
+The directed edges in $\mathcal{G}$ indicate the causal ordering (Javaloy, Sánchez-Martín, and Valera 2024) among feature groups, although some of these relations may have zero effect.
+To enable general analyses, we retain $S\rightarrow Y$ to account for direct effects in scenarios like healthcare.
+3.1 Disparities in Non-Binary Treatment Decisions
+Based on the causal graph in Fig.
+2, we use the SCM framework to measure how $S$ influences treatment decisions $Z$, capturing both the total and the direct disparities.
+For simplicity and without loss of generality, we assume a binary sensitive attribute, $S\in\{0,1\}$.
+[^1] Discussions on possible extensions to multi-valued and multiple sensitive features in Appendix B.3.
+Given a dataset $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$ consisting of instances that received a positive decision under a historical binary predictive policy, we define disparity measures using a disparity function $\Delta$, which depends on the type of treatment variable, e.g., using value differences for continuous treatments.
+Definition 1.
+Total Treatment Disparity (TTD) against decision-subjects with factual sensitive feature $s^{F}$, measures the expected disparity between the treatment $z^{F}$ they received and the treatment $z^{\mathrm{SCF}}$ they would have received if they belonged to a different sensitive group, i.e.,
+$\begin{split}\operatorname{TTD}(s^{F})&=\mathbb{E}_{\mathcal{D}}\left(\Delta(z^{\mathrm{SCF}},z^{F})\right),\,\mathrm{where}\\ z^{\mathrm{SCF}}&=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})))\end{split}$ (1)
+Following Kusner et al.
+(2017), this notion captures the total effect of $S$ on the treatment decisions $Z$.
+The sensitive counterfactual treatment $z^{\mathrm{SCF}}$ reflects, e.g., the loan terms a female applicant would receive if they were male with counterfactually adjusted covariates like income.
+Definition 2.
+Direct Treatment Disparity (DTD), against decision-subjects with factual sensitive feature $s^{F}$, measures the expected disparity between the treatment they received and the sensitive direct-path counterfactual treatment, i.e.,
+$\begin{split}\operatorname{DTD}(s^{F})&=\mathbb{E}_{\mathcal{D}}\left(\Delta(z^{\mathrm{SPCF}},z^{F})\right),\,\mathrm{where}\\ z^{\mathrm{SPCF}}&=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})),{\operatorname{do}}(X\rightarrow x^{F}))\end{split}$ (2)
+Following Nabi and Shpitser (2018); Chiappa (2019), this captures the direct effect of $S$ on $Z$.
+The sensitive direct-path counterfactual $z^{\mathrm{SPCF}}$ reflects, e.g., the loan terms a female applicant would have received if they identified as male while maintaining their original covariates like income.
+Relation to fair binary predictions.
+Ensuring fairness in binary decisions determines who receives a positive decision (e.g., loan granted) but does not directly address disparities in more complex, non-binary treatment assignments.
+We illustrate this empirically in Sec.
+5.
+Still, drawing parallels with predictive fairness concepts can help interpret the implications of treatment disparities.
+For example, $\operatorname{DTD}$ captures disparities caused by the explicit use of sensitive attributes $S$ in assigning $Z$.
+While this may be acceptable in settings like healthcare, it is generally inappropriate in domains like lending, echoing the notion of disparate treatment (Dwork et al.
+2012).
+Similarly, zero $\operatorname{TTD}$ requires $Z$ to be independent of $S$, analogous to demographic parity (Dwork et al.
+2012), which requires binary predictions to be independent of $S$.
+As with predictive fairness, the appropriateness of treatment parity depends on context.
+When the true outcome $Y$ is unobserved (e.g., hiring), treatment disparities may be considered discriminatory.
+In contrast, when $Y$ is observable and the ground-truth (e.g., loan repayment), separation-based fairness frameworks (Hardt, Price, and Srebro 2016; Zafar et al.
+2017) suggest that outcome differences may justify some disparities.
+Hence, analyzing the downstream impact on $Y$ is essential for assessing the unfairness of treatment disparities.
+3.2 Downstream Effects of Treatment Disparities
+While Def.
+1 and 2 help measure treatment disparities, it is crucial to assess the downstream impact of these disparities on the outcome $Y$ (e.g., when $Y$ is a ground truth as in lending) to determine if they are discriminatory.
+Next, we introduce two novel metrics to quantify the downstream outcome effects of total and direct treatment disparities.
+Definition 3.
+Total Treatment Disparity Effect (TTD-E) measures the downstream effect of the sensitive counterfactual treatment $z^{\mathrm{SCF}}$, computed in Eq.
+1, as the probability of a label change from the factual value $y^{F}$, i.e.,
+$\begin{split}\operatorname{TTD-E}(s^{F},y^{F})&=\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}\left(y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))\neq y^{F}\right)\right],\end{split}$ (3)
+where $\mathbb{I}$ is an indicator function.
+This measures how the outcome would change if the decision-maker applied the sensitive counterfactual treatment $z^{\mathrm{SCF}}$, e.g., how a female applicant’s repayment would differ if her covariates stayed the same but treatment was adjusted by the total effect of $S$.
+Definition 4.
+Direct Treatment Disparity Effect (DTD-E) measures the downstream effect of the sensitive path counterfactual treatment $z^{\mathrm{SPCF}}$, computed in Eq.
+2, as the probability of a label change from the factual value $y^{F}$, i.e.,
+$\operatorname{DTD-E}(s^{F},y^{F})=\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}\left(y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SPCF}}))\neq y^{F}\right)\right].$ (4)
+This measures how the outcome would change if the decision-maker applied the sensitive direct-path counterfactual treatment $z^{\mathrm{SPCF}}$, e.g., how a female applicant’s repayment would differ if her covariates stayed the same but only the treatment was adjusted by the direct effect of $S$.
+Relation to fair binary predictions.
+Ensuring parity in treatment decisions complements parity in binary decisions, and fairness in predictions alone cannot eliminate unfair effects of treatment disparities on outcomes.
+However, connecting to predictive fairness concepts helps identify when treatment disparities lead to discrimination through unfair decision outcomes.
+For example, under the notion of disparate treatment (Dwork et al.
+2012), any direct use of sensitive attribute $S$ in assigning treatment $Z$ is problematic, and its impact on outcome $Y$ is considered discriminatory.
+However, in some contexts like healthcare, DTD effects may be justified if they benefit $Y$.
+Similarly, disparities between counterfactual and factual treatments ($Z^{\mathrm{SCF}}$ and $Z^{F}$) can cause unfairness in $Y$.
+In predictive fairness, separation (Zafar et al.
+2017) allows decision disparities if they do not harm ground-truth outcomes.
+Likewise, treatment disparities (e.g., TTD) may be deemed discriminatory if they fail to improve $Y$, such as when factual treatment $z^{F}$ leads to worse outcomes than counterfactual treatment $z^{\mathrm{SCF}}$.
+3.3 Practical Implementation
+Prior work (Nabi and Shpitser 2018; Chiappa 2019) on path-specific counterfactuals for predictive fairness was either limited to linear settings or relied on imprecise approximations and regularization.
+To allow for a practical yet theoretically grounded method for computing path-specific counterfactuals, we leverage causal normalizing flows (CNF) (Javaloy, Sánchez-Martín, and Valera 2024), which have shown accurate causal estimates in societal decision-making settings (Majumdar and Valera 2024).
+Background on CNF.
+Given factual data $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$ and causal graph $\mathcal{G}$, CNF approximates the unknown SCM $\mathcal{M}$ using a single invertible normalizing flow model $T_{\psi}$.
+CNFs support partial causal graphs, where some features are grouped into multivariate blocks, requiring only a causal ordering across blocks.
+CNFs can also flexibly learn from data if certain causal relations have zero effect.
+Once trained, CNF estimates counterfactuals for a given factual instance $\vartheta^{F}=\{v_{1}^{F},\ldots,v_{K}^{F}\}$ under an intervention ${\operatorname{do}}(V_{j}\rightarrow\alpha)$ through three steps: (i) Abduction; infer exogenous variables as $u^{F}=T_{\psi}(\vartheta^{F})$; (ii) Action, modify $u^{F}$ to reflect the intervention, yielding $u^{CF}=\{u^{F}_{1:j-1},T_{\psi}({\operatorname{do}}(V_{j}\rightarrow\alpha)),u^{F}_{j+1:K}\}$; and, (iii) Prediction, estimate counterfactual $\vartheta^{CF}=T_{\psi}^{-1}(u^{CF})$.
+See Appendix B.2 for extended details.
+Next, we show how to adapt CNF to estimate treatment disparities.
+Estimating counterfactual treatments and outcomes.
+We can use a causal normalizing flows model trained using factual $\mathcal{D}$ to compute counterfactual quantities for some factual $\vartheta^{F}\in\mathcal{D}$.
+Importantly, we can readily utilize the existing counterfactual estimation method shown above to compute the sensitive total treatment counterfactual $z^{\mathrm{SCF}}$ as:
+$z^{\mathrm{SCF}}=T^{-1}_{\psi}(\langle T_{\psi}({\operatorname{do}}(s\rightarrow(1-s^{F}))),{u}^{F}_{X},{u}^{F}_{Z},{u}^{F}_{Y}\rangle)$ (5)
+[…]
+$y^{CF}({\operatorname{do}}(Z\rightarrow\hat{z}))=T^{-1}_{\psi}(\langle{u}^{F}_{S},{u}^{F}_{X},T_{\psi}({\operatorname{do}}(z\rightarrow\hat{z})),{u}^{F}_{Y}\rangle),$ (6)
+where $\hat{z}$ denotes respectively $z^{\mathrm{SCF}}$ from Def.
+1, and $z^{\mathrm{SPCF}}$ from Def.
+2.
+However, note that to compute the effect of the counterfactual treatments on the outcome, we need to perform counterfactual approximations in two steps, where the first step measures the counterfactual treatment and the second step measures the effect on the outcome.
+Extension for path-specific disparity measures.
+To analyze our path-specific measures in Def.
+2 and 4, we cannot directly apply the existing CNF counterfactual estimation process since it does not work for multiple sequential interventions.
+Moreover, since we want to compute path-specific measures in non-linear settings, we cannot directly apply the existing linear approach in (Chiappa 2019) either.
+Hence, we provide a novel procedure that extends the existing non-linear causal normalizing flows counterfactual approach to approximate path-specific counterfactuals.
+For computing $z^{\mathrm{SPCF}}$ in Def. 2 and 4, we need to perform sequential interventions in the CNF following the causal ordering of the features to obtain:
+$\begin{split}&z^{\mathrm{SPCF}}=T^{-1}_{\psi}(\langle{u}^{CF}_{S},\hat{{u}}_{X}^{F},{u}^{F}_{Z},{u}^{F}_{Y}\rangle),\,\mathrm{where}\\ &{u}^{CF}_{S}=T_{\psi}({\operatorname{do}}(S\rightarrow(1-s^{F}))),\,\mathrm{and}\,\,\\ &\hat{{u}}_{X}^{F}=T_{\psi}((1-s^{F}),{\operatorname{do}}(X\rightarrow x^{F}))\end{split}$ (7)
+Hence, to compute $z^{\mathrm{SPCF}}$, we perform two interventions following the causal ordering.
+First we use ${\operatorname{do}}(S\rightarrow(1-s^{F}))$ on $\vartheta^{F}$ to get $\vartheta^{\mathrm{SCF}}$.
+Then, we use ${\operatorname{do}}(X\rightarrow x^{F})$ on $\vartheta^{\mathrm{SCF}}$, combine the different exogenous $u$ from the separate intervention steps and generate $z^{\mathrm{SPCF}}$.
+Further details and pseudocode regarding the different counterfactual computations using CNF can be found in Appendix B.6.
+Assumptions and considerations.
+Our causal framework is designed to incorporate advances in causal generative modeling flexibly, but its current CNF-based implementation relies on key assumptions.
+Our framework assumes overlap, i.e., all relevant treatment-covariate combinations have non-zero support, a condition that generally holds in practice (Appendix B.4).
+Consistent with prior work (Khemakhem et al.
+2021; Sánchez-Martin, Rateike, and Valera 2022; Kusner et al.
+2017; Chiappa 2019), we also assume no hidden confounders and full observability of treatments, covariates, and sensitive attributes from the decision-maker’s perspective.
+However, in real-world settings, decision-makers may not observe all covariates impacting the outcome, introducing potential confounding.
+However, this would only impact the direct disparity measures (Definitions 2 and 4) since they need interventions on $X$.
+Importantly, since decision-makers only intervene on treatment decisions in our framework, robustness can be practically tested through small-scale interventional studies on subpopulations.
+Future work can also explore integrating recent advancements for tackling confounding (Almodóvar et al.
+2025).
+We provide a more detailed discussion in Appendix B.4.
+Appendix D presents synthetic data experiments validating the framework’s effectiveness against an oracle.
+4 Mitigating Historical Treatment Unfairness
+Our framework identifies when historical data reflects discriminatory treatment policies, i.e., disparities in treatments $Z$ that are not justified by corresponding benefits in ground-truth outcomes $Y$.
+To prevent such biases from being perpetuated in data-driven decision-making, we propose an automated pre-processing procedure.
+Although pre-processing approaches may lack formal guarantees, they remain a practical and effective mitigation strategy (Mutlu, Yousefi, and Ozmen Garibay 2022).
+Our method augments the biased dataset $\mathcal{D}$ to produce a fairer version, $\mathcal{D}^{\mathrm{fair}}$, enabling more equitable non-binary decisions (Sec.
+4.1).
+We further show that using $\mathcal{D}^{\mathrm{fair}}$ can mitigate the effects of treatment disparities on risk score estimation, a central component in many decision-making pipelines (Sec.
+4.2).
+4.1 Fairness in Non-Binary Decision-Making
+Automated decision-making systems often rely on historical data, including sensitive attributes, individual covariates, and past treatments to train classifiers that predict outcomes: $h_{y}:\{S,X,Z\}\mapsto Y$.
+These predictions are used to make binary decisions, such as granting loans (Corbett-Davies et al.
+2017).
+However, biases in the decision-maker’s non-binary past treatment policy (Def.
+14) can propagate into future automated decisions.
+As we show later in Sec.
+5 and Appendix E, optimizing for predictive fairness (Dwork et al.
+2012; Hardt, Price, and Srebro 2016) does not eliminate treatment disparities or their downstream effects.
+To prevent automated systems from reinforcing such disparities, we propose a pre-processing method that uses our causal framework to transform biased historical data $\mathcal{D}$ into a treatment-fair dataset $\mathcal{D}^{\mathrm{fair}}$.
+We define a treatment policy $\pi(Z)$ that adjusts treatment assignments for disadvantaged individuals $S^{F}=s^{-}$, those who previously received worse treatment without benefiting in outcome (Def.
+3).
+For these cases, $\pi(Z)$ assigns the sensitive counterfactual treatment $z^{\mathrm{SCF}}$ (Def.
+1).
+The fair dataset is then constructed as:
+$\begin{split}\mathcal{D}^{\mathrm{fair}}&=\left\{s^{F},x^{F},z^{F},y^{F}\right\}_{s=s^{+}}\bigcup\\ &\left\{s^{F},x^{F},z^{\mathrm{SCF}},y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))\right\}_{s=s^{-}}\end{split}$ (8)
+For individuals in group $S^{F}=s^{-}$, we replace the historically assigned treatment with $z^{\mathrm{SCF}}$ and simulate the counterfactual outcome under this intervention (Def.
+1 and 3).
+Despite its simplicity, $\pi(Z)$ corrects disparities without harming outcomes, ensuring $P\left[Y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))=1\mid S^{F}=s^{-}\right]\geq P\left[Y^{F}=1\mid S^{F}=s^{-}\right].$ The resulting $\mathcal{D}^{\mathrm{fair}}$ removes past disparate treatment and its effects, enabling the training of predictive models that do not perpetuate such disparities.
+4.2 Fair Risk Score Estimation
+Risk scores are widely used in algorithmic decision-making to inform lending decisions (Hurley and Adebayo 2016; Obermeyer et al.
+2019).
+For example, banks often rely on risk scores from third-party institutions such as FICO or SCHUFA to assess creditworthiness and justify loan terms (Doroghazi 2020; Toh 2023; Loqbox 2024).
+These scores are typically estimated from historical data $\mathcal{D}$ and past decision outcomes (Maiden 2024; MyFico 2018) as $R(X,Z)=P(Y=0\mid X,Z).$ However, disparities in prior non-binary treatment decisions can bias both outcomes and risk scores.
+Crucially, this bias may persist even if treatment decisions are excluded from the estimation, due to the causal effect of treatment $Z$ on outcome $Y$.
+Since the treatment decisions are beyond an applicant’s control (Kleinberg, Mullainathan, and Raghavan 2016; Coston et al.
+2020), we argue for mitigating the unfair effects of past treatment policies on risk score estimation.
+To this end, we formulate a fair risk score estimate as $R_{\text{fair}}(x,s)=\mathbb{E}_{z^{\prime}\sim\pi(Z)}\left[P\left(y^{CF}(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow z^{\prime}))=0\right)\right],$ where we use $\pi(Z)$ to marginalize out disparate effects of $Z$.
+To ensure the marginalization does not systematically disadvantage any group, we set $\pi(Z)=P_{\mathcal{D}^{\mathrm{fair}}}(Z\mid S=s^{+})$, using treatments from the advantaged group.
+Treatment distribution overlap across groups (Appendix B.4) ensures our interventions remain realistic while removing disparate effects.
+Data
+SCF
+Measure Disparity Effect($\%$)
+Amount (K\$) $Y^{F}=0/1$
+Hist. EOD Hist. EOD
+[…]
+DTD(-E) +7+7 +7+7 $2.5/0.2$ $3.6/0.1$
+[…]
+DTD(-E) 6-6 7-7 $1.8/0.1$ $2.6/0.0$
+[…]
+DTD(-E) +3+3 +3+3 $1.6/0.1$ $0.0/0.0$
+[…]
+DTD(-E) 4-4 3-3 $2.4/0.1$ $0.0/0.0$
+[…]
+Data
+SCF
+Measure Disparity Effect ($\%$)
+Annuity (INR) Amount (INR) $Y^{F}=0$ $Y^{F}=1$
+Hist. EOD Hist. EOD Hist. EOD Hist. EOD
+Home Credit F $\rightarrow$ M TTD(-E) +1005.26+1005.26 +1031.86+1031.86 +7834.47+7834.47 +8606.90+8606.90 $0.22$ $0.18$ $0.25$ $0.13$
+DTD(-E) +708.24+708.24 +694.17+694.17 797.40-797.40 972.88-972.88 $0.22$ $0.00$ $0.08$ $0.05$
+M $\rightarrow$ F TTD(-E) 957.37-957.37 924.41-924.41 6778.88-6778.88 7310.10-7310.10 $2.46$ $3.08$ $0.05$ $0.04$
+DTD(-E) 604.94-604.94 529.54-529.54 +857.12+857.12 +1069.00+1069.00 $1.36$ $1.54$ $0.08$ $0.06$
+Duration (months) Amount (DM) $Y^{F}=0$ $Y^{F}=1$
+German Credit F $\rightarrow$ M TTD(-E) +1.15+1.15 $0.94$ +272.15+272.15 +189.89+189.89 $4.25$ $0.00$ $2.70$ $3.64$
+DTD(-E) +0.42+0.42 +0.36+0.36 +133.39+133.39 +116.74+116.74 $4.25$ $0.00$ $1.35$ $1.82$
+M $\rightarrow$ F TTD(-E) 0.87-0.87 0.67-0.67 237.97-237.97 212.44-212.44 $12.33$ $9.10$ $0.97$ $0.81$
+DTD(-E) 0.58-0.58 0.42-0.42 189.68-189.68 189.68-189.68 $12.33$ $4.54$ $0.48$ $0.81$
+Table 2: Treatment disparities and outcome effects in S2.
+For sensitive counterfactuals (SCF), female (F) and male (M), $Y^{F}=0$: outcome changed defaulted $\rightarrow$ repaid, $Y^{F}=1$ the reverse.
+Reporting median INR (Indian Rupees) and DM (Deutsche Marks).
+Hist.
+all test-loans were granted (past policy), EOD only uses test-loans granted by the equalized odds predictor.
+5 Use Case: Lending Decisions
+In this section, we evaluate our framework by (i) auditing historical data to identify treatment disparities and their potentially discriminatory effects on outcomes, and (ii) assessing the effectiveness of our mitigation strategies.
+We evaluate using four real-world lending datasets that span diverse geographic regions, loan structures, and two representative decision-making scenarios (experimental details in Appendix C and additional results in Appendix E):
+  • Decision agreement as label (S.1): The ground-truth $Y$ captures a decision’s immediate outcome, where $Y=1$ shows agreement on treatment terms, e.g., a loan approved by the bank and terms accepted by the applicant.
+We evaluate this scenario using the 2017 HMDA dataset from New York and Texas (FFIEC 2017).
+  • Decision outcome as label (S.2): The ground-truth $Y$ captures a decision’s downstream outcome, e.g., $Y=1$ for loan repaid and $Y=0$ for default.
+We analyze this scenario using Home Credit (Montoya, Odintsov, and Kotek 2018) and German Credit (Hofmann 1994) data.
+5.1 Measuring Treatment Disparities
+We begin by auditing disparities in historical treatment assignments in test data.
+First, we examine disparities under historical binary decisions using the full test set (Hist.), where all individuals received positive decisions.
+We also evaluate the impact of predictive fairness by analyzing treatment disparities for positively predicted individuals under an equalized odds (EOD) predictor.
+Results for other predictors in Appendix E.2, analysis of unsplit data in Appendix E.1.
+S.1: Agreement of the decision-making process as label.
+In both HMDA datasets, we follow the identification and pre-processing procedures for sensitive attributes, covariates, and treatment decisions outlined in (Cooper et al.
+2024).
+We consider loan amount and preapproval status as treatment decisions within the bank’s authority.
+Table 1 reports the treatment disparity analysis results.
+Takeaway.
+Although loan agreement rates do not differ significantly by gender, our analysis (ref.
+Hist) shows that banks assign more conservative loan terms to female applicants, suggesting potential treatment unfairness.
+Offering higher loan amounts to women would not reduce agreement rates, challenging the justification for such disparities.
+We also observe that predictive fairness (e.g., EOD) does not eliminate treatment disparities.
+For example, Table 1 shows that in the NY dataset, an EOD-compliant predictor still yields disparities: females, if treated as males, would have received \$27K higher loan amounts without any negative impact on $Y$, indicating potential treatment unfairness.
+These differences are largely explained by applicant covariates, with minimal direct disparity observed.
+Our observations raise important questions about the rationale behind these disparate treatment decisions.
+Specifically, we ask: Are disparities in loan terms for female applicants justified by downstream outcomes, such as repayment behavior, and thus defensible as a business necessity?
+This motivates our second scenario examining treatment disparities, repayment outcomes, and potential discrimination.
+S.2: Decision outcome as label.
+In the Home Credit dataset, we analyze two treatment decisions: loan amount and annuity amount, while in the German Credit dataset, we examine three treatment decisions: loan amount, loan duration, and installment rate with results summarized in Table 2.
+Takeaway.
+In Scenario S.2, our results show female applicants experience negative treatment disparities not explained by downstream outcomes.
+Conversely, while males receive perceived positive treatment, they are ultimately negatively impacted by its downstream effects.
+Notably, our findings align with similar insights observed for German Credit in a simpler linear setting (Kanubala, Valera, and Gupta 2024), reinforcing the validity of our results.
+Despite appearing favorable, loan terms offered to certain males may increase their default risks and lower their creditworthiness over time.
+Applying the treatment decisions of their female counterparts to male applicants could mitigate this discrimination and improve repayment outcomes.
+Importantly, this scenario also further illustrates that predictive fairness alone does not mitigate treatment disparities or their effects on outcomes.
+In German Credit, 9.1% of male applicants who received loans from the EOD predictor but later defaulted would have repaid under counterfactual treatments.
+5.2 Mitigating Treatment Discrimination
+Our results in Table 2 show that banks exhibit negative treatment disparity against female applicants.
+However, males are potentially discriminated against since they experience a negative downstream effect on repayment ability, and treating them as females improves repayment performance.
+Hence, treating all applicants conservatively as females may enhance the overall repayment ability of borrowers, benefiting both the banks and the applicants.
+Dataset Group LGD (INR/DM) ESI (INR/DM)
+Home
+Credit
+Female $586692.67$ $290208.73$
+Male ($\mathcal{D}$) $576590.35$ $422114.31$
+Male ($\mathcal{D}^{\mathrm{fair}}$) 568157.43 391110.73
+German
+Credit
+Female $1220.51$ $312.13$
+Male ($\mathcal{D}$) $1163.88$ $495.65$
+Male ($\mathcal{D}^{\mathrm{fair}}$) 940.03 393.86
+Table 3: Fairness analysis in S.2, comparing bank’s LGD and applicant’s ESI (both: lower better) for various policies.
+$\mathcal{D}$ is factual data, in $\mathcal{D}^{\mathrm{fair}}$ all applicants treated as females.
+Fair non-binary decision-making.
+Following Sec.
+4.1, we examine the creation of a treatment-fair dataset aimed at automating future fair non-binary decision-making tasks.
+Based on our Scenario S.2 analyses, we identify males as the group needing adjusted treatment to generate our fair dataset, $\mathcal{D}^{\mathrm{fair}}$.
+To understand the impact of $\mathcal{D}^{\mathrm{fair}}$, we aim to compare the fair dataset with the original (unfair) dataset on two key metrics: the bank’s Loss Given Default ($\operatorname{LGD}$(Schuermann 2004) and the applicant’s Expected Simple Interest ($\operatorname{ESI}$(Ross, Westerfield, and Jordan 2014).
+Appendix B.5 contains further details on stakeholder losses.
+Takeaway.
+Table 3 shows at the data-level how adjusting male treatments reduces bias and benefits stakeholders by lowering $\operatorname{LGD}$ for banks and $\operatorname{ESI}$ for male borrowers, highlighting the value of fairer treatment.
+Appendix E.3 further confirms that training binary predictors on $\mathcal{D}^{\mathrm{fair}}$ improves outcome utilities while preserving predictive performance.
+Fair risk score estimation.
+Following Sec.
+4.2, we now analyze treatment-fair risk score estimates.
+Based on insights from scenario S.2, we focus on male applicants and apply different treatment distributions $\pi(Z)$ to generate risk estimates.
+We approximate $\pi(Z)$ for males under two conditions: (i) a fair interventional distribution, derived from the empirical distribution of $Z$ conditioned on female applicants to produce “fair” risk scores, and (ii) the factual “unfair” distribution, based on the empirical distribution of $Z$ conditioned on male applicants.
+For female applicants, $\pi(Z)$ is approximated using the empirical distribution of $Z$ conditioned on female applicants.
+We compare the empirical cumulative distribution functions (CDFs) of the resulting risk scores for male and female applicants in Fig.
+3.
+Takeaway.
+While males generally exhibit lower risk scores than females for German Credit, the pattern reverses for Home Credit.
+However, historical treatment practices have led to an unfair overestimation of males’ risk scores, especially in German Credit.
+This bias is corrected when the risk scores are recalculated under the fair interventional distribution, adjusting for potentially unfair treatments.
+Integrating these fair risk scores into decision-making pipelines mitigates the residual impacts of historical treatment biases, paving the way for more fair decision-making processes.
+Figure 3: Fair risk score estimation in S.2 for females and males under different treatment distributions $\pi(Z)$.
+Unfair estimates use the factual distribution; fair estimates apply an interventional distribution, treating all applicants as female.
+6 Discussion
+Through a novel causal framework, our work highlighted the overlooked role of non-binary treatment decisions in fairness analyses.
+This section notes several open challenges (additional discussion in Appendix B.4).
+Robust causal frameworks.
+Our approach assumes access to ground-truth outcomes and no hidden confounders, assumptions that often fail in practice.
+Hidden confounders (Kilbertus et al.
+2020a) and proxy labels for unmeasured outcomes (Mhasawade, D’Amour, and Pfohl 2024) can bias fairness measures, particularly when different stakeholders (e.g., decision-makers vs.
+auditors) apply the framework.
+Incorporating emerging methods (Almodóvar et al.
+2025) is a key step toward addressing these limitations.
+Considering diverse stakeholders.
+Although our pre-processing approach mitigates treatment discrimination, it lacks theoretical guarantees of optimality.
+Future work should benchmark our method against existing interventions (Coston et al.
+2020).
+Moreover, such approaches may not satisfy differing stakeholder objectives, raising a central question: What defines a fair treatment policy that accounts for all stakeholders?
+For instance, banks may prioritize repayment, while applicants seek lower interest rates (O’Neil and Gunn 2020).
+Achieving fairness thus requires a holistic approach to learning policies that balance these utilities.
+Collecting treatment data.
+While our framework generalizes to various domains (Appendix B.1), we evaluated only in lending owing to major data challenges in other domains.
+Many datasets conflate covariates with treatments, omit treatment details (e.g., COMPAS (Angwin et al.
+2022) lacks bail terms; lending datasets exclude interest rates), or oversimplify complex treatments into binary variables (e.g., IHDP (Madras et al.
+2019) reduces care to a single binary indicator).
+Moreover, most datasets are subject to selective labeling (Lakkaraju et al.
+2017), where outcomes are observed only when both decision-makers and individuals agree on terms, ignoring potential agreement-phase discrimination.
+These limitations underscore the need for transparent data collection across domains that includes full treatment assignments (Centre for Public Data 2023) and agreement-phase decisions to enable more comprehensive analyses.
+Conclusion.
+Our work underscores the importance of analyzing the fairness of non-binary treatment decisions given positive binary decisions, e.g., loan approvals.
+Our novel treatment disparity measures extend existing fairness frameworks toward a more comprehensive evaluation of algorithmic decision-making.
+Our findings show that fairness cannot rely on technical solutions alone; it requires socio-technical approaches that account for diverse stakeholder utilities.
+By capturing the nuanced impacts of treatment decisions, our framework advances aligning algorithmic systems with societal values, working toward jointly addressing fairness in binary predictions and non-binary treatment decisions.
+Acknowledgements
+We thank Adrián Javaloy for providing invaluable guidance and feedback regarding the causal generative models.
+This work has been funded by the European Union (ERC-2021-STG, SAML, 101040177).
+However, the views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
+Neither the European Union nor the granting authority can be held responsible.
+Appendix A Extended Related Work
+Fairness in lending.
+Fairness analyses in automated lending decisions have focused mainly on statistical methods (Hardt, Price, and Srebro 2016; Hurlin, Pérignon, and Saurin 2022; Liu et al.
+2018), emphasizing metrics like equality of opportunity, as noted in surveys (Kozodoi, Jacob, and Lessmann 2022; Fuster et al.
+2022).
+However, these approaches often overlook biases embedded in the ground-truth labels, which stem from unfair treatment during loan processing.
+While prior work has separately examined treatment discrimination (Sackey and Amponsah 2018; Agier and Szafarz 2010) and biased outcomes in lending decisions (Munnell et al.
+1996; Alesina, Lotti, and Mistrulli 2013; Montoya et al.
+2020), the causal link between biased treatment and adverse outcomes remains underexplored.
+Our novel framework bridges this gap and jointly analyzes treatment discrimination and its causal effects on lending outcomes.
+Causal notions of predictive fairness.
+Causal predictive fairness notions examine the causal influence of sensitive features on decisions and outcomes.
+Kusner et al.
+(2017) initiated the study of the total causal effect of sensitive attributes, followed by work on path-specific effects (Nabi and Shpitser 2018; Chiappa 2019).
+Subsequent research extended this to multi-level features (Mhasawade and Chunara 2021) and outcome fairness (Plecko and Bareinboim 2023).
+However, these approaches largely focused on binary decisions or outcomes, overlooking decision-makers’ complex treatment decisions.
+While Madras et al.
+(2019) proposed a causal treatment fairness framework, it reduced non-binary treatments to binary variables, which limits its expressiveness.
+We address this gap by incorporating the decision-maker’s perspective to provide a fine-grained analysis of treatment decisions and their causal effects.
+Furthermore, we improve the path-specific counterfactual accuracy in societal contexts using the state-of-the-art CNFs (Javaloy, Sánchez-Martín, and Valera 2024), outperforming earlier variational approaches (Chiappa 2019; Madras et al.
+2019).
+Multi-stage selection problems.
+Multi-stage selection problems (Emelianov et al.
+2019; Jia et al.
+2024) explore automating decision-making with binary decisions across stages, revealing additional candidate covariates for subsequent decisions.
+Concerning this line of work, our causal framework addresses a fundamentally different context governed by different mechanisms.
+Our work focuses on a single-stage decision process where binary decisions (e.g., loan approval) are made in one step and treatment decisions are directly controlled by decision-makers.
+We complement existing studies by analyzing the fairness of assigned multivariate, non-binary treatments and their causal impact on decision outcomes.
+Appendix B Methodology Details
+B.1 Treatment Decisions Across Different Domains
 Domain Sensitives $S$ Covariates $X$
@@
 Decisions $Z$
-Outcomes $Y$
 Binary
 Decisions
+Outcomes $Y$
 Lending gender, age, race
@@
 duration, interest rate…)
-loan repaid loan approved
+loan approved loan repaid
 Justice
@@
 conditions…)
-re-offence bail approved
+bail approved re-offence
 Health
@@
 (medication, surgery…)
-re-hospitalization hospitalization
+hospitalization re-hospitalization
 Insurance gender, age, race
@@
 (cost, coverage…)
-claim approval policy approved
+policy approved claim approval
 Hiring gender, age, race
@@
 (working hours, salary…)
-job performance job offered
-Data-driven approaches increasingly leverage historical data to automate decision-making processes across societal domains, such as finance [4, 49], healthcare [20], and criminal justice [13].
-These systems typically reduce complex decision-making to a binary classification task, predicting either ground-truth outcomes (e.g., loan repayment, recidivism after bail) or imperfect proxies (e.g., healthcare expenditures as a proxy for health status) [44, 53].
-However, real-world decisions are rarely as simple as a binary choice; they involve multiple, non-binary treatment decisions.
-As shown in Table 1, decision-makers often determine not only whether to give a positive decision (provide treatment, e.g., approve loan) but also its specific terms (e.g., loan amount, duration, interest rate).
-Similar oversimplifications occur in other domains, such as bail decisions [61].
-This simplistic approach overlooks the impact of non-binary treatment decisions on outcomes, leading to fairness analyses that fail to provide a holistic view of existing systems [46].
-Despite growing research on fairness in algorithmic decision-making, existing analyses simplify complex decision-making mechanisms in three key ways.
-First, many (if not most) studies [69, 21, 11] include non-binary treatments as covariates of the decision-subject (covariate examples in Table 1).
-This assumption obscures the mechanisms that assign treatment decisions, which are not controlled by decision-subjects but by decision-makers and, thus, may reveal discriminatory patterns.
-This assumption thus limits the ability of these works to account for all sources of unfairness and their implications.
-Second, alternative works [39, 12, 57] examine the fairness implications of treatment decisions but assume they are binary, ignoring the multivariate, non-binary nature of treatment decisions (such as, e.g., loan or bail amount).
-Finally, only a few works [2, 17, 3] have focused on disparities in non-binary treatment decisions across different groups.
-Still, they lack a generalizable framework, providing only empirical insights in financial domains and, more importantly, neglecting the discriminatory downstream effects of these disparities on outcomes.
-To bridge gaps in the existing approaches, we propose a fundamental paradigm shift in fairness analysis of automated decision-making systems that brings fairness considerations to non-binary treatment decisions and their downstream outcome effects.
-Our paradigm shift leverages the following key insight: decision-makers do not simply make binary decisions (whether to give treatment) but also control the assignment of adjacent non-binary treatment decisions, which can introduce disparities and negatively impact outcomes.
-To incorporate this insight into fairness analysis, we introduce a broadly applicable causal framework that allows decision-makers to analyze the fairness of underlying decision-making mechanisms systematically.
-This systematic analysis allows our framework to provide novel measures of treatment disparities, assess their discriminatory effects on outcomes, and enable mitigation strategies to prevent biases in non-binary treatment decisions from perpetuating unfairness in automated systems.
-Our approach explicitly differentiates between a decision-subject’s covariates and a decision-maker’s treatment decisions in the causal data-generating process, as illustrated in Figure 1.
-This causal formulation allows us to measure disparities in non-binary treatment decisions $Z$ induced by sensitive attributes $S$, both directly (solid red path) and indirectly through covariates $X$ (dotted red path).
-Furthermore, our framework quantifies how these disparities propagate to outcomes $Y$ via the causal path from $Z$ to $Y$ (dashed orange path), thereby inducing treatment discrimination.
-Our framework supports mitigation in scenarios impacted by treatment discrimination through two pre-processing strategies.
-These techniques generate treatment-fair counterfactual datasets, enabling fairer non-binary automated decisions and more equitable risk score estimation to inform future decision-making.
-Our framework advances fairness analyses by leveraging path-specific counterfactuals to quantify direct and indirect disparities in non-binary treatment decisions and their impact on outcomes.
-While path-specific counterfactuals provide a powerful tool for causal fairness analyses, existing methods [52, 8] suffer from significant limitations, such as being restricted to linear data generation processes or requiring inefficient Monte Carlo estimation with strong regularization.
-To overcome these challenges, we build on the state-of-the-art in causal inference, the causal normalizing flows [26], and propose a novel path-specific counterfactual generation method.
-This advancement enables our framework to be theoretically grounded while being practical to be applied in real-world decision-making systems.
-Finally, we validate our framework using four real-world lending datasets across two decision-making scenarios.
-Our empirical results demonstrate that explicitly accounting for non-binary treatment decisions helps provide more holistic fairness assessments, identifies systemic biases, and supports mitigation efforts.
-These findings underscore the importance of addressing treatment discrimination in achieving fairer outcomes, bridging a critical gap in fairness research, and laying the foundation for more robust causal fairness analyses.
-$X$$Z$$Y$$S$
-Figure 1: Causal graph of the data generation process, distinguishing decision-subject covariates $X$ and treatment decisions $Z$.
-Under the assumption that a positive binary decision has already been made.
-Red paths (solid: direct, dotted: indirect) show causal paths to $Z$ influenced by sensitive attribute $S$, highlighting potential sources of disparities.
-The orange dashed path from $Z$ to $Y$ captures the potentially discriminatory effects on the outcome.
-2 Background
-Our framework for measuring and mitigating treatment discrimination in decision-making relies on causal analyses. This section briefly outlines the core causal concepts used throughout the paper.
-Structural causal models (SCMs).
-An SCM [56] $\mathcal{M}=(V,U,F)$ is defined by the observational features $V$, independent, unobserved, exogenous variables $U$ (no hidden confounding) and structural equations $F$.
-A structural equation for some $V_{i}$ gives the functional mapping from the corresponding exogenous $U_{i}$ and causal parents $\operatorname{pa}(V_{i})$ to the value of $V_{i}$ as $V_{i}=F_{i}(\operatorname{pa}(V_{i}),U_{i})$.
-These cause-effect relations in $\mathcal{M}$ can also be represented by a directed acyclic graph $\mathcal{G}$ like in Figure 1, where the nodes represent $V$, and the directed edges represent the cause-effect relations.
-Performing interventions.
-The SCM $\mathcal{M}$ helps us estimate the effects of hypothetical interventions $\mathcal{I}$ on features.
-For instance, to analyze the effect of the intervention $\mathcal{I}$ setting $V_{i}=\alpha$, we use the ${\operatorname{do}}$-operator [56] ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ on $\mathcal{M}$ that generates the intervened SCM $\mathcal{M}^{\mathcal{I}}$.
-Counterfactuals.
-In the absence of hidden confounders, this intervened SCM $\mathcal{M}^{\mathcal{I}}$ can also be used to perform counterfactual analyses for a particular factual decision-subject with features $\vartheta^{F}$ to ask “what would be the counterfactual features $\vartheta^{CF}$ had we intervened with ${\operatorname{do}}(V_{i}\rightarrow\alpha)$, and all exogenous factors $u^{F}$ remaining unchanged.” Counterfactuals $\vartheta^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))$ are computed following the three steps [56]: i).
-abduction: estimating $u^{F}$ from $\vartheta^{F}$ using $\mathcal{M}$, ii).
-action: performing ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ to get $\mathcal{M}^{\mathcal{I}}$, iii).
-prediction: compute $\vartheta^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))$ from $\mathcal{M}^{\mathcal{I}}$ using $u^{F}$ and setting $V_{i}=\alpha$.
-In simple terms, computing counterfactuals estimates the total effect of an intervention on the downstream causal features of the intervened variable in the causal graph $\mathcal{G}$ for the factual decision-subject $\vartheta^{F}$.
-Path-specific counterfactuals.
-In certain scenarios, it is crucial to understand the effects of a particular feature $V_{i}$ on a downstream feature $V_{j}$ along different causal paths in $\mathcal{G}$.
-This analysis can be done by computing path-specific counterfactuals [8], which use multiple ${\operatorname{do}}$ operations to control for specific causal paths.
-Consider the two paths in an example causal structuring between variables $\langle V_{i},V_{j},V_{k}\rangle$, where we have $V_{i}\rightarrow V_{k}$ and $V_{i}\rightarrow V_{j}\rightarrow V_{k}$.
-For a factual instance $\vartheta^{F}=\langle v_{i}^{F},v_{j}^{F},v_{k}^{F}\rangle$, we can compute the path-specific counterfactual effect of ${\operatorname{do}}(V_{i}\rightarrow\alpha)$ along these paths.
-For the path $V_{i}\rightarrow V_{k}$, we measure the counterfactual effect using $v_{k}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha),{\operatorname{do}}(V_{j}\rightarrow v_{j}^{F}))$.
-Similarly, for the path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$, we use $v_{k}^{CF}(v_{i}^{F},{\operatorname{do}}(V_{j}\rightarrow v_{j}^{CF}({\operatorname{do}}(V_{i}\rightarrow\alpha))))$.
-For the path $V_{i}\rightarrow V_{j}\rightarrow V_{k}$, we use the counterfactual $v_{j}^{CF}$ obtained from the intervention on $V_{i}$ but keep $V_{i}=v_{i}^{F}$ fixed along the other path.
-Causal fairness analyses.
-The SCM framework has been extensively used to causally study fairness problems in ML [35, 8].
-In these scenarios, target labels $Y$ (e.g., bail, loan repayment) are considered causally affected by the sensitive feature $S$ (e.g., gender/sex, race).
-We can estimate the total effect of $S$ on $Y$ by finding sensitive counterfactuals $y^{\mathrm{SCF}}=y^{CF}({\operatorname{do}}(S\rightarrow\alpha))$, as was shown in counterfactual fairness [35].
-Similarly, Chiappa [8] used path-specific counterfactuals $y^{\mathrm{PSCF}}$ to analyze the potentially discriminatory effect of $S$ on $Y$ along different causal paths in $\mathcal{G}$ for linear SCMs.
-In this work, we extend the concept of counterfactual fairness and path-specific counterfactual fairness for treatment decisions and their downstream effects.
-3 Measuring treatment disparities and their downstream effects
-In this section, we introduce our causal framework to address the limitations of the existing fair decision-making literature by explicitly measuring disparities in treatment decisions and their downstream effects and mitigating potential discriminatory consequences.
-To this end, we assume that the data generation process that produces the observed (factual) historical data $\mathcal{D}$ follows the causal graph depicted in Figure 1, where:
+job offered job performance
+Table 4: Examples of data collected across various decision-making domains, highlighting the distinction between sensitive attributes $S$, decision-subject covariates $X$, treatment decisions $Z$, and outcomes $Y$.
+The multivariate, non-binary $Z$ have been represented previously in the literature as a positive binary decision (in gray).
+Table 4 details treatment decisions and covariates across various domains relevant to algorithmic fairness.
+For each domain, we identify the treatment decision features, which are those that are under the decision-maker’s control and directly influence the outcome.
+The table also specifies the binary decision for each domain.
+Importantly, these binary decisions may serve as either ground truth labels or proxies for the underlying outcome of interest.
+For example, the criminal justice domain: while automated systems typically predict bail approval decisions, the true objective is assessing defendant risk.
+Judges use this risk assessment to make the binary bail decision and consequently determine bail terms.
+However, focusing solely on the fairness of the binary decisions of bail approval might fail to address the unfairness in the bail terms.
+Existing research in the social sciences and medicine has shown potential biases in treatment decisions.
+For instance, legal scholars have shown potential bias issues in bail treatment decisions (Arnold, Dobbie, and Yang 2018), while some quantitative research has shown unexplained gender biases in insurance costs (Pernagallo, Punzo, and Torrisi 2024).
+Similarly, Castilla (2008) empirically analyzed potential gender disparities in salary hikes in organizations.
+In healthcare, potential biases were also shown to exist in several treatment factors (Hershman et al.
+2005).
+Our proposed causal framework can be readily used to study treatment disparities and outcome effects across all these representative decision-making domains.
+Importantly, our framework can allow for a more comprehensive evaluation of such real-world scenarios using state-of-the-art computational methods.
+However, accessing appropriate datasets across most domains remains a significant challenge, as stated by existing research (Pernagallo, Punzo, and Torrisi 2024).
+Even when some data is available, they lack sufficient fine-grained information about treatment decisions provided by the decision-makers.
+With future advancements in data availability, our proposed causal framework can be used to study potential treatment disparities and fairness issues across multiple domains.
+B.2 Background on Causal Normalizing Flows
+Causal normalizing flows (CNF) (Javaloy, Sánchez-Martín, and Valera 2024) enable causal inference from observational data under minor assumptions.
+It provides an ideal practical methodology to estimate unknown causal relations from observational data and an accurate technique to estimate interventional and counterfactual quantities.
+Here, we provide a brief background on CNFs and refer the reader to the original paper for full details.
+Assumptions
+The work assumes the following for its theoretical grounding.
+First, it assumes a diffeomorphic data generation process.
+Hence, the structural equation functions $F$ in the SCM are invertible, and the functions and their inverse are differentiable.
+Second, it assumes the causal mechanism has no feedback loops; hence, it can be represented by a directed acyclic graph (DAG).
+Third, it assumes causal sufficiency.
+Hence, the exogenous variables for each feature are mutually independent, i.e., $p(U)=\prod_{i}p(u_{i})$.
+Finally, it assumes that a causal ordering is given across the features.
+Autoregressive architecture
+Under the assumptions, the authors show that SCM mechanisms correspond to a special class of autoregressive functions known as triangular monotone increasing (TMI) maps.
+Through TMI maps, the SCM is represented by triangular, invertible functions.
+Each feature in $V$ is a monotone function of its parents $pa(V)$ and its exogenous noise term $u$.
+These maps can be estimated using autoregressive normalizing flows (ANFs), where each conditional is modeled as a monotone transformation, forming the causal normalizing flows model.
+The CNF defines a parameterized function $T_{\psi}$ that maps exogenous variables $U$ to observed features $V$, serving as a generative model for the features.
+Leveraging the autoregressive structure, the inverse $T_{\psi}^{-1}$ maps $V$ back to the exogenous space $U$.
+Since we assume the causal topological order is known, each conditional in the flow corresponds directly to a structural equation in the SCM.
+By modeling the SCM in this way, CNFs can flexibly approximate the joint distribution over observed variables while remaining consistent with the underlying causal structure.
+For further details on the architecture and training procedure, we refer to the main paper (Javaloy, Sánchez-Martín, and Valera 2024).
+Discussion on Theoretical Guarantees
+Under the stated assumptions and leveraging the flow-based approximation of the SCM, the authors show that if the CNF $(T_{\psi},p(U))$ reproduces the observational distribution generated by the true SCM, then the exogenous variables learned by the flow differ from the true exogenous noise variables only by component-wise invertible transformations.
+Specifically, for each feature $i$, the corresponding exogenous factor learned by the CNF is a transformation $h_{i}(u_{i})$ of the true exogenous variable $u_{i}$, where $h_{i}$ is an invertible function.
+This result establishes that CNFs guarantee identifiability of exogenous variables up to invertible transformations, and by successfully isolating the exogenous information, they ensure causal consistency in the learned generative process.
+Interventions and do-operation
+While the traditional three-step process of Pearl (Pearl et al.
+2000) manipulates the features directly for interventions, CNFs perform interventions by manipulating the latent exogenous variables.
+The process is discussed in Section 3.3.
+For generating counterfactuals, we first use the flows model to map the observed features $\vartheta^{F}$ to exogenous factors as $u=T_{\psi}(\vartheta^{F})$.
+Next, for the intervention ${\operatorname{do}}(V_{j}\rightarrow\alpha)$, we recompute the exogenous factor specifically for feature $j$, hence, $u_{j}=T_{\psi}({\operatorname{do}}(V_{j}\rightarrow\alpha))$.
+Hence, in the exogenous space, only for the exact intervened feature, the exogenous variable is updated.
+For all other features, they remain the same.
+Finally, one uses the inverse of the flows model $T^{-1}_{\psi}$ to go from post-interventional $u$ to $\vartheta^{CF}$.
+Further details on the $\operatorname{do}$}-operation in the CNF can be found in Appendix C in (Javaloy, Sánchez-Martín, and Valera 2024).
+Handling partial causal graphs
+A key strength of CNFs, which we leverage in our work, is their ability to learn from partial causal graphs, where the complete causal graph is unknown, but the causal ordering between subsets of observed features is available.
+Each subset of features with unknown internal ordering can be treated as a block, referred to in (Javaloy, Sánchez-Martín, and Valera 2024) as a strongly connected component (SCC).
+Given knowledge of the causal ordering between these blocks, the CNF applies autoregressive flow modeling across the blocks.
+Within each block, the joint distribution is modeled without assuming any specific ordering.
+Consequently, the CNF retains its theoretical guarantees and identifiability properties at the block level, while identifiability is not ensured within individual blocks.
+In our causal framework, we adopt this formulation by treating all covariates $X$ and all treatment decisions $Z$ as separate blocks.
+Through our counterfactual analysis of treatment and outcome disparities, we evaluate block-level counterfactuals for covariates and treatments to study their causal effects on the downstream outcome $Y$.
+B.3 Extending to Multiple Sensitive Attributes
+Our definitions for measuring treatment disparities and their outcome effects (Definitions 1-4) can be naturally extended to scenarios where we have multi-valued and multiple sensitive attributes, e.g, non-binary gender identities, race/ethnicity, religion, etc.
+Let us consider $K$ different sensitive attributes, where each attribute $S_{i}$ has unique values $\mathcal{S}_{i}$.
+Note that this naturally considers binary and multi-valued categorical sensitive attributes.
+The combined set of all possible sensitive attribute values can be denoted as $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times\dots\times\mathcal{S}_{K}$.
+To compute counterfactual treatment decisions and outcomes, a central challenge is interpreting what constitutes a meaningful counterfactual across multiple sensitive attributes.
+Multiple plausible counterfactuals may exist (depending on which attributes are considered), and we outline two different strategies here.
+  • Single flip.
+We change one sensitive attribute at a time (e.g., only gender), holding the others fixed to their factual values.
+For example, what treatment would an individual receive if their gender were different but their race and religion stayed the same?
+For attribute $S_{i}$, the counterfactual treatment can be defined for any value $s^{\prime}_{i}$ as:
+    $z^{\mathrm{SCF}}=z^{CF}({\operatorname{do}}(S_{i}\rightarrow s^{\prime}_{i})),\forall\,s^{\prime}_{i}\in\mathcal{S}_{i}\setminus\{s^{F}_{i}\}$
+  • Joint flip.
+We simultaneously change all sensitive attributes to a new disadvantaged group (changing the value set of $S$ from $s^{F}$ to a different value set $s^{\prime}$).
+For example, an individual who is factually a White male Christian would be compared to a Black female Muslim.
+Formally, the counterfactual treatment would be:
+    $z^{\mathrm{SCF}}=z^{CF}({\operatorname{do}}(S\rightarrow s^{\prime})),\forall\,s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}$
+Our framework accommodates both counterfactual strategies, enabling interventions ${\operatorname{do}}(S\rightarrow s^{\prime})$ across multiple sensitive attributes using the estimated CNF model.
+However, each choice corresponds to a distinct fairness question of interest, whether one is concerned with the effects of individual attributes or disparities across entire groups.
+The work of (Wastvedt, Huling, and Wolfson 2024; Alvarez and Ruggieri 2025) defines counterfactuals involving multiple sensitive attributes through a joint flip, whereas (Yang, Loftus, and Stoyanovich 2020) defines counterfactuals relative to a chosen reference subgroup.
+Treatment disparities.
+Since we now need to consider potentially multiple sensitive counterfactuals across different sensitive attributes and values, we have to extend our definitions to aggregate disparities properly.
+For treatment-level disparities (Definitions 1 and 2), we can apply an aggregation operator $\mathcal{A}$ over the absolute treatment differences $\big|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})\big|$ before taking the expectation over the dataset $\mathcal{D}$.
+For instance, the multi-valued TTD would be:
+$\operatorname{TTD}_{\text{M}}(s^{F})=\mathbb{E}_{\mathcal{D}}\Big[\mathcal{A}\big(\{|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})|\}_{s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}}\big)\Big].$ (9)
+Following (Wastvedt, Huling, and Wolfson 2024), we can consider different types of aggregate functions.
   1. 1.
-    The treatment decision $Z$ is a causal child of both $X$ and (potentially) $S$.
-We assume that the causal mechanism that generated $Z$ is under the control of the decision-maker (e.g., we assume that banks decide not only whether or not to grant a loan but also the loan terms) and may be discriminatory.
-The direct causal path from $S$ to $Z$ is shown in solid red, while the indirect path mediated by $X$ is shown in dashed red.
+    Average:
+    $\mathcal{A}_{\text{AVG}}(s^{F})=\frac{1}{(|\mathcal{S}|-1)|\mathcal{S}|/2}\sum_{s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}}\big|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})\big|$
   2. 2.
-    The treatment decision $Z$ is a causal parent of the outcome variable $Y$ (e.g., we assume loan terms can impact the repayment ability of applicants).
-Hence, $S$ can influence $Y$ mediated by $Z$ (orange dashed line in Figure 1), leading to the possible propagation of treatment disparities to the outcome.
-Moreover, we maintain the direct $S\rightarrow Y$ since such effects may occur in specific scenarios, e.g., in healthcare.
-Remarks.
-The edges of our causal graph $\mathcal{G}$ (Figure 1) indicate the structural causal ordering [26] among the different groups of features, where the cause-effect relation for some edges may be zero.
-Moreover, although the figure shows a single sensitive feature for simplicity, our $\mathcal{G}$ can consider multiple sensitive features as independent root nodes.
-We treat $X$ and $Z$ in $\mathcal{G}$ as multivariate blocks, focusing our causal analyses on relations between these “blocks” without making explicit assumptions about their causal relations.
-3.1 Disparities in non-binary treatment decisions
-Given our proposed causal graph in Figure 1, we use the SCM framework to measure disparities in treatment decisions by measuring counterfactual and path-specific counterfactual effects of the sensitive feature (e.g., gender) on the treatment decision variables.
-For simplicity and without loss of generalization, here we assume the sensitive feature on which we intervene to be binary, i.e., $S\in\{0,1\}$.
-[^1] Our measures can easily be extended for multi-valued categorical sensitive attributes, e.g., by conducting group-pairwise comparisons.
-Intersectional disparity measurements across multiple sensitive attributes are left for future work.
-Given an observed dataset $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$, disparity function $\mathcal{F}$, and aggregation function $\mathcal{A}_{\mathcal{D}}$ regarding $\mathcal{D}$, we define two treatment disparity measures.
-Definition 1.
-Total Treatment Disparity (TTD), against decision-subjects with (factual) sensitive feature $s^{F}$, measures the expected difference between the treatment $z^{F}$ they received and the treatment $z^{\mathrm{SCF}}$ they would have received if they belonged to a different sensitive group, i.e.,
-$\operatorname{TTD}(s^{F})=\mathcal{A}_{\mathcal{D}}\left(\mathcal{F}(z^{\mathrm{SCF}},z^{F})\right),\,\mathrm{where}\,z^{\mathrm{SCF}}=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})))$ (1)
-Related to [35], the ${\operatorname{do}}$-operator on the sensitive feature $S$ gives the sensitive counterfactual treatment $z^{\mathrm{SCF}}$ a decision-subject would have received if their sensitive group were different.
-Definition 2.
-Direct Treatment Disparity (DTD), against decision-subjects with (factual) sensitive feature $s^{F}$, measures the expected difference between the treatment they received and the sensitive (direct) - path counterfactual treatment, which assumes a ${\operatorname{do}}(S\rightarrow(1-s^{F})))$ only on the direct path from $S$ to $Z$, i.e.,
-$\begin{split}\operatorname{DTD}(s^{F})=\mathcal{A}_{\mathcal{D}}\left(\mathcal{F}(z^{\mathrm{SPCF}},z^{F})\right),\,\mathrm{where}\,z^{\mathrm{SPCF}}=z^{CF}({\operatorname{do}}(S\rightarrow(1-s^{F})),{\operatorname{do}}(X\rightarrow x^{F}))\end{split}$ (2)
-Leveraging path-specific counterfactuals [52, 8], we compute the sensitive path counterfactual treatments $z^{\mathrm{SPCF}}$ a decision-subject would have received for different sensitive group memberships while keeping original covariates $x^{F}$ fixed.
-Remarks.
-First, the nature of $\mathcal{A},\mathcal{F}$ in the above definitions depends on the nature of the treatment decision variables.
-For instance, when treatment decisions are continuous, $\mathcal{A}$ would be the expectation $\mathbb{E}$ over $\mathcal{D}$, and $\mathcal{F}$ would denote the differences across counterfactual treatments.
-Second, by drawing parallels with predictive fairness concepts, we can better understand the fairness implications of treatment disparities.
-For instance, $\operatorname{DTD}$ highlights disparities caused by directly incorporating $S$ into treatments, which may be acceptable in areas such as healthcare but undesirable in contexts such as lending, aligning with the predictive fairness concept of disparate treatment [15].
-Similarly, zero $\operatorname{TTD}$ requires that treatments $Z$ be independent of $S$, showing parallels with demographic parity [15], which demands that predictions do not depend on $S$.
-However, like predictive fairness, the relevance of perfect treatment parity depends on context.
-Similar to demographic parity scenarios, when ground-truth outcome $Y$ is unavailable (e.g., admissions, hiring), treatment disparities may be considered discriminatory.
-On the other hand, as separation-based frameworks [21, 69] show, the outcome $Y$ can explain predictive disparities in certain contexts, e.g., when $Y$ is a ground truth as in lending.
-Hence, studying the downstream impact on $Y$ is imperative to understand if treatment disparities are discriminatory.
-3.2 Downstream effects of treatment disparities
-While Definitions 1 and 2 help measure treatment disparities, it is crucial to assess the downstream impacts of these disparities on the outcome $Y$ (e.g., when $Y$ is a ground truth as in lending) to determine whether they are discriminatory.
-Hence, we introduce two novel measures to quantify the counterfactual downstream outcome effects of total and direct treatment disparities.
-Definition 3.
-Total Treatment Disparity Effect (TTD-E) measures the downstream effect of the sensitive counterfactual treatment $z^{\mathrm{SCF}}$, computed in Eq.
-1, as the probability of a label change from the factual value $y^{F}$, i.e.,
-$\begin{split}\operatorname{TTD-E}(s^{F},y^{F})&=\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}\left(y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))\neq y^{F}\right)\right],\end{split}$ (3)
-where $\mathbb{I}$ is an indicator function.
-This definition helps us understand how the total causal effect of $S$ on the treatment decisions $Z$ ends up impacting the decision outcomes $Y$.
-Intuitively, for any decision-subject belonging to $s^{F}$ with covariates $x^{F}$, the decision-maker provides the sensitive counterfactual $z^{\mathrm{SCF}}$ treatments to measure the outcome change.
-Definition 4.
-Direct Treatment Disparity Effect (DTD-E) measures the downstream effect of the sensitive path counterfactual treatment $z^{\mathrm{SPCF}}$, computed in Eq.
-2, as the probability of a label change from the factual value $y^{F}$, i.e.,
-$\operatorname{DTD-E}(s^{F},y^{F})$ $=\mathbb{E}_{\mathcal{D}}\left[\mathbb{I}\left(y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SPCF}}))\neq y^{F}\right)\right],$ (4)
-where $\mathbb{I}$ is an indicator function.
-This definition helps us understand how the direct causal effect of $S$ on the treatment decisions $Z$ ends up impacting the decision outcomes $Y$.
-Intuitively, for any decision-subject belonging to $s^{F}$ with covariates $x^{F}$, the decision-maker provides the sensitive path counterfactual $z^{\mathrm{SPCF}}$ treatments to measure the outcome change.
-Remarks.
-As before, we establish connections to predictive fairness concepts to understand when treatment disparities can result in discrimination through unfair decision outcomes.
-For specific contexts, we can relate to the perspective of disparate treatment [15] and consider any direct use of $S$ in assigning $Z$ problematic and any consequent effect on $Y$ as discriminatory.
-In other scenarios, e.g., healthcare, such effects may not be discriminatory if the effect of DTD on $Y$ is beneficial.
-Similarly, disparities between treatment decisions $z^{\mathrm{SCF}}$ and $z^{F}$ can lead to unfairness in $Y$.
-For instance, in predictive fairness, separation [69] deems decision disparities justifiable if they do not result in adverse ground-truth outcomes.
-Similarly, treatment disparities regarding $S$ may be deemed discriminatory if they do not have a positive impact on the ground-truth $Y$, e.g., when assigning $z^{F}$ results in negative outcomes ($y^{F}$), whereas $z^{\mathrm{SCF}}$ would have led to more favorable outcomes ($y^{CF}$).
-3.3 Practical implementation
-Prior works that computed path-specific counterfactuals for predictive fairness [52, 8] are restricted to linear settings or must rely on imprecise Monte Carlo approximations and complex regularization.
-To compute path-specific counterfactuals in more practical settings, we propose a novel framework that helps estimate our proposed disparity measures.
-Assuming access only to factual (potentially unfair) data $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$ and the causal graph $\mathcal{G}$ (Figure 1), our framework builds on the causal normalizing flows (CNF) [26].
-CNF has demonstrated accurate causal estimates in societal decision-making settings [41].
-Hence, by building our framework based on CNF, we develop a practical yet theoretically grounded approach to compute path-specific counterfactuals.
-3.3.1 Background on causal normalizing flows
-Causal normalizing flows (CNF) approximate the true and unknown SCM $\mathcal{M}$ using a single, invertible normalizing flows model $T_{\psi}$ with parameters $\psi$.
-Note that CNF can handle partial causal graphs, in which each random variable in the graph may be multivariate without explicitly modeling the causal relationships across its dimensions.
-Moreover, given a causal ordering across the random variables, CNF is highly flexible in learning the presence or absence of causal effects from observational data.
-Hence, even with data collected from a (potentially unfair) decision-making process, the CNF model provides an ideal practical framework to estimate our discrimination measures regarding treatment decisions and their downstream effects.
-Importantly, a causal normalizing flows model fitted on observational data can be used to estimate the counterfactual datapoint $\vartheta^{CF}$ resulting from a causal intervention $({\operatorname{do}}(V_{j}\rightarrow\alpha))$ on a given factual datapoint $\vartheta^{F}=\{v_{1}^{F},\ldots,v_{K}^{F}\}$ via:
-  1. 1.
-    abduction step computing the exogenous values for $\vartheta^{F}$ as ${u}^{F}=T_{\psi}(\vartheta^{F})$;
-  2. 2.
-    action step constructing the exogenous values for $\vartheta^{CF}$ as ${u}^{CF}({\operatorname{do}}(V_{j}\rightarrow\alpha))=\{{u}^{F}_{1:j-1},T_{\psi}({\operatorname{do}}(V_{j}\rightarrow\alpha)),{u}^{F}_{j+1:K}\}$;
+    Maximum absolute error:
+    $\mathcal{A}_{\text{MAX}}(s^{F})=\max_{s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}}\big|\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})\big|$
   3. 3.
-    prediction step estimating the resulting counterfactual feature values from the exogenous variables as $\vartheta^{CF}=T_{\psi}^{-1}\left({u}^{CF}({\operatorname{do}}(V_{j}\rightarrow\alpha))\right)$.
-Refer to Appendix A for more details on causal normalizing flows. Next, we show how to adapt the causal normalizing flows for our non-binary decision-making fairness analysis framework.
-3.3.2 Approximating counterfactual treatments and outcomes
-We can use a causal normalizing flows model trained using factual $\mathcal{D}$ to compute counterfactual quantities for some factual $\vartheta^{F}\in\mathcal{D}$.
-Importantly, we can readily utilize the existing counterfactual estimation method shown above to compute the sensitive total treatment counterfactual $z^{\mathrm{SCF}}$ as:
-$z^{\mathrm{SCF}}=T^{-1}_{\psi}(\langle T_{\psi}({\operatorname{do}}(s\rightarrow(1-s^{F}))),{u}^{F}_{X},{u}^{F}_{Z},{u}^{F}_{Y}\rangle)$ (5)
-Similarly, we can directly use the approximation process in Section 3.3.1 to compute the counterfactual labels in Definitions 3 and 4 resulting from an treatment intervention ${\operatorname{do}}(Z\rightarrow\hat{z})$, as:
-$y^{CF}({\operatorname{do}}(Z\rightarrow\hat{z}))=T^{-1}_{\psi}(\langle{u}^{F}_{S},{u}^{F}_{X},T_{\psi}({\operatorname{do}}(z\rightarrow\hat{z})),{u}^{F}_{Y}\rangle),$ (6)
-where $\hat{z}$ denotes respectively $z^{\mathrm{SCF}}$ from Definition 1, and $z^{\mathrm{SPCF}}$ from Definition 2.
-However, note that to compute the effect of the counterfactual treatments on the outcome, we need to perform counterfactual approximations in two steps, where the first step measures the counterfactual treatment and the second step measures the effect on the outcome.
-Extension for path-specific disparity measures.
-To analyze our path-specific measures in Definitions 2 and  4, we cannot apply the existing causal normalizing flows counterfactual estimation process directly since it does not work for multiple sequential interventions.
-Moreover, since we want to compute path-specific measures in non-linear settings, we cannot directly apply the existing linear approach in [8] either.
-Hence, we provide a novel procedure that extends the existing non-linear causal normalizing flows counterfactual approach to approximate path-specific counterfactuals.
-For computing $z^{\mathrm{SPCF}}$ in Definitions 2 and 4, we need to perform sequential interventions in the causal normalizing flows model following the causal ordering of the features to obtain:
-$\begin{split}&z^{\mathrm{SPCF}}=T^{-1}_{\psi}(\langle{u}^{CF}_{S},\hat{{u}}_{X}^{F},{u}^{F}_{Z},{u}^{F}_{Y}\rangle),\,\mathrm{where}\\ &{u}^{CF}_{S}=T_{\psi}({\operatorname{do}}(S\rightarrow(1-s^{F}))),\,\mathrm{and}\,\,\hat{{u}}_{X}^{F}=T_{\psi}(s^{CF},{\operatorname{do}}(X\rightarrow x^{F}))\end{split}$ (7)
-Hence, to compute $z^{\mathrm{SPCF}}$, we perform two interventions following the causal ordering.
-First we use ${\operatorname{do}}(S\rightarrow(1-s^{F}))$ on $\vartheta^{F}$ to get $\vartheta^{\mathrm{SCF}}$.
-Then, we use ${\operatorname{do}}(X\rightarrow x^{F})$ on $\vartheta^{\mathrm{SCF}}$, combine the different exogenous $u$ from the separate intervention steps and generate $z^{\mathrm{SPCF}}$.
-Further details and pseudocode regarding the different counterfactual computations using CNF can be found in Appendix B.
-[^2] Code can be accessed on GitHub.
-3.3.3 Assumptions and considerations
-While our implementation currently utilizes causal normalizing flows, our causal framework is inherently flexible, allowing for the seamless integration of future advancements in causal generative modeling to generate counterfactual quantities.
-However, examining the assumptions and certain considerations specific to the current implementation is crucial for understanding its practical implications.
-First, we highlight the causal assumptions central to our causal framework.
-Consistent with previous works in causal modeling [30, 62, 26] and causal fairness [35, 8, 31], we assume the causal data generation process has no hidden confounders.
-We base our analysis on the decision-maker’s perspective and assume our framework has access to all treatment decisions $Z$ and all covariates $X$ and sensitive $S$ used to assign $Z$.
-These considerations eliminate confounding between $Z,Y$ and $X,Z$.
-However, in many real-world scenarios, decision-makers may be unable to account for all $X$ that impact $Y$, leading to unseen confounding amongst $X,Y$.
-This hidden confounding does not impact total treatment disparity (TTD) but could influence direct treatment disparity (DTD), which relies on interventions on $X$.
-Alternatively, future work could consider an alternative external auditor perspective, in which case a lack of transparency on the treatment assignment procedure leads to not accounting for all $Z$ or all $X$ that were used to assign $Z$, thereby introducing confounding across $X,Z,Y$.
-While some studies [31, 7] have attempted to provide fairness estimation bounds under confounding, they require specific modeling considerations or do not account for multivariate, non-binary treatment decisions.
-Other works [68, 67, 45] have explored causal inference using instrumental and proxy variables to address hidden confounding.
-However, these methods only estimate population-level average effects and lack the decision-subject specificity needed to generate counterfactuals in the presence of hidden confounders.
-Therefore, further research is necessary to understand the impact of confounding on treatment fairness in real-world scenarios and to develop more robust estimation methods for generating counterfactuals in the presence of confounding.
-Beyond confounding, another causal assumption central to our framework and CNF is overlap [54].
-We assume that intervened values in the SCM are plausible and have non-zero support in the original causal data-generating process.
-However, this assumption is reasonable and holds in many real-world scenarios.
-For instance, the treatment $Z$ distributions across different sensitive $S$ for real-world datasets (Appendix C) show significant overlap.
-Second, to ensure accurate treatment fairness measures, it is essential to consider the accuracy of the generated counterfactual entities.
-Assuming underlying causal assumptions hold in a real-world scenario, [26] have already shown how causal normalizing flows can generate highly accurate counterfactuals.
-However, testing for the validity of the assumptions in real-world settings is challenging.
-Although we do not explicitly pursue this avenue in this work, our framework and setup offer a unique advantage: it allows decision-makers to assess and improve counterfactual quality through real-world studies.
-Specifically, our counterfactuals only modify treatments, which are entirely under the control of decision-makers.
-Hence, decision-makers can adjust treatment policies (e.g., adopting fair counterfactual treatments) and conduct real-world interventional studies on small subpopulations.
-This possibility enables decision-makers to conduct real-world analyses of counterfactual treatment accuracy, ultimately improving our framework’s applicability.
-4 Mitigating treatment discrimination in historical data
-In this section, we focus on settings where historical data serves a data-driven decision-making pipeline with one of the following purposes: i) automating the (non-binary) decision-making processes or ii) informing decision-making via risk score estimation.
-Unfortunately, as discussed in previous sections, historical data may have been collected under discriminatory treatment policies, where disparities in past treatments $Z$ led to undesired harm in the outcome $Y$.
-In such scenarios, we want to ensure that past discriminatory effects from treatment disparities do not propagate through automated decision-making processes and cause unfair future decisions.
-To enable future fair decisions, we propose an extension of our framework, providing an automated pre-processing procedure to transform biased historical datasets $\mathcal{D}$ to treatment-fair datasets $\mathcal{D}^{\mathrm{fair}}$.
-Our pre-processing approach helps prevent the propagation of past treatment discrimination in future automated systems and further improves them through fair risk score estimates.
-In the following sections, we assume an observed and historically collected unfair dataset, $\mathcal{D}=\{s^{F},x^{F},z^{F},y^{F}\}$, comprising a representative sample [^3] We assume the dataset consists of independent and identically distributed samples drawn from an underlying data-generation process.
-Relaxations of this assumption, such as algorithmic decision-making under selective labels, are discussed in [36].
-of the random variables $S$, $X$, $Z$, and $Y$, representing sensitive attributes, covariates, non-binary treatment decisions, and outcomes, respectively.
-4.1 Fairness in non-binary decision-making
-To automate decision-making processes, decision-makers often use historical data to train a binary classifier, $h_{y}:={S,\tilde{X}}\mapsto Y$, where $S$, $\tilde{X}$, and $Y$ represent, respectively, sensitive attributes, input features, and the target label.
-To the best of our knowledge, prior works assume that $\tilde{X}=\{X,Z\}$ contains the applicant covariates $X$ (e.g., income, savings, etc.) along with the non-binary treatment decisions $Z$ (e.g., loan terms).
-In this context, the objective is to achieve both accuracy and fairness by learning a function that produces predictions $\hat{Y}\in\{0,1\}$, which act as proxies for the true target labels $Y\in\{0,1\}$.
-However, if the decision-maker’s policy to assign treatments is unfair, the resulting automated decision-making process will also be unfair.
-We address this problem by extending our framework with an automated procedure to transform historically treatment-biased data.
-This procedure allows the decision-makers to assign fair treatments and ultimately make fair (non-binary) treatment decisions.
-The idea is that fair treatments should be assigned to lead to better, and thus fairer, outcomes $Y$.
-Predictive algorithms trained on these fair treatments and outcomes would lead to fairer subsequent decisions.
-More in detail, we first identify a fair treatment policy $\pi_{z}:=X\mapsto Z$.
-However, assigning non-binary treatments is often inherently unfair, complex, or unknown, making the selection of an appropriate $\pi(Z)$ that is optimal and fair for all stakeholders a challenging task.
-To address this issue, we propose an approach that ensures automated treatment assignments are non-discriminatory while improving decision outcomes for disadvantaged subpopulations.
-Specifically, when we observe treatment disparities lead to discriminatory effects on outcomes, we leverage our causal framework and the generative capacity of the estimated SCM to augment existing potentially unfair data to construct a fair dataset $\mathcal{D}^{\mathrm{fair}}$.
-In this automated transformation process, for applicants in the disadvantaged group $S^{F}=s^{-}$ who receive less favorable treatments, we propose intervening in their treatment decisions $Z$ to construct $\mathcal{D}^{\mathrm{fair}}$ as follows:
-$\begin{split}\mathcal{D}^{\mathrm{fair}}=\left\{s^{F},x^{F},z^{F},y^{F}\right\}_{s=s^{+}}\bigcup\left\{s^{F},x^{F},z^{\mathrm{SCF}},y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))\right\}_{s=s^{-}}\end{split}$ (8)
-where $z^{\mathrm{SCF}}$ follows Definition 1.
-For applicants in the original (unfair) dataset $\mathcal{D}$ with $S=s^{-}$, we intervene on $Z$ by replacing it with $z^{\mathrm{SCF}}$, derived from the advantaged group.
-Consequently, the intervened $Z$ and the resulting $Y$ are replaced with their counterfactual values, as specified in Definitions 1 and 3.
-This intervention on the disadvantaged subpopulation ensures that the revised treatment assignments $z^{\mathrm{SCF}}$ remove treatment discrimination and lead to improved decision outcomes, satisfying $P[Y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}))=1\mid S^{F}=s^{-}]\geq P[Y^{F}=1\mid S^{F}=s^{-}]$.
-This process ensures that disadvantaged applicants receive treatment decisions that eliminate bias while leading to fair decision outcomes.
-In summary, to automate fair non-binary decision-making processes based on an unfair dataset, one can use our causal framework to generate a fair dataset and then use it to learn a fair treatment policy $\pi_{z}$ and a fair binary classifier $h_{y}$.
-4.2 Fair risk score estimation
-Risk scores are a critical tool in algorithmic decision-making that decision-makers use to inform decisions and treatment policies [23, 53, 6].
-For example, in lending, banks rely on risk score estimates from third-party institutions such as FICO and SCHUFA [14, 64] to understand an applicant’s creditworthiness, helping them inform the loan decisions and justify interest rate variations [65, 38].
-Risk scores in decision-making are usually estimated by leveraging historically collected data $\mathcal{D}$ and past decision outcomes [40, 51].
-Unfortunately, potential disparities in prior treatment decisions may have a discriminatory influence in past data and outcomes.
-Hence, when estimating risk scores as the probability of outcome $y^{F}$ using past data covariates $x^{F}$ and past treatments $z^{F}$, the unfairness can propagate and lead to biased risk score estimates $R_{\mathrm{unfair}}$, often modeled as:
-$R_{\mathrm{unfair}}(x,z)={P}(y^{F}=0\mid X=x^{F},Z=z^{F}),$ (9)
-Importantly, these risk score estimations may remain biased even if we remove treatment decisions from the estimation process due to the causal effect of the treatment $Z$ in the outcome $Y$.
-Therefore, directly incorporating these risk score estimates as additional applicant features in decision-making will perpetuate or amplify unfairness.
-Risk score estimates should be free from potential past treatment discrimination to ensure decision-makers can make well-informed, fair decisions.
-For instance, in insurance underwriting, fair risk scores allow providers to offer premium rates, mitigate systemic inequities, and prevent perpetuating historical biases in insurance coverage decisions.
-To ensure fairness, we propose that risk scores designed to estimate an applicant’s inherent risk should exclude historical treatment decisions, as these factors go beyond an applicant’s control and may reflect past biases [33, 12].
-Hence, for an applicant, we formulate fair risk score estimates $R_{\mathrm{fair}}$ as:
-$\begin{split}R_{\text{fair}}(x,s)=\mathbb{E}_{z^{\prime}\sim\pi(Z)}\left[P\left(y^{CF}(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow z^{\prime}))=0\right)\right],\end{split}$ (10)
-where $\pi(Z)$ denotes an interventional distribution over treatments, which we use to marginalize out the impacts of treatments on the risk estimation.
-These fair risk scores, $R_{\text{fair}}$, can then be integrated into decision-making pipelines with covariates ($X$) to make future fair treatment decisions.
-The interventional treatment distribution $\pi(Z)$ in our risk score estimation ensures that $R_{\mathrm{fair}}$ is both fair and treatment-agnostic.
-Thus, $\pi(Z)$ cannot be any arbitrary distribution; instead, the treatments in $\pi(Z)$ should not systematically disadvantage any sensitive group through adverse downstream effects.
-As a simple approach, when we observe past treatment disparities are discriminatory, we set $\pi(Z)$ as the interventional treatment distribution of the group $S=s^{+}$ in $\mathcal{D}^{\mathrm{fair}}$.
-Since the group $S=s^{+}$ historically received favorable treatment assignments and outcomes, our choice of $\pi(Z)$ helps maximize positive outcomes for both groups and mitigates the effects of treatment discrimination.
-Moreover, since real-world scenarios show overlap across the factual treatments of the intervened groups (Appendix C), our selection of $\pi(Z)$ results in a distribution over realistic treatment values.
-We leave the challenging exploration of alternative fair interventional distributions that are globally fair policies for future work.
-5 Use case : Lending decisions
-In this section, we evaluate the effectiveness of our proposed causal framework for measuring and mitigating non-binary treatment discrimination in decision-making processes.
-We use four real-world datasets across diverse lending scenarios spanning different societal contexts, geographic regions, and loan structures.
-Given the comprehensive practical foundation for evaluation, our framework shows how historical data can contain treatment disparities and its potentially discriminatory impact on decision outcomes.
-Additionally, we demonstrate the application of our mitigation strategies to address this treatment discrimination.
-By focusing on two scenarios, our results explore potential historical biases in treatment decisions, particularly the loan terms set by banks.
-  • Agreement of the decision-making process as label (S.1): The ground-truth label $Y$ represents the immediate outcome of the decision-making process, with $Y=1$ indicating a positive agreement on treatment terms between the decision-maker and the applicant.
-For instance, in lending scenarios, $Y=1$ means the bank approved the loan, and the applicant accepted the terms.
-We analyze this scenario using the HMDA (Home Mortgage Disclosure Act) dataset, focusing on data from New York and Texas for 2017 [9, 18].
-  • Decision outcome as label (S.2): The ground-truth label $Y$ represents the downstream outcome of a decision, indicating whether it was right ($Y=1$) or wrong ($Y=0$).
-For example, in lending scenarios, $Y=1$ means the loan was fully repaid, while $Y=0$ indicates default.
-We analyze this scenario using the Home Credit[48] and German Credit [22] datasets as case studies.
-First, we use the novel fairness metrics from our framework in Section 3 to quantify treatment disparities in historical lending data.
-Through our path-specific counterfactual analyses, we demonstrate how sensitive attributes influence treatment decisions both directly and totally and subsequently affect the outcomes, potentially adversely causing discrimination.
-Next, we mitigate the discrimination using two strategies: (i) generating treatment-fair counterfactual datasets to automate unbiased decision-making and (ii) adjusting risk score estimation to reduce the impact of unfair treatments.
-We detail the experimental setup and datasets in Appendix C and provide additional results on synthetic datasets in Appendix D to analyze the accuracy of our causal estimation-based framework.
-5.1 Measuring treatment disparities
-5.1.1 S.1: Agreement of the decision-making process as label.
-Table 2: Measuring treatment disparities in S.1 and outcome effects.
-$Y^{F}=0$ indicates a label change from negative to positive agreement, $Y^{F}=1$ the reverse.
-Results show median U.S.
-Dollars (\$currency-dollar\\$\$).
-[…]
-Amount (×$\times 10^{3}$\$) $Y^{F}=0$ $Y^{F}=1$
-HMDA-NY Female $\rightarrow$ Male TTD(-E) $+29.46$ $2.00$ $0.40$
-DTD(-E) $+6.65$ $2.00$ $0.20$
-Male $\rightarrow$ Female TTD(-E) $-29.93$ $4.00$ $0.16$
-DTD(-E) $-5.88$ $2.00$ $0.14$
-HMDA-TX Female $\rightarrow$ Male TTD(-E) $+19.43$ $2.00$ $0.28$
-DTD(-E) $+2.67$ $2.00$ $0.13$
-Male $\rightarrow$ Female TTD(-E) $-19.78$ $3.00$ $0.18$
-DTD(-E) $-3.65$ $3.00$ $0.10$
-In both HMDA datasets, we follow the identification and pre-processing procedures for sensitive attributes, covariates, and treatment decisions outlined in [10].
-We consider “gender” a sensitive attribute, while loan amount and preapproval status are treatment decisions within the bank’s decision-making authority.
-We summarize the results of the treatment discrimination analyses in Table 2.
-Results in HMDA.
-We analyze the treatment disparity measures ($\operatorname{TTD}$ and $\operatorname{DTD}$) for factual males and females across the two states.
-Our analyses show that female applicants, if treated like their male counterfactuals, would receive a substantial increase in loan amounts, e.g., in NY, approximately \$currency-dollar\\$\$29K more based on $\operatorname{TTD}$ (and \$currency-dollar\\$\$6K based on $\operatorname{DTD}$) on average.
-Conversely, if male applicants were treated like females, their loan amounts would decrease by a similar amount.
-Examining the effects on the outcome agreement rates, we observe similar $\operatorname{TTD-E}$ and $\operatorname{DTD-E}$ values for both factual females and males, with agreement rates increasing slightly more when males are treated as females.
-The HMDA-TX dataset shows similar, though less pronounced, treatment disparities compared to the NY dataset.
-Takeaway.
-Although we find no significant differences in loan agreement rates between genders, banks tend to adopt a more conservative approach in determining treatments for female applicants, suggesting potential treatment discrimination.
-Our findings show that offering higher loan amounts to female applicants would not negatively impact their agreement outcomes, casting doubt on the rationale behind such differential treatment.
-Furthermore, the observed disparities seem primarily mediated through applicant covariates, with minimal direct disparity.
-The lack of direct disparities and minimal impact on the immediate outcome (agreement rates) raises important questions about the root causes of the differential treatments.
-Specifically, one might ask: Is the disparity in loan terms against female applicants justified by the downstream outcome, such as repayment behavior, and thus justified by business necessity?
-This question forms the basis for our analyses in the second scenario, where we investigate the link between treatment disparities, repayment outcomes, and discrimination.
-Table 3: Measuring treatment disparities in S.2 and outcome effects.
-$Y^{F}=0$ indicates a label change from loan defaulted to repaid, $Y^{F}=1$ the reverse.
-Results show median INR (Indian Rupees) and DM (Deutsche Marks).
-[…]
-Annuity (INR) Amount (INR) $Y^{F}=0$ $Y^{F}=1$
-Home Credit Female $\rightarrow$ Male TTD(-E) $+1008.12$ $+7885.72$ $0.24$ $0.30$
-DTD(-E) $+715.91$ $-816.94$ $0.27$ $0.00$
-Male $\rightarrow$ Female TTD(-E) $-942.92$ $-6659.68$ $2.30$ $0.00$
-DTD(-E) $-594.56$ $+758.46$ $1.30$ $0.00$
-Duration (months) Amount (DM) $Y^{F}=0$ $Y^{F}=1$
-German Credit Female $\rightarrow$ Male TTD(-E) $+1.29$ $+253.34$ $2.00$ $3.50$
-DTD(-E) $+0.46$ $+164.66$ $2.00$ $3.00$
-Male $\rightarrow$ Female TTD(-E) $-0.79$ $-243.11$ $8.00$ $1.00$
-DTD(-E) $-0.59$ $-212.39$ $8.00$ $0.60$
-5.1.2 S.2: Decision outcome as label.
-In the Home Credit dataset, we analyze two treatment decisions: the loan amount and annuity amount, while in the German Credit dataset, we examine three treatment decisions: loan amount, loan duration, and installment rate.
-In both datasets, we use “gender” as the sensitive attribute to evaluate potential disparities in treatment decisions.
-We summarize our treatment disparities results in Table 3.
-Results in Home Credit.
-Our analyses show that if females were treated like their male counterfactuals, they would have received higher annuities and loan amounts.
-At the same time, the reverse pattern would apply to male applicants when treated as females.
-Interestingly, we see contrasting trends between total and direct disparities regarding loan amounts.
-Total disparity indicates conservative lending practices where female applicants are granted lower loan amounts and annuities.
-In contrast, direct disparity, attributed solely to gender, shows that female applicants receive lower loan amounts but higher annuities, suggesting elevated interest rates.
-These findings suggest that while banks may have implemented some affirmative treatment policies, disparities mediated through covariates influence lending decisions significantly.
-Importantly, we see slight downstream effects on repayment outcomes across groups, indicating that outcome differences do not justify the observed treatment disparities against female applicants, particularly in terms of $\operatorname{DTD}$.
-Hence, in this context, the unjustified treatment disparities may be considered discriminatory.
-Results in German Credit.
-We observe a similar trend with the bank adopting more conservative practices toward female applicants.
-If females were treated like males, they would receive longer loan durations and higher loan amounts, with the reverse pattern observed for males.
-Significant disparity is seen in loan amounts, driven mainly by direct effects.
-This observation suggests that applicant covariates do not fully explain treatment disparities, and gender may have a direct influence, exposing a potential problematic mechanism.
-Interestingly, there is no disparity in installment rate treatments regarding $\operatorname{TTD}$ or $\operatorname{DTD}$, so we exclude this feature from Table 3 for clarity.
-When examining the impact of treatment disparities on outcomes, a key observation arises: treating males like their female counterparts would result in 8.0% of defaulting males repaying their loans.
-In contrast, only 1.0% of male repayers would default under this counterfactual treatment.
-Hence, our counterfactual analysis shows that the treatment disparities are not justified by outcomes, raising concern on potential treatment discrimination.
-Takeaway.
-Our results in Scenario S.2 show that, across different datasets, female applicants experience negative treatment disparities that cannot be explained by downstream outcome differences.
-Furthermore, while males receive perceived positive treatment disparity, they are ultimately negatively impacted and discriminated by its downstream effects on outcomes.
-Notably, our findings align with similar insights observed for German Credit in a simpler linear setting [28], reinforcing the validity of our results.
-Despite appearing favorable, the loan terms offered to certain males may increase their default risks and lower creditworthiness over time.
-Applying the treatment decisions of their female counterparts to male applicants could mitigate this discrimination and improve repayment outcomes.
-Next, we further explore the mitigation of this discrimination in treatment decisions.
-5.2 Mitigating treatment discrimination
-Our results in Table 3 show that banks exhibit negative treatment disparity against female applicants.
-However, males are potentially discriminated against since they experience a negative downstream effect on repayment ability, and treating them as females improves repayment performance.
-Hence, treating all applicants conservatively as females may enhance the overall repayment ability of borrowers, benefiting both the banks and the applicants.
-5.2.1 Fair non-binary decision-making
-Following Section 4.1, we examine the creation of a treatment-fair dataset aimed at automating future fair non-binary decision-making tasks.
-Based on our Scenario S.2 analyses, we identify males as the group needing adjusted treatment to generate our fair dataset, $\mathcal{D}^{\mathrm{fair}}$.
-To understand the impact of $\mathcal{D}^{\mathrm{fair}}$, we aim to compare the fair dataset with the original (unfair) dataset on two key metrics: the bank’s Loss Given Default ($\operatorname{LGD}$[63] and the applicant’s Expected Simple Interest ($\operatorname{ESI}$[59].
-Detailed information on the metric calculations and dataset-specific implementations can be found in Appendix C.2.
-Takeaway.
-The results for Home Credit and German Credit, summarized in Table 4, show that adjusting treatment decisions for male applicants addresses treatment biases and yields tangible financial benefits for both stakeholders.
-The adjusted treatments result in an improved $\operatorname{LGD}$ for banks, indicating reduced financial risk.
-Simultaneously, male borrowers can reduce loan costs owing to the lower resultant $\operatorname{ESI}$.
-These findings underscore the potentially broad impact of fairer treatment practices to improve the conditions for all stakeholders.
-5.2.2 Fair risk score estimation
-Following Section 4.2, we now examine how we can more fairly inform decision-making processes by estimating treatment-fair risk scores.
-Based on insights from scenario S.2, we focus on male applicants and apply different treatment distributions $\pi(Z)$ to generate fair risk estimates as defined in Equation 10.
-We approximate $\pi(Z)$ for males under two conditions: (i) a fair interventional distribution, derived from the empirical distribution of $Z$ conditioned on female applicants to produce “fair” risk scores, and (ii) the factual “unfair” distribution, based on the empirical distribution of $Z$ conditioned on male applicants.
-For female applicants, $\pi(Z)$ is approximated using the empirical distribution of $Z$ conditioned on female applicants.
-We apply these policies to the Home Credit and German Credit datasets and compare the empirical cumulative distribution functions (CDFs) of the resulting risk scores for male and female applicants in Figure 2.
-Takeaway.
-The figure shows that, while males generally exhibit lower risk scores than females for German Credit, the pattern reverses for Home Credit.
-However, historical treatment practices have led to an unfair overestimation of males’ risk scores, especially in German Credit.
-This bias is corrected when the risk scores are recalculated under the fair interventional distribution, adjusting for potentially unfair treatments.
-Integrating these recalibrated, fair risk scores into decision-making pipelines mitigates the residual impacts of historical treatment biases, paving the way for more fair decision-making processes.
-Table 4: Fairness in non-binary decision-making in S.2.
-Comparison of bank’s LGD and applicant’s ESI (both: lower better) under different policies.
-$\mathcal{D}$ is factual data, $\mathcal{D}^{\mathrm{fair}}$ data is obtained by treating all applicants as females.
-Dataset Sensitive Group LGD (INR/DM) ESI (INR/DM)
-Home Credit Females $586692.67$ $290208.73$
-Males ($\mathcal{D}$) $576590.35$ $422114.31$
-Males ($\mathcal{D}^{\mathrm{fair}}$) 568157.43 391110.73
-German Credit Females $1220.51$ $312.13$
-Males ($\mathcal{D}$) $1163.88$ $495.65$
-Males ($\mathcal{D}^{\mathrm{fair}}$) 940.03 393.86
-(a) Home Credit
-(b) German Credit
-Figure 2: Fair risk score estimation in S.2.
-Comparison of risk score estimates for females and males under various treatment distributions $\pi(Z)$.
-Unfair estimates are based on factual distribution, while fair estimates use interventional treatment distribution where all applicants are treated as females.
-6 Related work
-Fairness in lending.
-Fairness analyses in automated lending decisions have focused mainly on statistical methods [21, 24, 37], emphasizing metrics like equality of opportunity [21], as noted in surveys [34, 19].
-However, these approaches often overlook biases embedded in the ground-truth labels, which stem from unfair treatment during loan processing.
-While prior work has separately examined treatment discrimination [60, 1] and biased outcomes in lending decisions [50, 3, 47], the causal link between biased treatment and its impact on outcomes remains underexplored.
-Our novel framework bridges this gap and jointly analyzes treatment discrimination and its causal effects on lending outcomes.
-Causal fairness notions in algorithmic predictive systems.
-Causal fairness in algorithmic decision-making examines the causal influence of sensitive features on decision labels and outcomes.
-Counterfactual fairness [35] introduced the study of the total causal effect of sensitive attributes, while subsequent work explored effects along specific pathways, distinguishing fair from unfair pathways [52, 8].
-Later frameworks expanded to include multi-level attributes [43] and focused on fairness in decision outcomes [57].
-However, these approaches primarily address biases in binary labels or outcomes, neglecting complex mechanisms like treatment decisions (e.g., loan terms) controlled by decision-makers.
-Although [39] introduced a causal framework for treatment fairness, it simplified multi-faceted treatments into binary variables, limiting its granularity.
-Our work advances this by incorporating the decision-maker’s perspective to analyze treatment decisions and their causal effects.
-We develop a novel framework for fine-grained fairness analyses using path-specific counterfactuals [8].
-Additionally, following [41], we enhance counterfactual accuracy in societal contexts through state-of-the-art causal normalizing flows [26], surpassing previous variational methods [8, 39].
-Multi-stage selection in decision-making.
-Multi-stage selection problems [16, 27] explore automating decision-making with binary decisions across stages, revealing additional candidate covariates for subsequent decisions.
-Concerning this line of work, our causal framework addresses a fundamentally different context governed by different mechanisms.
-Our work focuses on a single-stage decision process where binary decisions (e.g., loan approval) are made in one step, with treatments directly controlled by the decision-maker.
-In contrast, multi-stage problems rely on observing additional applicant covariates.
-We complement existing studies by analyzing fairness in automated decision-making, specifically the fairness of assigned multivariate, non-binary treatments and their causal impact on decision outcomes.
-7 Discussion
-This work highlights the importance of incorporating non-binary treatment decisions in fairness analyses for automating decision-making systems.
-We proposed a novel causal framework to measure disparities in treatment decisions across sensitive groups and its impact on downstream outcomes.
-We also demonstrate how our framework can help mitigate unfairness and promote fairer decisions.
-However, ensuring its practical and reliable implementation requires addressing and discussing several critical challenges.
-Ensuring robust causal frameworks for real-world scenarios.
-Our analyses assume access to ground-truth labels for outcomes and the absence of hidden confounders in the causal generative process.
-However, these assumptions often fail in real-world settings and specific use cases (e.g., decision-maker vs auditor as user).
-Hidden confounders can bias treatment fairness estimates while using proxy labels for unobservable outcomes [66], which introduces additional complexities.
-Addressing these challenges by accounting for hidden confounders is a promising direction for future research.
-Designing mitigation strategies for diverse stakeholders.
-Similar to pre-processing approaches in predictive fairness, our pre-processing mitigation methods for treatment discrimination prove empirically effective but they do not guarantee optimality.
-Moreover, these approaches may not fully align with the diverse and often conflicting objectives of all stakeholders, raising a crucial question: What defines a fair treatment policy, and how should it be formulated?
-For instance, banks prioritize repayment probability in loan approval, while applicants seek lower interest rates [55].
-Addressing fairness in such settings requires a holistic approach that considers all stakeholders’ utilities and relies on carefully crafted utility functions to capture these objectives reliably.
-Hence, future work should compare our methods to existing simple treatment interventions [12] while also focusing on novel in-processing methods to develop context-aware policies that can balance fairness and diverse stakeholder utilities.
-Improving data collection to enable treatment discrimination analysis.
-As shown in Table 1, while our framework focuses on loan datasets, it is broadly applicable across decision-making domains.
-However, its practical application is constrained by significant data limitations.
-Many datasets fail to distinguish covariates $X$ from treatments $Z$, omit critical treatment details (e.g., COMPAS [5] lacks bail terms, lending datasets exclude loan interest rates), or oversimplify multifaceted treatments (e.g., IHDP [39] reduces childcare factors to binary).
-Moreover, most datasets are collected under selective labeling settings [36, 32], where ground-truth outcomes are observed only for cases with a favorable agreement between the decision-maker and the decision-subject.
-This setting is problematic, as it overlooks unknown agreement-phase discrimination that may have occurred during the agreement process.
-These gaps emphasize the need for more transparent and detailed data collection practices that explicitly capture treatment decisions, including those made during the agreement phase, to enable comprehensive and fair analyses.
-Concluding remarks.
-Our work introduces a paradigm shift by integrating fairness considerations into non-binary treatments conditional on positive binary decisions (e.g., granting a loan).
-By proposing treatment disparity measures, we complement existing literature to advance a more holistic analysis of fairness.
-Our novel perspectives and empirical analyses demonstrate that fairness in algorithmic decision-making remains unresolved, necessitating interdisciplinary socio-technical approaches to address the needs of diverse stakeholders.
-By accounting for the complexity of decisions and their impact on stakeholders’ utility, our work signifies a major step toward aligning algorithmic systems with societal expectations.
-Our approach also paves the way for future research to adapt established predictive fairness methods to redefine and mitigate treatment unfairness.
-Acknowledgements
-We thank Adrián Javaloy for providing invaluable guidance regarding incorporating the causal generative models and feedback on our written manuscript.
-This work has been funded by the European Union (ERC-2021-STG, SAML, 101040177).
-However, the views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
-Neither the European Union nor the granting authority can be held responsible.
-Author contributions
-Conceptualization: IV, DK; Formalization: All authors; Methodology: AM, IV; Related work: AM, DK; Data curation: DK, KG; Code implementation: AM, KG; Experimental evaluation: All authors; Writing - Original draft: All authors with AM as main contributor to Sections 2, 3, 6, 7, Appendix B), DK to Sections 56,Appendices CD; Writing - Review and editing: All authors.; Supervision: IV.
-Appendix A Causal normalizing flows
-Causal normalizing flows [26] enable causal inference from observational data under minor assumptions like invertible and differentiable structural equations, acyclic (known) graphs $\mathcal{G}$, and causal sufficiency.
-Under these assumptions, causal flows represent a SCM $\mathcal{M}$ as Triangular Monotonic Increasing (TMI) maps and model them using single, invertible normalizing flows models $T_{\psi}$ with parameters $\psi$.
-Note that causal flows can handle partial causal graphs, in which each random variable in the graph may be multivariate without explicitly modelling the causal relationships across its dimensions.
-Thus, it provides an ideal practical framework to estimate our measures of fairness with respect to treatment decisions and the downstream effects from available (potentially unfair) data collected from a decision-making process.
-A.1 Hyper-Parameter tuning and training of causal normalising flows
-For the causal normalizing flows [26], we use the original work as inspiration and tune for the flow models’ inner dimensionality and learning rate Table 5.
-All other parameters are fixed, which are as follows: Adam optimizer with plateau scheduler, activation as ELU, layer type was Neural Spline Flow (NSF) and all the experiments were run for a maximum of $1000$ epochs.
-For other parameters, we refer to the GitHub code  [^4] https://github.com/psanch21/causal-flows.
-The number of nodes is specified in Table 6.
-An example of causal graph input to causal normalizing flows is given in Figure 3.
-Table 5: Hyperparameters for causal Flows. The log-prob of the observational data distribution estimate was used to select the best hyperparameters.
-Dataset dim-inner lr log-prob
-HMDA-NY $32\times 32$ $10^{-2}$61.67
-HMDA-TX $32\times 32$ $10^{-2}$63.49
-Home Credit $64$ $10^{-2}$41.87
-German Credit $64$ $10^{-3}$14.68
-Synthetic Loan $16\times 16\times 16\times 16$ $10^{-3}$16.51
-Table 6: Number of root nodes, covariates, and treatment variables for training causal normalizing flows model. The dimensionality of the outcome label is 1.
-Dataset #rootnodes #covariates #treatment
-HMDA-NY 2 12 2
-HMDA-TX 2 12 2
-Home Credit 2 9 2
-German Credit 2 15 3
-Synthetic Loan 2 3 2
-A.2 Hardware
-All experiments of hyper-parameter tuning and training of causal normalizing flows are done on an Ubuntu workstation.
-The code primarily uses Pytorch CPU and Python.
-We set the seed for PyTorch and Python to a fixed number and ran all experiments.
-A fixed seed is also used to split datasets into train, validation, and test sets - 80%, 10%, 10%.
-This seed is fed as input in a parameter file to train and run evaluations.
-AG
-[…]
-hammer
-EISLDYSensitive $S$Outcome $Y$Block $X$Block $Z$
-Figure 3: Causal graph for synthetic loan dataset introduced in Appendix D.
-While training the causal normalising flows, covariates $X$ and treatment decisions $Z$ are treated as multivariate blocks without any explicit assumptions on the cause-effect directionality within each block.
-Here, we have multiple sensitive attributes (independent root nodes) but for this work, we concentrate on intervening on just one sensitive attribute (“gender”).
-A similar analysis can be done by intervening on the attribute “age”.
-Appendix B Counterfactual computation pseudocodes
+    Variance:
+    $\mathcal{A}_{\text{VAR}}(s^{F})=\frac{1}{(|\mathcal{S}|-1)|\mathcal{S}|/2}\sum_{s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}}\Delta(z^{\mathrm{SCF}}_{s^{\prime}},z^{F})^{2}$
+Definition 2 can be extended similarly by computing path-specific counterfactuals $z^{\mathrm{SPCF}}$ instead of $z^{\mathrm{SCF}}$ for all $\mathcal{S}\setminus\{s^{F}\}$ and using the proposed aggregation methods.
+Outcome disparities.
+We first collect the flips in the outcome label $Y$ caused by the counterfactual treatments:
+$Y_{\text{flip}}=\Big\{\mathbb{I}[y^{CF}({\operatorname{do}}(Z\rightarrow z^{\mathrm{SCF}}_{s^{\prime}}))\neq y^{F}]\;\mid\;s^{\prime}\in\mathcal{S}\setminus\{s^{F}\}\Big\}.$
+We then apply an outcome-level aggregation operator $\mathcal{A}_{Y}$ over these indicators to summarize the disparity for the individual, and finally take the expectation over the dataset:
+$\operatorname{TTD\text{-}E}_{\text{M}}(s^{F},y^{F})=\mathbb{E}_{\mathcal{D}}\Big[\mathcal{A}_{Y}(Y_{\text{flip}})\Big].$ (10)
+For the worst-case outcome disparity, $\mathcal{A}_{Y}$ is the identity applied to the indicator corresponding to the treatment with the maximum $\Delta$ (corresponding to $\operatorname{arg\,max}$ over the counterfactual treatment assignments).
+For other aggregations like average or variance in counterfactual outcomes, $\mathcal{A}_{Y}$ would be the corresponding aggregation function ($\operatorname{mean}$ or $\operatorname{Var}$).
+One can use this idea to extend Definition 4 similarly to compute the direct effects across all sensitive values $\mathcal{S}$.
+Intersectionality.
+As discussed, our framework supports analysis involving multiple sensitive attributes.
+However, sensitive attributes may not always act independently, hence their effects can interact in additive, non-additive, or even compounded ways, influencing treatment and outcome disparities in complex patterns (Crenshaw 2013; Kanubala and Valera 2025; Xenidis 2018; Makkonen 2002).
+In particular, disparities arising from intersectionality go beyond simple multi-sensitive attribute analysis.
+Intersectionality involves interdependent and interlocking systems of social influence and oppression, where the combined effect of attributes cannot be fully decomposed into separate or additive components (Crenshaw 2013).
+As such, capturing such intersectional sources of unfairness requires richer causal representations and remains an important direction for future work.
+(a) German Credit
+(b) Home Credit
+Figure 4: Treatment decision feature distributions across social groups show overlap, indicating it as a reasonable causal assumption for real-world settings.
+B.4 Extended Discussion on Assumptions, Limitations, and Practical Considerations
+We provide an extended discussion on the implicit assumptions of our existing causal framework and practical considerations for deployment.
+Partial causal knowledge.
+Our approach requires only partial knowledge of the underlying causal mechanisms—an assumption that is significantly weaker than that of many existing causal fairness methods (Kusner et al.
+2017; Chiappa 2019), which typically assume full access to the causal graph.
+Specifically, we assume only the causal ordering among $S$, $X$, $Z$, and $Y$, as depicted in Figure 2.
+Both $X$ and $Z$ are treated as feature blocks, and we do not require knowledge of the internal causal ordering among covariates or treatment components.
+By leveraging CNFs to learn causal relations directly from observed data, our framework preserves theoretical guarantees even under partial causal knowledge, thereby improving its practical applicability.
+No hidden confounders.
+Consistent with prior work in causal inference (Khemakhem et al.
+2021; Sánchez-Martin, Rateike, and Valera 2022; Javaloy, Sánchez-Martín, and Valera 2024) and causal fairness (Kusner et al.
+2017; Chiappa 2019), we assume a data-generating process with no hidden confounders.
+From the decision-maker’s perspective, this corresponds to having complete access to the treatments $Z$, covariates $X$, and sensitive attributes $S$ used to assign $Z$, thereby removing confounding between $(Z,Y)$ and $(X,Z)$.
+In practice, however, decision-makers may lack full knowledge of all relevant covariates $X$ influencing $Y$, introducing hidden confounding between $X$ and $Y$.
+Such confounding does not affect our metrics for total disparities (Definitions 1 and 3) but can bias our direct disparity measures (Definitions 2 and 4), which rely on interventions on $X$.
+External auditing and confounding.
+Under an external auditor perspective, transparency in treatment assignment may be limited, and some components of $Z$ or the covariates $X$ used to determine $Z$ might be unobserved, introducing confounding among $X$, $Z$, and $Y$.
+Existing studies (Kilbertus et al.
+2020a; Byun et al.
+2024) provide fairness bounds under confounding, but they rely on strong assumptions or do not accommodate multivariate, non-binary treatments.
+Other approaches (Xu et al.
+2021; Wang and Blei 2021; Miao et al.
+2023) employ instrumental or proxy variables to mitigate hidden confounding, yet they are restricted to population-level effects and lack the individual-level specificity required for counterfactual analysis.
+This underscores the need for further research on fairness under confounding and the development of robust, individualized counterfactual methods.
+Incorporating recent techniques that enable reliable causal inference under specific types of confounding (Almodóvar et al.
+2025) into our framework would be a promising direction for future work.
+Overlap.
+Beyond confounding, another key assumption underlying our framework and CNFs is overlap (Oberst et al.
+2020).
+We assume that the intervened values in the SCM have non-zero support within the original data-generating process.
+This assumption generally holds in practice.
+For instance, our considered real-world lending datasets show substantial overlap in treatment distributions across sensitive groups.
+As seen in Figure 4, we have significant overlap between the distributions of treatment decisions for female and male applicants.
+Hence, assigning treatment decisions of the sensitive counterfactual is a realistic consideration for these datasets.
+Nevertheless, certain real-world contexts may exhibit lack of overlap between treatment distributions, for which extending our causal framework remains an important avenue for future research.
+Counterfactual validity.
+Accurate fairness evaluation depends on the fidelity of generated counterfactuals.
+Under valid causal assumptions, CNFs can produce highly accurate counterfactuals (Javaloy, Sánchez-Martín, and Valera 2024).
+However, validating these assumptions in real-world environments remains challenging.
+Our framework offers a practical advantage in this regard: since counterfactuals in our setting involve only treatment interventions — which are under the decision-maker’s control — they can be empirically evaluated via interventional studies.
+By deploying counterfactual treatment policies to specific subpopulations and measuring the resulting average effects, practitioners can assess and refine the quality of counterfactuals and the causal framework in practice.
+Mitigation approaches.
+Our mitigation strategy focuses on pre-processing to address treatment unfairness.
+Such techniques have proven effective in binary predictive fairness (Kamiran and Calders 2012) and provide a natural foundation for mitigating treatment disparities.
+However, pre-processing methods may not suit all scenarios and often lack theoretical guarantees.
+Future work should explore in-processing strategies capable of jointly optimizing fairness and utility objectives.
+Designing such approaches is non-trivial, as fair policy learning must account for binary decision fairness, non-binary treatment fairness, and stakeholder utility optimization.
+Achieving these objectives likely requires multi-objective optimization (Padh et al.
+2021) for holistic fairness.
+Moreover, since treatment decisions and outcomes are typically conditioned on prior binary decisions (e.g., loan approvals), data may exhibit selection bias (Lakkaraju et al.
+2017).
+Addressing this challenge may require online learning or policy adaptation methods (Kilbertus et al.
+2020b; Rateike et al.
+2022).
+Finally, future work should also assess whether in-processing methods can fully meet the multi-dimensional fairness requirements in decision-making (Favier and Calders 2025), or whether entirely new paradigms are needed.
+B.5 Defining Stakeholder Losses
+Here, we define the stakeholder loss functions we utilize to measure the utilities of decision-makers and decision-subjects for our mitigation strategy.
+Loss Given Default (LGD):
+We estimate the loss incurred by the bank in the event an applicant defaults.
+Quantifying the loss will require additional analysis of other variables, such as the value of collateral, instalments payments etc (Investopedia 2024).
+There are several different ways to estimate $LGD$; however, a common approach considers the exposure at default and the recovery rate which is:
+LGD $=(1-\text{recovery rate})\times\text{exposure at default (EAD)}]$ (11)
+In our estimation, we assume that EAD is equivalent to the credit amount loaned out.
+EAD is a crucial metric that estimates the total loss a bank is exposed to when a loan defaults.
+However, it is important to note that the calculation of EAD can be more complex, as it may involve factors beyond the simple loan amount, such as commitment details (Risk 2012).
+The recovery rate, on the other hand, is a risk-adjusted measure that reflects the portion of the loan that the bank expects to recover after default.
+In our analysis, we considered it as the outcome of the loan (i.e., the actual repayment).
+Finally, we formalize the expected loss given default for the bank as:
+LGD $=\mathbb{E}_{\mathcal{D}^{\prime}}[(1-y)\times\text{amount}]$ (12)
+where $y$ represents the applicant’s repayment probability.
+Expected Simple Interest (ESI):
+We estimate the loss of the individuals using the simple interest formula:
+$\textbf{SI}=\frac{(\text{amount}\times\text{rate}\times\text{time})}{100}$
+For the Home Credit dataset, we used slightly modified formulae to estimate the ESI for Home Credit, as the dataset did not include information on the rate and time, making ESI estimation less straightforward.
+We estimate ESI using annuity and loan amount as follows: we assume duration=15 years.
+Annuity can be written as :
+$\text{annuity}=\frac{(\text{amount + interest})}{\text{total durations(months)}}=\frac{(\text{amount + interest})}{12\times\text{duration}}$
+Using this the simple interest can be calculated as :
+$\textbf{ESI}\text{(Home Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}[\text{annuity}\times 12\times\text{duration}-\text{amount}]$ (13)
+In our formulae for the German Credit dataset, the loan duration is in months and we assume in our implementation that everyone was given the same interest rate of 10%, hence enabling us to rewrite the ESI as follows:
+$\textbf{ESI}\text{(German Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}\left[y\times\text{amount}\times\frac{\text{rate}}{100}\times\frac{\text{time}}{12}\right]$ (14)
+B.6 Counterfactual Computation Pseudocodes
 This section provides the different function pseudocode for computing the various counterfactual and path-specific counterfactual entities using the causal normalizing flows.
@@
 Moreover, note that for our default consideration of binary sensitive attributes, $s^{CF}$ can be considered as $(1-s^{F})$ following the notation in the main paper.
-B.1 Generating sensitive counterfactual
-This function computes the sensitive counterfactual following the causal relationships when ${\operatorname{do}}(S\rightarrow s^{CF})$.
+Generating sensitive counterfactual
+Algorithm 1 computes the sensitive counterfactual following the causal relationships when ${\operatorname{do}}(S\rightarrow s^{CF})$.
 For this computation, we perform the single $\operatorname{do}$} operation on $S$ as ${\operatorname{do}}(S\rightarrow s^{CF})$.
 Algorithm 1 compute_sensitive_counterfactual
-Input: Factual data $D^{F}$
+Input: Factual data $\mathcal{D}^{F}$
 Parameters: Flows model, data preparator, sensitive index $i_{S}$ (default: 0)
-Output: Sensitive counterfactual data $D^{SCF}({\operatorname{do}}(S\rightarrow s^{CF}))$
-1:Get sensitive feature values $s^{F}$ from $D^{F}$.
+Output: Sensitive counterfactual data $\mathcal{D}^{SCF}({\operatorname{do}}(S\rightarrow s^{CF}))$
+1:Get sensitive feature values $s^{F}$ from $\mathcal{D}^{F}$.
 2:Get sensitive counterfactual feature value $s^{SCF}$ for intervention.
 3:Get data scaler from preparator.
-4:$D^{SCF}$leftarrow model.compute_CF ($D^{F}$, $i_{S}$, $s^{SCF}$, scaler)
-5:return $D^{SCF}$. $\triangleright$ clamp and floor categorical features before saving.
-B.2 Computing downstream effect of counterfactual treatments
-This function helps to compute the downstream effect on $Y$ if we perform ${\operatorname{do}}(Z\rightarrow\hat{z})$, where $\hat{z}$ is some interventional value.
-For example, we use this function to :
-  • understand the downstream effect of Total Treatment Discrimination (TTD) and
-  • understand the downstream effect of Direct Treatment Discrimination (DTD).
+[…]
+5:return $\mathcal{D}^{SCF}$. $\triangleright$ clamp and floor categorical features before saving.
+Computing downstream effect of counterfactual treatments
+[…]
 Algorithm 2 compute_path_specific_downstream_
 effect_treatment_intervention
-Input: Factual data $D^{F}$, counterfactual data $\hat{D}$
+Input: Factual data $\mathcal{D}^{F}$, counterfactual data $\hat{\mathcal{D}}$
 Parameters: Flows model, data preparator
-Output: $D(s^{F},x^{F},{\operatorname{do}}({Z\rightarrow\hat{z}}))$
-1:Get treatment indices $i_{\{T\}}$ from preparator
-2:Get values $\hat{z}$ from $\hat{D}$ using $i_{\{T\}}$
+Output: $)\mathcal{D}(s^{F},x^{F},{\operatorname{do}}({Z\rightarrow\hat{z}}$))
+1:Get treatment indices $i_{\{Z\}}$ from preparator
+2:Get values $\hat{z}$ from $\hat{\mathcal{D}}$ using $i_{\{Z\}}$
 3:Get scaler from preparator
 4:Clamp and floor $\hat{z}$ categorical features to proper value ranges
-5:$D^{F}({\operatorname{do}}(Z\rightarrow\hat{z}))$ = model.compute_CF($D^{F}$, $i_{\{T\}}$, $\hat{z}$, scaler)
-6:return $D(S^{F},X^{F},{\operatorname{do}}(Z\rightarrow\hat{z}))$triangleright Extract $Y(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow\hat{z}))$ and clamp, floor.
-B.3 Computing direct effect of sensitive on the target
-This function computes the direct effect $S$ on $Y$, i.e.
+[…]
+6:return $)\mathcal{D}(S^{F},X^{F},{\operatorname{do}}(Z\rightarrow\hat{z}$)) $\triangleright$ Extract $)Y(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow\hat{z}$)) and clamp, floor.
+Computing direct effect of sensitive on the target
+Algorithm 3 computes the direct effect $S$ on $Y$, i.e.
 ${\operatorname{do}}(S\rightarrow s^{CF})$ on the target $Y$ by computing $y({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$.
@@
 _direct_label
-Input: Factual data $D^{F}$, sensitive counterfactual data $D^{SCF}$ from Algorithm 1
+Input: Factual data $\mathcal{D}^{F}$, sensitive counterfactual data $\mathcal{D}^{SCF}$ from Algorithm 1
 Parameters: Flows model, data preparator
-Output: $D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$
+Output: $\mathcal{D}({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$
 1:Get scaler from preparator
-2:$U^{F}$ = model.flow.transform($D^{SCF}$) $\triangleright$ Helps get $U^{F}_{Y}$
-3:$U^{SCF}$ = model.flow.transform ($D^{SCF}$) $\triangleright$ Helps get $U^{SCF}_{S}$
+[…]
+[…]
 4:Get intervention indices $i_{\{X,T\}}$
-5:Clamp, floor categorical dimensions in $D^{F}[i_{\{X,T\}}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $D^{F}$
-6:Set data_tmp = clone($D^{SCF}$)
-7:Set data_tmp$[i_{\{X,T\}}]$ = intervention_vals$[i_{\{X,T\}}]$
-8:$U^{\mathcal{I}}$ = model.flow.transform (data_tmp) $\triangleright$ Helps get $U^{\mathcal{I}}_{\{X,Z\}}$
+5:Clamp, floor categorical dimensions in $\mathcal{D}^{F}[i_{\{X,T\}}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $\mathcal{D}^{F}$
+6:Set data_tmp = clone($\mathcal{D}^{SCF}$)
+[…]
+[…]
 9:Get combined $U_{\text{comb}}=[U^{SCF}_{S},U^{\mathcal{I}}_{\{X,Z\}},U^{F}_{Y}]$
-10:$D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F},Z\rightarrow z^{F}))$ = model.flow.transform.inverse($U_{\text{comb}}$)
-11:return $D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$triangleright Extract $y({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$ and clamp, floor.
-B.4 Computing path-specific treatment
-With this function we calculate the path-specific counterfactuals i.e.
-we compute the path-specific treatment $z^{PS}=z({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$.
+[…]
+11:return $\mathcal{D}({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$ $\triangleright$ Extract $y({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X,Z\rightarrow x^{F},z^{F}))$ and clamp, floor.
+Computing path-specific treatment
+Algorithm 4 calculates the path-specific counterfactuals i.e., the path-specific treatment $z^{PS}=z({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$.
 Algorithm 4 compute_path_specific_treatment
-Input: Factual data $D^{F}$, sensitive counterfactual data $D^{SCF}$ from Algorithm 1
+Input: Factual data $\mathcal{D}^{F}$, sensitive counterfactual data $\mathcal{D}^{SCF}$ from Algorithm 1
 Parameters: Flows model, data preparator
-Output: $D^{PS}=D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$
+Output: $\mathcal{D}^{PS}=D({\operatorname{do}}(S\rightarrow s^{CF}),{\operatorname{do}}(X\rightarrow x^{F}))$
 1:Get scaler from preparator
-2:$U^{F}$ = model.flow.transform($D^{SCF}$) $\triangleright$ Helps get $U^{F}_{Y},U^{F}_{Z}$
-3:$U^{SCF}$ = model.flow.transform ($D^{SCF}$) $\triangleright$ Helps get $U^{SCF}_{S}$
+[…]
+[…]
 4:Get intervention indices $i_{X}$
-5:Clamp, floor categorical dimensions in $D^{F}[i_{X}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $D^{F}$
-6:Set data_tmp = clone($D^{SCF}$)
+5:Clamp, floor categorical dimensions in $\mathcal{D}^{F}[i_{X}]$, save in intervention_vals $\triangleright$ Pre-apply clone on $\mathcal{D}^{F}$
+6:Set data_tmp = clone($\mathcal{D}^{SCF}$)
 7:Set data_tmp[$i_{X}$] = intervention_vals[$i_{X}$]
-8:$U^{\mathcal{I}}$ = model.flow.transform(data_tmp) $\triangleright$ Helps get $U^{\mathcal{I}}_{X}$
+[…]
 9:Get combined $U_{\text{comb}}=[U^{SCF}_{S},U^{\mathcal{I}}_{X},U^{F}_{Z},U^{F}_{Y}]$
-10:$D^{PS}$ = model.flow.transform.inverse($U_{\text{comb}}$)
-11:return $D^{PS}$triangleright Extract $z^{PS}$, clamp and floor categorical features before saving.
+[…]
+11:return $\mathcal{D}^{PS}$ $\triangleright$ Extract $z^{PS}$, clamp and floor categorical features before saving.
 Computing downstream effect of $z^{PS}$
 Note that once we have $z^{PS}$ from Algorithm 4, we can compute the downstream effect, i.e., $Y(s^{F},x^{F},{\operatorname{do}}(Z\rightarrow z^{PS}))$ using Algorithm 2.
-Appendix C Datasets, preprocessing and other details
-C.1 Datasets and their preprocessing
-  1. 1.
+Appendix C Experimental Setup
+C.1 Datasets and Preprocessing
   • HMDA: This dataset regarding mortgage applications covers several states in the United States, but we focused on just two states (New York and Texas) in the year 2017.
-Following similar preprocessing steps as done in [10], we binarized the gender column as $\{Male:1,Female:0\}$.
+Following similar preprocessing steps as done in (Cooper et al.
+2024), we binarized the gender column as $\{Male:1,Female:0\}$.
 Also, we binarized the race with originally $8$ possible values but utilized only $\{NativeWhite:0,White:1\}$.
-For the action taken with $8$ possible values, we filtered out ${3-8}$.
+For the action taken with $8$ possible values, we filtered out 38{3-8}.
 We kept the outcome as a $1$ Loan originated to mean the applicant and the bank decision agreed, and $2$, which is the application approved but not accepted, to mean a disagreement.
@@
 After the preprocessing, we had a total dataset of size $399354$ with $17$ features.
-  2. 2.
   • Home Credit: We downloaded this dataset from the kaggle competition.
-We used the application training datasets and selected features for our study based on the feature importance as done in [42].
+We used the application training datasets and selected features for our study based on the feature importance as done in (Matthys 2019).
 We utilized the resulting features selected by their random forest classifier and information entropy models.
-Based on this, we selected the following features : “EXT SOURCE 2, EXT SOURCE 3, EXT SOURCE 1, DAYS BIRTH, NAME EDUCATION TYPE, CODE GENDER, DAYS EMPLOYED, NAME INCOME TYPE, ORGANIZATION TYPE, AMT CREDIT, AMT ANNUITY, AMT GOODS PRICE, REGION POPULATION RELATIVE, TARGET".
-Based on this subset of data, we dropped the missing values in the dataset and converted the days of birth and days of employment to years using days of $\frac{\text{days of birth/employment}}{-365}$.
-We divided it by $-365$ because these values were negative.
+Based on this, we selected the following features : “EXT SOURCE 2, EXT SOURCE 3, EXT SOURCE 1, DAYS BIRTH, NAME EDUCATION TYPE, CODE GENDER, DAYS EMPLOYED, NAME INCOME TYPE, ORGANIZATION TYPE, AMT CREDIT, AMT ANNUITY, AMT GOODS PRICE, REGION POPULATION RELATIVE, TARGET”.
+Based on this subset of data, we dropped the missing values in the dataset and converted the days of birth and days of employment to years using $\frac{\text{days of birth/employment}}{-365}$.
+We divided it by 365-365 because these values were negative.
 The complete dataset has a size of $98859$ with $14$ features.
-  3. 3.
   • German Credit: We obtained this dataset from the Statlog website and restructured the personal status feature by consolidating the information on gender and marital status into binary gender categories.
@@
 The dataset had no missing values.
-(a) German Credit
-(b) Home Credit
-Figure 4: Treatment decision feature distributions across social groups show overlap, indicating it as a reasonable causal assumption for real-world settings.
 Remarks.
@@
 Hence, we use the difference between treatment values as $\mathcal{F}$ and the expectation $\mathcal{A}$ over each dataset $\mathcal{D}$ when considering disparity measures in Definitions 1 and 2.
-Furthermore, from Figure 4, we note that treatment feature distributions show significant overlap across the two sensitive groups.
-This observation shows that our overlap assumption is a reasonable one and holds in real-world settings.
-Table 7: Different loan approval datasets we used in the study with their sensitive attributes ($S$), treatment decisions ($Z$), outcome label ($Y$) and dataset size.
 Dataset
@@
 ($1000,21$)
-C.2 Losses
-  1. 1.
-    Loss Given Default (LGD): We estimate the loss incurred by the bank in the event an applicant defaults.
-Quantifying the loss will require additional analysis of other variables such as the value of collateral, instalments payments etc [25].
-There are a number of different ways to estimate $LGD$ however a common approach considers the exposure at default and the recovery rate that is:
+Table 5: Different loan approval datasets we used in the study with their sensitive attributes ($S$), treatment decisions ($Z$), outcome label ($Y$) and dataset size.
+C.2 Hyper-Parameter Tuning and Training of Causal Normalising Flows
+For the causal normalizing flows (Javaloy, Sánchez-Martín, and Valera 2024), we use the original work as inspiration and tune for the flow models’ inner dimensionality and learning rate Table 6.
+All other parameters are fixed, which are as follows: Adam optimizer with plateau scheduler, activation as ELU, layer type was Neural Spline Flow (NSF) and all the experiments were run for a maximum of $1000$ epochs.
+For other parameters, we refer to the GitHub code  [^2] https://github.com/psanch21/causal-flows.
+The number of nodes is specified in Table 7.
+An example of causal graph input to causal normalizing flows is given in Figure 5.
+Dataset dim-inner lr log-prob
+HMDA-NY $32\times 32$ $10^{-2}$ -61.67
+HMDA-TX $32\times 32$ $10^{-2}$ -63.49
+Home Credit $64$ $10^{-2}$ -41.87
+German Credit $64$ $10^{-3}$ -14.68
+Synthetic Loan $16\times 16\times 16\times 16$ $10^{-3}$ -16.51
+Table 6: Hyperparameters for causal Flows. The log-prob of the observational data distribution estimate was used to select the best hyperparameters.
+Dataset #rootnodes #covariates #treatment
+HMDA-NY 2 12 2
+HMDA-TX 2 12 2
+Home Credit 2 9 2
+German Credit 2 15 3
+Synthetic Loan 2 3 2
+Table 7: Number of root nodes, covariates, and treatment variables for training causal normalizing flows model. The dimensionality of the outcome label is 1.
+C.3 Hardware
+All experiments of hyper-parameter tuning and training of causal normalizing flows are done on an Ubuntu workstation with an Intel CPU.
+The code primarily uses PyTorch CPU and Python.
+We set the seed for PyTorch and Python to a fixed number and ran all experiments.
+A fixed seed is also used to split datasets into train, validation, and test sets - 80%, 10%, 10%.
+This seed is fed as input in a parameter file to train and run evaluations.
+AG
+EISLDYSensitive SSOutcome YYBlock XXBlock $Z$
+Figure 5: Causal graph for synthetic loan dataset introduced in Appendix D.
+While training the causal normalising flows, covariates $X$ and treatment decisions $Z$ are treated as multivariate blocks without any explicit assumptions on the cause-effect directionality within each block.
+Here, we have multiple sensitive attributes (independent root nodes) but for this work, we concentrate on intervening on just one sensitive attribute (“gender”).
+A similar analysis can be done by intervening on the attribute “age”.
+Appendix D Synthetic Validation Evaluation
+In this section, we consider a synthetic data generation process where we have access to the ground-truth generative process.
+Using this dataset, we aim to validate the performance of our causal framework, comparing it to the ground-truth oracle that has access to the exact causal relations.
+This analysis shows that, considering the underlying assumptions hold, our causal framework leveraging the CNF can capture unknown causal relations and reliably help audit existing treatment disparities and also mitigate potential discriminatory effects of such treatment decision disparities.
+D.1 Synthetic Dataset Generation
+Taking inspiration from (Karimi et al.
+2020; Majumdar and Valera 2024), we consider the 7-variable loan dataset modeling the credit dataset (Hofmann 1994) with an added outcome variable $Y$ (Equation 6).
+Similar to (Karimi et al.
+2020; Majumdar and Valera 2024), the equations are non-linear with non-additive exogenous variables.
+Gender $G$ and age $A$ are root-nodes, loan amount $L$, education-level $E$, income $I$, and savings $S$ are actionable with equations shown below.
+Additionally, we added a link between the savings $S$ to the loan amount and thresholded the outcome variable $Y$ directly on $G$ using two different thresholding scenarios.
+$\sigma(\cdot)$ represents the sigmoid function in the following equations.
+$\mathcal{N}(a,b)$ denotes the Gaussian distribution where a is the mean and b is the variance of the distribution, $\sqrt{b}$ denoting the standard deviation.
+Finally, $\operatorname{Gamma}(a,b)$ denotes the Gamma distribution where a is the concentration/shape and b is the scale, 1/b being the rate.
 […]
-    In our estimation, we assume that EAD is equivalent to the credit amount loaned out.
-EAD is a crucial metric that estimates the total loss a bank is exposed to when a loan defaults.
-However, it is important to note that the calculation of EAD can be more complex, as it may involve factors beyond the simple loan amount, such as commitment details [58].
-    The recovery rate, on the other hand, is a risk-adjusted measure that reflects the portion of the loan that the bank expects to recover after default.
-In our analysis, we considered it as the outcome of the loan (i.e., the actual repayment).
-    Finally, we formalize the expected loss given default for the bank as:
-    LGD $=\mathbb{E}_{\mathcal{D}^{\prime}}[(1-y)\times\text{amount}]$ (12)
-    where $y$ represents the applicant’s repayment probability.
-  2. 2.
-    Expected Simple Interest (ESI):
-    We estimate the loss of the individuals using the simple interest formula:
-    $\textbf{SI}=\frac{(\text{amount}\times\text{rate}\times\text{time})}{100}$
-    For the Home Credit dataset, we used slightly modified formulae to estimate the ESI for Home Credit, as the dataset did not include information on the rate and time, making ESI estimation less straightforward.
-We estimate ESI using annuity and loan amount as follows: we assume duration=15 years.
-Annuity can be written as :
-    annuity=(amount + interest)total durations(months)=(amount + $\text{annuity}=\frac{(\text{amount + interest})}{\text{total durations(months)}}=\frac{(\text{amount + interest})}{12\times\text{duration}}$
-    Using this the simple interest can be calculated as :
-    ESI(Home $\textbf{ESI}\text{(Home Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}[\text{annuity}\times 12\times\text{duration}-\text{amount}]$ (13)
-    In our formulae for the German Credit dataset, the loan duration is in months and we assume in our implementation that everyone was given the same interest rate of 10%, hence enabling us to rewrite the ESI as follows:
-    ESI(German $\textbf{ESI}\text{(German Credit)}=\mathbb{E}_{\mathcal{D}^{\prime}}\left[y\times\text{amount}\times\frac{\text{rate}}{100}\times\frac{\text{time}}{12}\right]$ (14)
-Appendix D Synthetic loan
-D.1 Synthetic dataset generation
-Taking inspiration from [29], we consider the 7-variable loan dataset modeling the credit dataset [22] with an added outcome variable $Y$.
-Similar to [29], the equations are non-linear with non-additive exogenous variables.
-The nodes, gender $G$ and age $A$ are root-nodes (non-actionable), loan amount $L$, education-level $E$, income $I$, and savings $S$ are actionable with equations shown below.
-Additionally, we added a link between the savings $S$ to the loan amount and thresholded the outcome variable $Y$ directly on $G$ using two different thresholding scenarios.
-The symbol $\sigma(\cdot)$ represents the sigmoid function in the following equations.
-Distribution N (a,b) denotes the Gaussian distribution where a is the mean and b is the variance of the distribution, $\sqrt{b}$ denoting the standard deviation.
-Finally, Gamma (a, b) denotes the Gamma distribution where a is the concentration/shape and b is the scale, 1/b being the rate.
-Root nodes : (15)
-$f_{G}:G=U_{G},\quad U_{G}~{}\sim\operatorname{Bernoulli}(0.5)$
-$f_{A}:A=-35+U_{A},\quad U_{A}~{}\sim\operatorname{Gamma}(10,3.5)$
-Covariates:
-$f_{E}:E=-0.5+\sigma(-1+0.5G+\sigma(0.1A)+U_{E}),$
-$\quad\quad U_{E}\sim\mathcal{N}(0,0.25)$
-$f_{I}:I=-4+0.1(A+35)+2G+GE+U_{I},U_{I}~{}\sim~{}\mathcal{N}(0,4)$
-$f_{S}:S=-4+1.5\mathbb{I}\{I>0\}I+U_{S},U_{S}~{}\sim~{}\mathcal{N}(0,5)$
-Treatments:
-$f_{L}:L=1+0.01(A-5)(5-A)+2(1-G)+\beta S+U_{L},$
-$\quad\quad U_{L}~{}\sim~{}\mathcal{N}(0,10),\beta\in\{0,0.03\}$
-$f_{D}:D=-1+0.1A+3(1-G)+L+U_{D},U_{D}~{}\sim~{}\mathcal{N}(0,9)$
-Outcome: two possibilities
-$f_{Y}:Y=\mathbb{I}\{\sigma(\delta(-L-D)+0.3(I+S+\alpha IS))$
-$\geq 0.5+\gamma(1-G)\}\quad\textbf{OR}$
-$f_{Y}:Y=\mathbb{I}\{\sigma(\delta(-L-D)+0.3(I+S+\alpha IS)$
-$+U_{Y}*\gamma(1-G))\geq 0.5\},U_{Y}~{}\sim~{}\mathcal{N}(0,5)$
-[…]
-$\alpha=1,\textbf{if}\left[\mathbb{I}\{I>0\}\land\mathbb{I}\{S>0\})\right]=1,$
-$\alpha=-1,\textbf{otherwise}$
-[…]
+Figure 6: Synthetic data generating process.
+Dataset Sensitive counterfactual Measure Treatment Disparity Effect on Repayment (%)
+Flows Oracle Flows Oracle
+Amount Duration Amount Duration $Y^{F}=0$ $Y^{F}=1$ $Y^{F}=0$ $Y^{F}=1$
+Balanced Female $\rightarrow$ Male TTD(-E) 1.91-1.91 4.44-4.44 1.95-1.95 4.94-4.94 $0.19$ $0.00$ $0.18$ $0.00$
+DTD(-E) 1.62-1.62 4.25-4.25 2.00-2.00 5.00-5.00 $0.17$ $0.00$ $0.18$ $0.00$
+Male $\rightarrow$ Female TTD(-E) +2.14+2.14 +4.64+4.64 +1.95+1.95 +4.95+4.95 $0.00$ $0.20$ $0.00$ $0.20$
+DTD(-E) +1.90+1.90 +4.05+4.05 +2.00+2.00 +5.00+5.00 $0.00$ $0.17$ $0.00$ $0.20$
+Imbalanced Female $\rightarrow$ Male TTD(-E) 1.92-1.92 4.61-4.61 1.95-1.95 4.95-4.95 $0.09$ $0.00$ $0.09$ $0.00$
+DTD(-E) 1.55-1.55 3.75-3.75 2.00-2.00 5.00-5.00 $0.07$ $0.00$ $0.09$ $0.0$
+Male $\rightarrow$ Female TTD(-E) +2.00+2.00 +4.69+4.69 +1.95+1.95 +4.95+4.95 $0.00$ $0.21$ $0.00$ $0.21$
+DTD(-E) +2.21+2.21 +5.45+5.45 +2.00+2.00 +5.00+5.00 $0.00$ $0.24$ $0.00$ $0.21$
 Table 8: Measuring treatment disparities and discrimination for synthetic data.
@@
 Results are reported as median values.
-Dataset Sensitive counterfactual Measure Treatment Disparity Effect on Repayment (%)
-Flows Oracle Flows Oracle
-Amount Duration Amount Duration $Y^{F}=0$ $Y^{F}=1$ $Y^{F}=0$ $Y^{F}=1$
-Balanced Female $\rightarrow$ Male TTD(-E) $-1.91$ $-4.44$ $-1.95$ $-4.94$ $0.19$ $0.00$ $0.18$ $0.00$
-DTD(-E) $-1.62$ $-4.25$ $-2.00$ $-5.00$ $0.17$ $0.00$ $0.18$ $0.00$
-Male $\rightarrow$ Female TTD(-E) $+2.14$ $+4.64$ $+1.95$ $+4.95$ $0.00$ $0.20$ $0.00$ $0.20$
-DTD(-E) $+1.90$ $+4.05$ $+2.00$ $+5.00$ $0.00$ $0.17$ $0.00$ $0.20$
-Imbalanced Female $\rightarrow$ Male TTD(-E) $-1.92$ $-4.61$ $-1.95$ $-4.95$ $0.09$ $0.00$ $0.09$ $0.00$
-DTD(-E) $-1.55$ $-3.75$ $-2.00$ $-5.00$ $0.07$ $0.00$ $0.09$ $0.0$
-Male $\rightarrow$ Female TTD(-E) $+2.00$ $+4.69$ $+1.95$ $+4.95$ $0.00$ $0.21$ $0.00$ $0.21$
-DTD(-E) $+2.21$ $+5.45$ $+2.00$ $+5.00$ $0.00$ $0.24$ $0.00$ $0.21$
-Table 9: Fairness in non-binary decision-making. Comparison of bank’s LGD and applicant’s ESI losses for balanced and unbalanced datasets under different policies.
-Sensitive Group Balanced Dataset Unbalanced Dataset
-LGD (DM) ESI (DM) LGD (DM) ESI (DM)
+Our framework using CNF can accurately capture disparities without access to ground-truth causal mechanisms.
+Sens. Group Balanced Data Unbalanced Data
+LGD ESI LGD ESI
 Males $3.09$ $0.58$ $3.30$ $0.58$
@@
 Females ($\mathcal{D}^{\mathrm{fair}}$) 3.14 $0.55$ 2.39 $0.47$
+Table 9: Fairness in non-binary decision-making. Comparison of the bank’s LGD and the applicant’s ESI losses (in DM) for balanced and unbalanced datasets under different treatments.
 D.2 Results
@@
 This allows us to compare the results given by our CNF-based causal framework with the oracle.
-D.2.1 Balanced dataset:
+Balanced dataset:
 Using the synthetic loan generation process described in the previous section, we set the parameters $\beta=0.03,\gamma=0.5,\delta=1,\eta=5$ to construct a balanced dataset containing $5000$ samples.
@@
 Designing such policies will enable a more fair and efficient decision-making pipeline.
-D.2.2 Unbalanced dataset
+Unbalanced dataset
 To create the imbalanced dataset, we adjusted the parameters to $\beta=0.03$, $\gamma=0.5$, $\delta=2$, and $\eta=5$, generating a dataset with $5000$ samples.
@@
 Unfair estimates are based on the factual distribution, while fair estimates use interventional treatment distribution where all applicants are treated as males.
+Appendix E Additional Real-World Evaluations
+Dataset
+Sensitive
+Counterfactual
+Measure Treatment Disparity Effect on Agreement($\%$)
+Amount ($\times 10^{3}$\$) $Y^{F}=0$ $Y^{F}=1$
+HMDA-NY Female $\rightarrow$ Male TTD(-E) +29.46+29.46 $2.00$ $0.40$
+DTD(-E) +6.65+6.65 $2.00$ $0.20$
+Male $\rightarrow$ Female TTD(-E) 29.93-29.93 $4.00$ $0.16$
+DTD(-E) 5.88-5.88 $2.00$ $0.14$
+HMDA-TX Female $\rightarrow$ Male TTD(-E) +19.43+19.43 $2.00$ $0.28$
+DTD(-E) +2.67+2.67 $2.00$ $0.13$
+Male $\rightarrow$ Female TTD(-E) 19.78-19.78 $3.00$ $0.18$
+DTD(-E) 3.65-3.65 $3.00$ $0.10$
+[…]
+Dataset
+Sensitive
+Counterfactual
+Measure Treatment Disparity Effect on Repayment($\%$)
+Annuity (INR) Amount (INR) $Y^{F}=0$ $Y^{F}=1$
+Home Credit Female $\rightarrow$ Male TTD(-E) +1008.12+1008.12 +7885.72+7885.72 $0.24$ $0.30$
+DTD(-E) +715.91+715.91 816.94-816.94 $0.27$ $0.00$
+Male $\rightarrow$ Female TTD(-E) 942.92-942.92 6659.68-6659.68 $2.30$ $0.00$
+DTD(-E) 594.56-594.56 +758.46+758.46 $1.30$ $0.00$
+Duration (months) Amount (DM) $Y^{F}=0$ $Y^{F}=1$
+German Credit Female $\rightarrow$ Male TTD(-E) +1.29+1.29 +253.34+253.34 $2.00$ $3.50$
+DTD(-E) +0.46+0.46 +164.66+164.66 $2.00$ $3.00$
+Male $\rightarrow$ Female TTD(-E) 0.79-0.79 243.11-243.11 $8.00$ $1.00$
+DTD(-E) 0.59-0.59 212.39-212.39 $8.00$ $0.60$
+Table 11: Measuring treatment disparities in S.2 and outcome effects.
+Analysis conducted on the entire, unsplit dataset.
+$Y^{F}=0$ indicates a label change from loan defaulted to repaid, $Y^{F}=1$ the reverse.
+Results show median INR (Indian Rupees) and DM (Deutsche Marks).
+Results on entire data consistent with trends in main paper.
+E.1 Analyzing Treatment Disparities on the Entire Data
+We repeat our evaluations conducted in the main paper regarding the auditing of historical treatment disparities and their unfair effect on outcomes.
+Here, we consider only the historical decision policy that gave loans and analyze the whole proportion of each lending dataset.
+[…]
+Results in Home Credit.
+We report the results in Table 11.
+Our analyses show that if females were treated like their male counterfactuals, they would have received higher annuities and loan amounts.
+The reverse pattern would apply to male applicants when treated as females.
+Interestingly, we see contrasting trends between total and direct disparities regarding loan amounts.
+Total disparity indicates conservative lending practices where female applicants are granted lower loan amounts and annuities.
+In contrast, direct disparity, attributed solely to gender, shows that female applicants receive lower loan amounts but higher annuities, suggesting elevated interest rates.
+These findings suggest that while banks may have implemented some affirmative treatment policies, disparities mediated through covariates influence lending decisions significantly.
+Importantly, we see minor downstream effects on outcomes across groups.
+Hence, the outcome differences do not justify the observed treatment disparities against females, particularly regarding $\operatorname{DTD}$, highlighting potential discrimination.
+Results in German Credit.
+From the results in Table 11, we observe a similar trend with the bank adopting more conservative practices toward female applicants.
+If females were treated like males, they would receive longer loan durations and higher loan amounts, with the reverse pattern observed for males.
+Significant disparity is seen in loan amounts, driven mainly by direct effects.
+This observation suggests that applicant covariates do not fully explain treatment disparities, and gender may have a direct influence, exposing a potential problematic mechanism.
+Interestingly, there is no disparity in installment rate treatments regarding $\operatorname{TTD}$ or $\operatorname{DTD}$, so we exclude this feature from Table 2 for clarity.
+When examining the impact of treatment disparities on outcomes, a key observation arises: treating males like their female counterparts would result in 8.0% of defaulting males repaying their loans.
+In contrast, only 1.0% of male repayers would default under this counterfactual treatment.
+Hence, our counterfactual analysis shows that the treatment disparities are not justified by outcomes, indicating potential discrimination.
+Dataset Predictor Accuracy DP EOD
+German Credit Accuracy 70 0.06 0.18
+DP 67.5 0.03 0.17
+EOD 66.5 0.02 0.15
+Home Credit Accuracy 69.2 0.14 0.13
+DP 70.8 0.0 0.05
+EOD 68.2 0.01 0.02
+HMDA-NY Accuracy 74.0 0.01 0.01
+DP 72.0 0.00 0.00
+EOD 72.0 0.00 0.01
+HMDA-TX Accuracy 77.4 0.01 0.01
+DP 77.1 0.00 0.00
+EOD 77.1 0.00 0.00
+Table 12: Accuracy and fairness (demographic parity: DP, equalized odds: EOD) measures on the test-set of unmitigated (Accuracy) and fairness post-processed (DP, EOD) predictors for binary predictions of loan granting.
+Dataset
+Binary Policy
+(accepted people)
+Sensitive
+Counterfactual
+Measure Treatment Disparity Effect on Agreement($\%$)
+Amount ($\times 10^{3}$\$) $Y^{F}=0$ $Y^{F}=1$
+HMDA-NY
+Historical
+Female $\rightarrow$ Male TTD(-E) +29.68+29.68 $2.18$ $0.41$
+DTD(-E) +6.77+6.77 $2.51$ $0.20$
+Male $\rightarrow$ Female TTD(-E) 29.98-29.98 $3.88$ $0.15$
+DTD(-E) 5.87-5.87 $1.77$ $0.14$
+Accuracy
+Female $\rightarrow$ Male TTD(-E) +27.92+27.92 $4.39$ $0.09$
+DTD(-E) +6.37+6.37 $5.26$ $0.06$
+Male $\rightarrow$ Female TTD(-E) 27.75-27.75 $2.34$ $0.05$
+DTD(-E) 6.50-6.50 $2.34$ $0.03$
+Dem. Parity
+Female $\rightarrow$ Male TTD(-E) +28.03+28.03 $5.48$ $0.08$
+DTD(-E) +6.52+6.52 $4.11$ $0.06$
+Male $\rightarrow$ Female TTD(-E) 27.68-27.68 $2.56$ $0.04$
+DTD(-E) 6.58-6.58 $2.56$ $0.03$
+Eq. Odds
+Female $\rightarrow$ Male TTD(-E) +27.96+27.96 $4.82$ $0.07$
+DTD(-E) +6.48+6.48 $3.61$ $0.05$
+Male $\rightarrow$ Female TTD(-E) 27.68-27.68 $2.56$ $0.04$
+DTD(-E) 6.59-6.59 $2.56$ $0.03$
+HMDA-TX
+Historical
+Female $\rightarrow$ Male TTD(-E) +19.38+19.38 $1.60$ $0.29$
+DTD(-E) +2.70+2.70 $1.64$ $0.14$
+Male $\rightarrow$ Female TTD(-E) 19.81-19.81 $3.01$ $0.17$
+DTD(-E) 3.65-3.65 $2.38$ $0.10$
+Accuracy
+Female $\rightarrow$ Male TTD(-E) +20.98+20.98 $0.00$ $0.02$
+DTD(-E) +2.95+2.95 $0.00$ $0.02$
+Male $\rightarrow$ Female TTD(-E) 21.31-21.31 $0.00$ $0.01$
+DTD(-E) 3.35-3.35 $0.00$ $0.01$
+Dem. Parity
+Female $\rightarrow$ Male TTD(-E) +20.98+20.98 $0.00$ $0.02$
+DTD(-E) +2.95+2.95 $0.00$ $0.02$
+Male $\rightarrow$ Female TTD(-E) 21.32-21.32 $0.00$ $0.01$
+DTD(-E) 3.35-3.35 $0.00$ $0.01$
+Eq. Odds
+Female $\rightarrow$ Male TTD(-E) +20.98+20.98 $0.00$ $0.02$
+DTD(-E) +2.95+2.95 $0.00$ $0.02$
+Male $\rightarrow$ Female TTD(-E) 21.32-21.32 $0.00$ $0.01$
+DTD(-E) 3.35-3.35 $0.00$ $0.01$
+[…]
+Dataset
+Binary Policy
+(accepted people)
+Sensitive
+Counterfactual
+Measure Treatment Disparity Effect on Repayment($\%$)
+Annuity (INR) Amount (INR) $Y^{F}=0$ $Y^{F}=1$
+Home Credit
+Historical
+Female $\rightarrow$ Male TTD(-E) +1005.26+1005.26 +7834.47+7834.47 $0.22$ $0.25$
+DTD(-E) +708.24+708.24 797.40-797.40 $0.22$ $0.08$
+Male $\rightarrow$ Female TTD(-E) 957.37-957.37 6778.88-6778.88 $2.46$ $0.05$
+DTD(-E) 604.94-604.94 +857.12+857.12 $1.36$ $0.08$
+Accuracy
+Female $\rightarrow$ Male TTD(-E) +1024.06+1024.06 +8527.55+8527.55 $0.34$ $0.14$
+DTD(-E) +684.38+684.38 946.94-946.94 $0.00$ $0.05$
+Male $\rightarrow$ Female TTD(-E) 930.77-930.77 7427.25-7427.25 $3.28$ $0.03$
+DTD(-E) 516.16-516.16 +1115.90+1115.90 $1.64$ $0.03$
+Dem. Parity
+Female $\rightarrow$ Male TTD(-E) +1024.06+1024.06 +8527.55+8527.55 $0.34$ $0.14$
+DTD(-E) +687.54+687.54 952.23-952.23 $0.00$ $0.05$
+Male $\rightarrow$ Female TTD(-E) 922.00-922.00 7263.01-7263.01 $2.93$ $0.03$
+DTD(-E) 537.34-537.34 +1046.03+1046.03 $1.58$ $0.06$
+Eq. Odds
+Female $\rightarrow$ Male TTD(-E) +1031.86+1031.86 +8606.90+8606.90 $0.18$ $0.13$
+DTD(-E) +694.17+694.17 972.88-972.88 $0.00$ $0.05$
+Male $\rightarrow$ Female TTD(-E) 924.41-924.41 7310.10-7310.10 $3.08$ $0.04$
+DTD(-E) 529.54-529.54 +1069.00+1069.00 $1.54$ $0.06$
+Duration (months) Amount (DM) $Y^{F}=0$ $Y^{F}=1$
+German Credit
+Historical
+Female $\rightarrow$ Male TTD(-E) +1.15+1.15 +272.15+272.15 $4.25$ $2.7$
+DTD(-E) +0.42+0.42 +133.39+133.39 $4.25$ $1.35$
+Male $\rightarrow$ Female TTD(-E) 0.87-0.87 237.97-237.97 $12.33$ $0.97$
+DTD(-E) 0.58-0.58 189.68-189.68 $12.33$ $0.48$
+Accuracy
+Female $\rightarrow$ Male TTD(-E) +0.87+0.87 +175.69+175.69 $0.0$ $3.51$
+DTD(-E) +0.36+0.36 +112.62+112.62 $0.0$ $1.75$
+Male $\rightarrow$ Female TTD(-E) 0.67-0.67 171.70-171.70 $6.90$ $0.70$
+DTD(-E) 0.42-0.42 167.69-167.69 $3.45$ $0.70$
+Dem. Parity
+Female $\rightarrow$ Male TTD(-E) +0.87+0.87 +175.69+175.69 $0.0$ $3.51$
+DTD(-E) +0.36+0.36 +112.62+112.62 $0.0$ $1.75$
+Male $\rightarrow$ Female TTD(-E) 0.66-0.66 207.17-207.17 $8.70$ $0.81$
+DTD(-E) 0.39-0.39 174.30-174.30 $4.35$ $0.81$
+Eq. Odds
+Female $\rightarrow$ Male TTD(-E) +0.94+0.94 +189.89+189.89 $0.0$ $3.64$
+DTD(-E) +0.36+0.36 +116.74+116.74 $0.0$ $1.82$
+Male $\rightarrow$ Female TTD(-E) 0.67-0.67 212.44-212.44 $9.1$ $0.81$
+DTD(-E) 0.42-0.42 189.68-189.68 $4.54$ $0.81$
+Table 14: Measuring treatment disparities in S.2 and outcome effects.
+$Y^{F}=0$ indicates a label change from loan defaulted to repaid, $Y^{F}=1$ the reverse.
+Results show median INR (Indian Rupees) and DM (Deutsche Marks).
+The policy ”Historical” uses the whole test set since the data collected under the original policy is for people who were given the loan.
+Accuracy indicates that an accuracy-optimized predictor is used to give out loans to a subset of the test population.
+DP and EOD indicate fairness post-processed predictors used for the loan binary decision-making.
+Ensuring fairness of binary predictions cannot directly mitigate disparities in treatment decisions.
+E.2 Complementarity of Treatment and Predictive Fairness Across Various Predictors
+In this section, we analyze how different binary prediction models used for making binary predictions (giving loans) impact the treatment disparity values, assuming the treatment decision policy remains fixed to the past policy.
+Through this evaluation, we show that if the treatment decision policy remains the same, simply intervening for the fairness of binary decisions does not mitigate disparities that exist in treatment decisions.
+Here, other than considering the past historical policy that was used to collect the dataset (the entire dataset is for positive lending decisions; we do not have access to data of individuals rejected for a loan), we consider that the binary predictor is applied to simulate making automated binary decisions lending decisions (Corbett-Davies et al.
+2017).
+Practical considerations.
+We split each dataset 60-40 for train-test, then use the training data subset to train a logistic regression predictor that aims to make binary predictions that improve accuracy based on historically observed outcomes $Y$.
+Then, we use the fairlearn (Weerts et al.
+2023) package to post-process this predictor for (i) demographic parity and (ii) equalized odds, given the historical dataset.
+For post-processing, we set a particular seed to ensure reproducibility when the fairness algorithm uses randomization.
+For data pre-processing, we one-hot encode categorical features and use standard scaling on continuous features.
+The train-test data split was performed using stratified splitting regarding the label $Y$.
+The base logistic regression and latter fairness post-processing were optimized for balanced accuracy.
+The post-processing algorithm followed (Hardt, Price, and Srebro 2016), using a grid size of 5000.
+The different predictive metrics of the trained predictors are shown in Table 12.
+S.1 HMDA Analysis
+We provide the results in Table 13.
+Across the HMDA-NY and HMDA-TX datasets, the Accuracy, Demographic Parity (DP), and Equalized Odds (EOD) policy models show persistent gender-based treatment disparities that remain close in magnitude to those under the Historical policy.
+In both states, females receive significantly lower loan amounts than males—about \$28K in NY and \$21K in TX under TTD even after applying fairness-aware adjustments.
+However, the outcome-level effects of these disparities diverge between the two states.
+In NY, counterfactual analysis reveals that a meaningful percentage of males would have agreed with loan terms had they received female-like treatment up to 2.34% under the Accuracy model (TTD and DTD), with very low disagreement rates suggesting that female-assigned terms may result in more acceptable or calibrated offers.
+This agreement gain persists across DP and EOD models, albeit with slight variations, indicating that these fairness adjustments do little to alter the underlying outcome asymmetry.
+In contrast, the TX dataset exhibits minimal sensitivity in counterfactual agreement under all models beyond Historical.
+Agreement shifts are essentially zero under Accuracy, DP, and EOD in both directions, suggesting that while treatment disparities still exist, they do not meaningfully affect borrower agreement in Texas.
+Taken together, these results highlight that fairness-aware models do not eliminate gender-based disparities in treatment, and particularly in NY, the terms typically offered to females may lead to better borrower agreement outcomes—raising important considerations for fairness beyond simple parity in allocation.
+S.2 Home Credit Analysis
+We show the results in Table 14.
+For the Home Credit dataset, all binary policy models show consistent disparities in treatment between male and female applicants.
+Across models, male applicants consistently receive higher annuities and loan amounts compared to females (e.g., under the Historical model, Female $\rightarrow$ Male counterfactuals show increases of INR1005.26 in annuity and INR7834.47 in amount).
+However, we also consistently observe that a non-trivial percentage of male applicants would have repaid their loans if given the terms offered to their female counterfactuals.
+For example, under the Accuracy model, 3.28% of males would have repaid under female terms ($Y^{F}=0$), compared to just 0.03% who would have defaulted ($Y^{F}=1$), indicating that female-like terms may be more conservative, supporting higher repayment likelihoods.
+On the other hand, very small percentages of females would have repaid if given male terms (often ¡0.4%), indicating that the male treatment regime is not more advantageous in terms of outcomes.
+It may increase default risk.
+Overall, the fairness-aware policies (DP and EOD) reduce neither the disparities in loan amounts nor the trends of repayment outcomes.
+This specific analysis clearly shows that post-hoc binary prediction fairness adjustments do not fully correct for treatment disparities or their unfair impact on outcomes, critically highlighting the complementary nature of binary decision fairness and treatment decision fairness.
+S.2 German Credit Analysis
+The results in Table 14 for the German Credit dataset reflect the trends observed in the main paper: there are systematic differences in treatment terms and their outcome effects across gender.
+Across the different predictors tested, males are generally offered longer and larger loans.
+However, regarding repayment, we see that a large proportion of males would have repaid their loans under female-like terms.
+While binary predictive fairness-aware models (DP, EOD) reduce this repayment gap somewhat (e.g., $Y^{F}=0$ for males drops to 9.1% for EOD), the general trend of disparity and its potentially unfair effect on outcomes remains.
+Moreover, disparity values in treatment terms, e.g., in amount, increase for the EOD-fair predictor compared to a predictor optimized only for accuracy.
+These results suggest that fairness in lending should not be judged only by parity in binary decisions, but also regarding the treatment terms provided and how well the terms align with successful repayment.
+Data Predictor Train on Bin. Pred. Utility
+($Y=1$) Pre-proc. Female Male Overall
+(loans) Accuracy DP EOD LGD ESI LGD ESI LGD ESI
+Home
+Credit
+✗: 92.4%
+✓: 92.5%
+Accuracy ✗ (26258) 69.2 0.14 0.13 631167.83 446661.90 651661.23 4771788.94 636986.18 4538969.85
+✓ (26240) 69.1 0.13 0.12 630534.25 4444490.55 641463.72 4513145.93 633674.39 4464215.89
+DP ✗ (27072) 70.8 0.0 0.05 633404.55 4451011.27 632518.60 4728754.74 633112.83 4542463.90
+✓ (27027) 70.6 0.0 0.04 631624.03 4445472.78 622713.60 4479214.72 628695.43 4456562.79
+EOD ✗ (25864) 68.2 0.01 0.02 635313.03 4455243.59 639333.13 4742956.96 636622.24 4548941.77
+✓ (25500) 67.3 0.01 0.02 634544.57 4450423.15 630188.55 4491738.00 633127.24 4463865.90
+German
+Credit
+✗: 70%
+✓: 71.7%
+Accuracy ✗ (238) 70.0 0.06 0.18 290.79 274.62 733.49 471.94 608.86 416.39
+✓ (240) 69.2 0.04 0.21 282.36 277.70 464.74 373.13 412.31 345.70
+DP ✗ (213) 67.5 0.03 0.17 290.79 274.62 714.79 438.54 581.42 386.98
+✓ (239) 68.5 0.0 0.19 445.30 281.67 475.87 364.84 466.66 339.78
+EOD ✗ (210) 66.5 0.02 0.15 299.74 277.88 643.84 457.89 537.33 402.17
+✓ (229) 67.0 0.03 0.2 290.79 271.89 483.76 355.67 427.30 331.16
+Table 15: Analyzing test-set binary prediction metrics and treatment decision downstream utilities of bank and applicants (mean across female, male, and overall), comparing training+analysis on original data (✗) against our proposed pre-processed treatment-fair-augmented data (✓).
+$Y=1$ proportions and the number of granted loans across datasets remain similar regardless of pre-processing, allowing for comparison.
+Training on pre-processed data keeps similar binary predictive metrics but greatly improves the downstream overall utility of the bank (LGD) and applicants (ESI).
+E.3 Training Binary Predictors on Treatment-Fair Pre-Processed Data
+We evaluate the effects of training on our proposed treatment-fair pre-processed data on both the fairness and utility of downstream binary predictors designed to make binary lending decisions.
+Specifically, we compare training binary predictors $\hat{Y}=h_{y}(S,X,Z)$ on the original treatment-unfair data $\mathcal{D}$ and our pre-processed treatment-fair data $\mathcal{D}^{\mathrm{fair}}$.
+For each data, we split it into a 604060-40 train-test subsets, train (and post-process for fairness), and analyze on the test set.
+We again consider a logistic regression and use balanced accuracy.
+For fairness, we perform post-processing for demographic parity (DP) or equalized odds (EOD).
+The practical considerations mirror those shown in Appendix E.2.
+Table 15 reports results across the Home Credit and German Credit datasets, highlighting metrics related to predictive performance (accuracy, demographic parity (DP), and equalized odds (EOD)), as well as utility outcomes measured in terms of expected loss to the bank (LGD) and expected interest paid by applicants (ESI) for female applicants, male applicants, and over all applicants.
+As discussed in the main paper (Sec.
+5.2), we pre-process the data by augmenting fairer treatment assignments and their outcome effects on $Y$ for the male subgroup since our analyses revealed historically riskier loan terms offered to this group.
+Home Credit.
+Training on pre-processed data does not significantly affect binary predictive performance, supporting its use as a foundation for downstream model training.
+When the binary predictor is trained solely for accuracy, binary fairness metrics show slight improvement after pre-processing (DP: 0.14 to 0.13, EOD: 0.13 to 0.12), with a negligible drop in accuracy (69.2% to 69.1%).
+Similar trends hold for the DP- and EOD-trained predictors.
+Crucially, pre-processing successfully achieves fairer treatments and consistent improvements in downstream utility across predictors.
+Specifically, looking at the overall expected LGD and ESI across all applicants, we see that training the predictor on our pre-processed data reduces the expected bank loss and the expected interest paid by applicants, providing benefits for both stakeholders.
+Looking at each group, we see a negligible impact on female applicants and a major reduction for male applicants (expectedly since we intervene on this group for pre-processed $\mathcal{D}^{\mathrm{fair}}$), resulting in narrower utility gaps between the groups.
+German Credit.
+We observe similar trends where training binary predictors on pre-processed data generally does not significantly affect predictive performance.
+For example, the Accuracy-trained predictor shows an improvement in demographic parity (DP: 0.06 to 0.04) but a slight worsening in equalized odds (EOD: 0.18 to 0.21), alongside a minor drop in accuracy (70.0% to 69.2%).
+From a utility standpoint, we see our desired effect of mitigating treatment disparities: pre-processing consistently reduces both the bank’s expected losses (LGD) and applicants’ effective interest rates (ESI) overall, reflecting greater efficiency for the lender and fairer loan terms for borrowers.
+We see similar trends across the different predictors.
+Moreover, the gaps in LGD and ESI between female and male applicants have substantially reduced.
+These results indicate the benefit of training on the pre-processed data: the data-driven decision-making system can now provide loans with better terms that benefit all stakeholders.
+Takeaway.
+Pre-processing to correct treatment disparities does not significantly impact binary predictive performance.
+Still, it substantially improves treatment fairness and mitigates its downstream effects, making it a practical foundation for model training.
+Our approach specifically targets disparities in treatment and their resulting impact on financial outcomes.
+Training on pre-processed data consistently reduces both the bank’s expected losses (LGD) and applicants’ effective interest rates (ESI) overall, thereby benefiting all stakeholders.
+These improvements also lead to narrower utility gaps between female and male applicants, indicating a more equitable distribution of financial burden.
+While our results show that treatment-fair pre-processed data can be effectively used to train binary predictors with fairness constraints, improvements in treatment disparity do not always translate into gains in predictive fairness metrics, further indicating the complementary nature of fairness in binary decisions and non-binary treatment decisions.
+Future work should explore more nuanced in-processing methods that jointly address both treatment and predictive fairness to enable more comprehensive fairness improvements while accounting for the utilities of diverse stakeholders.